### v1

# Mapping 1,000+ Language Models via the Log-Likelihood Vector

Abstract
To compare autoregressive language models at scale, we propose using log-likelihood vectors computed on a predefined text set as model features. This approach has a solid theoretical basis: when treated as model coordinates, their squared Euclidean distance approximates the Kullback–Leibler divergence of text-generation probabilities. Our method is highly scalable, with computational cost growing linearly in both the number of models and text samples, and is easy to implement as the required features are derived from cross-entropy loss. Applying this method to over 1,000 language models, we constructed a “model map,” providing a new perspective on large-scale model analysis.
Mapping 1,000+ Language Models via the Log-Likelihood Vector
Momose Oyama1,2 Hiroaki Yamagiwa1 Yusuke Takase1  Hidetoshi Shimodaira1,2 1Kyoto University 2RIKEN {oyama.momose, hiroaki.yamagiwa, y.takase}@sys.i.kyoto-u.ac.jp, shimo@i.kyoto-u.ac.jp
1 Introduction
Language models have been evolving rapidly, and their community has expanded significantly. To understand this landscape and its future directions, it is essential to systematically analyze model similarity and positioning based on language modeling principles. On the Hugging Face Hub, models are categorized by name and attributes, while other studies assess similarity based on outputs Yax et al. (2024) or activations Zhou et al. (2025). Leaderboards Beeching et al. (2023); Fourrier et al. (2024); Chiang et al. (2024) are commonly used to assess model standings.
Since language models are probability models, we propose representing each model using coordinates that capture the geometric structure of the space of probability distributions. Concretely, we define a language model’s coordinates as its log-likelihood vector across a large collection of texts. Figure 1 shows a model map obtained through dimensionality reduction applied to the coordinates of 1,018 language models. This visualization reveals that models of the same type tend to cluster together, while models in close proximity often share the same primary text category, forming a continuous distribution across the map.
Figure 1: Map of 1,018 language models. Their log-likelihood vectors are visualized using t-SNE. (Top) Colors indicate model types. (Bottom) Colors indicate the model’s “primary text category,” the text category where the model achieves the highest standardized log-likelihood score among 17 text categories. See Section 4 for details.
meta-llama/Meta-Llama-3-8B KL (/byte) google/codegemma-2b KL (/byte) deepseek-ai/deepseek-llm-7b-base KL (/byte) Undi95/Meta-Llama-3-8B-hf 0.003 (0.000) deepseek-ai/deepseek-coder-1.3b-instruct 1654.767 (2.455) deepseek-ai/deepseek-moe-16b-base 130.715 (0.194) dfurman/Llama-3-8B-Orpo-v0.1 6.994 (0.010) bigcode/starcoderbase-1b 1812.768 (2.690) deepseek-ai/deepseek-llm-7b-chat 245.408 (0.364) migtissera/Tess-2.0-Llama-3-8B 28.829 (0.043) deepseek-ai/deepseek-coder-1.3b-base 2115.462 (3.139) deepseek-ai/deepseek-moe-16b-chat 247.889 (0.368) freewheelin/free-llama3-dpo-v0.2 67.783 (0.101) bigcode/gpt_bigcode-santacoder 2644.348 (3.924) deepseek-ai/DeepSeek-V2-Lite 306.530 (0.455) jondurbin/bagel-8b-v1.0 110.365 (0.164) deepseek-ai/deepseek-coder-6.7b-instruct 2675.004 (3.969) deepseek-ai/ESFT-vanilla-lite 307.487 (0.456) migtissera/Llama-3-8B-Synthia-v3.5 128.141 (0.190) Qwen/CodeQwen1.5-7B-Chat 2759.243 (4.094) deepseek-ai/DeepSeek-V2-Lite-Chat 585.090 (0.868) nvidia/Llama3-ChatQA-1.5-8B 128.776 (0.191) NTQAI/Nxcode-CQ-7B-orpo 2767.578 (4.106) mistralai/Mistral-7B-Instruct-v0.1 913.708 (1.356) ruslanmv/Medical-Llama3-8B 138.566 (0.206) Salesforce/codegen-6B-multi 2859.701 (4.243) statking/zephyr-7b-sft-full-orpo 1089.127 (1.616) FairMind/Llama-3-8B-4bit-UltraChat-Ita 199.748 (0.296) bigcode/starcoderbase-7b 2990.326 (4.437) Severian/ANIMA-Phi-Neptune-Mistral-7B 1140.161 (1.692) NousResearch/Hermes-2-Theta-Llama-3-8B 222.357 (0.330) deepseek-ai/deepseek-coder-6.7b-base 3282.352 (4.870) sethuiyer/Medichat-Llama3-8B 1157.599 (1.718)
Table 1: Top 10 nearest neighbors among the 1,018 language models for each model listed in the first row. The KL divergence is computed using formula (3) in Section 2. The values in parentheses represent the KL divergence measured in bits per byte of text (Section 3.5), obtained by multiplying the KL divergence by 0.001484. Tables for the nearest neighbors of the models labeled in the top panel of Fig. 1 are provided in Appendix I.
We find that the distances in our defined coordinate system accurately capture relationships among language models. On this map, each point represents a single model, with those having similar text-generation probability distributions appearing closer together and those with more distinct distributions positioned farther apart. In Section 2, we show that the squared Euclidean distance in this coordinate system approximates the Kullback-Leibler (KL) divergence among models. Table 1 lists the nearest neighbors for each language model. For example, many of the closest neighbors of meta-llama/Meta-Llama-3-8B AI@Meta (2024) also contain Llama-3 in their names.
Several studies have explored methods for comparing language models (see Appendix A). In particular, prior work on comparing generated text includes approaches that construct phylogenetic trees based on model-generated text (Yax et al., 2025) and approaches that measure differences in text-generation probabilities conditioned on given prompts using KL divergence (Melamed et al., 2024). However, these methods require generating text with each model, thus incurring the cost of pairwise distance computations, which becomes prohibitively expensive at large scale. By contrast, our method does not involve actual text generation; instead, we compute generation probabilities on a predefined text corpus. This enables us to derive model coordinates without pairwise comparisons, allowing efficient large-scale comparison of many models.
To gain insights from visualizing model attributes on the model map, in Section 4, we analyze various attributes [^1] Attributes include model type, primary text category, model performance, model size, and creation date. and their relationships. Additionally, by comparing log-likelihood and benchmark performance, we demonstrate the ability to detect data leakage. Then, in Section 5, we treat the log-likelihood vector as a feature and show how it can predict benchmark performance. Finally, in Section 6, we validate the theoretical relationship between model coordinates and KL divergence through experiments.
2 Mapping Language Models into the Space of Text Probability Distributions
In this section, we present our proposed method. Sections 2.2 and 2.3 introduce model feature vectors derived from text-generation probabilities. Section 2.4 demonstrates that the squared Euclidean distance in the coordinate system built using these features approximates the KL divergence between models. Section 2.5 offers an interpretation of the resulting model coordinates. An extension of this method, which defines model coordinates using the sequence of conditional probabilities for generating a given token sequence, is presented in Appendix E.
2.1 Autoregressive language models
Let $\mathcal{X}$ be the set of all possible texts, and let $\mathcal{V}$ be the token vocabulary. A text $x\in\mathcal{X}$ is represented as a sequence of tokens:
$x=(y_{1},\dots,y_{n}),\quad y_{t}\in\mathcal{V}.$
Denoting the maximum text length by $n_{\textrm{max}}$, we have $\mathcal{X}=\bigcup_{n=0}^{n_{\textrm{max}}}\mathcal{V}^{n}.$ We consider a set of $K$ language models $\{p_{i}\}_{i=1}^{K}$. With $y_{0}$ denoting the beginning-of-sequence (BOS) token, each language model $p_{i}$ predicts the next token $y_{t}$ given the preceding token sequence $y^{t-1}=(y_{0},\dots,y_{t-1})$. Thus, the conditional probability defined by $p_{i}$ is given by
$y_{t}\sim p_{i}(y_{t}\mid y^{t-1}),\quad t=1,\dots,n.$
Accordingly, the probability of a text $x$ under model $p_{i}$, denoted $x\sim p_{i}$, is
$p_{i}(x)=\prod_{t=1}^{n}p_{i}(y_{t}\mid y^{t-1}).$
In addition to the $K$ language models $p_{1},\dots,p_{K}$, we introduce a language model $p_{0}$ that represents an underlying distribution for theoretical purposes. We assume we have a dataset (corpus)
$D=(x_{1},x_{2},\dots,x_{N})\in\mathcal{X}^{N},$
consisting of $N$ texts, where each text is independently drawn from $p_{0}$.
2.2 Log-likelihood vector
For a model $p_{i}$, the probability of generating a text $x$ is denoted $p_{i}(x)$. Following the convention in statistical model selection, we refer to $p_{i}(x)$ as the likelihood of model $p_{i}$ given the text $x$. The log-likelihood is then defined as
$\ell_{i}(x)=\sum_{t=1}^{n}\log p_{i}(y_{t}\mid y^{t-1}).$
In language model implementations, $-\ell_{i}(x)$ corresponds to the cross-entropy loss for the text $x$, and $\exp(-\ell_{i}(x)/n)$ is known as the perplexity.
Our approach is straightforward. Given that the dataset $D$ consists of $N$ texts, we use the log-likelihood vector
$\bm{\ell}_{i}=(\ell_{i}(x_{1}),\dots,\ell_{i}(x_{N}))^{\top}\in\mathbb{R}^{N}$
as the feature vector for model $p_{i}$. The first step in our model analysis is to construct the log-likelihood matrix
$\bm{L}=(\bm{\ell}_{1},\dots,\bm{\ell}_{K})^{\top}\in\mathbb{R}^{K\times N}$
by stacking the vectors $\bm{\ell}_{i}$ for the $K$ models.
2.3 Double centering
As a preprocessing step for model analysis, we apply a technique called double centering Borg and Groenen (2005) to $\bm{L}$. First, we perform row-wise centering. The mean of each row, referred to as the mean log-likelihood, is given by
¯$\bar{\ell}_{i}=\sum_{s=1}^{N}\ell_{i}(x_{s})/N.$
Subtracting this value from each component of $\bm{\ell}_{i}$, we define the centered log-likelihood vector $\bm{\xi}_{i}=(\xi_{i1},\ldots,\xi_{iN})^{\top}\in\mathbb{R}^{N}$, where
$\xi_{is}:=\ell_{i}(x_{s})-\bar{\ell}_{i},\quad s=1,\ldots,N.$
Next, we apply column-wise centering to the matrix of centered feature vectors $(\bm{\xi}_{1},\ldots,\bm{\xi}_{K})^{\top}$. The mean vector is $\bar{\bm{\xi}}=\frac{1}{K}\sum_{i=1}^{K}\bm{\xi}_{i},$ and by subtracting this vector from each $\bm{\xi}_{i}$, we define the double-centered log-likelihood vector
$\bm{q}_{i}=\bm{\xi}_{i}-\bar{\bm{\xi}}.$
For further details, see Appendices B and C.
2.4 Kullback-Leibler divergence
[…]
$\mathrm{KL}(p_{i},p_{j})$ $=\sum_{x\in\mathcal{X}}p_{i}(x)\log\frac{p_{i}(x)}{p_{j}(x)}$
$=\mathbb{E}_{x\sim p_{i}}\bigl{(}\ell_{i}(x)-\ell_{j}(x)\bigr{)}.$ (1)
We assume the dataset $D$ is generated from an unknown underlying model $p_{0}$ and that the models $p_{i}$ and $p_{j}$ provide good approximations of $p_{0}$. Under this assumption, the KL divergence can be approximated as follows:
$2\,\mathrm{KL}(p_{i},p_{j})\approx\mathrm{Var}_{x\sim p_{0}}\bigl{(}\ell_{i}(x)-\ell_{j}(x)\bigr{)}.$ (2)
While the definition of KL divergence in (1) involves the expectation of $\ell_{i}(x)-\ell_{j}(x)$, the approximation in (2) takes the form of a variance. This result is somewhat surprising yet quite insightful. Notably, although KL divergence is not symmetric in the two models, the approximation in (2) is symmetric. We estimate (2) from the dataset $D$ as
$2\,\mathrm{KL}(p_{i},p_{j})\approx\|\bm{q}_{i}-\bm{q}_{j}\|^{2}/N.$ (3)
Thus, if we regard the model coordinates of $p_{i}$ as $\bm{q}_{i}/\sqrt{N}$ by scaling with $N$, then the squared Euclidean distance between two points approximates $2\,\mathrm{KL}(p_{i},p_{j})$.
The main results, namely (2) and (3), are proved in Appendix D using the theory of exponential family of distributions (Barndorff-Nielsen, 2014; Efron, 1978, 2022; Amari, 1982), similar to the discussion on the relationship between the norm of embeddings and KL divergence (Oyama et al., 2023). Although the concepts of model map and model coordinates have been discussed in statistics (Shimodaira, 1993; Shimodaira and Cao, 1998; Shimodaira, 2001), and there have been a few applications of model maps (Shimodaira and Hasegawa, 2005; Shimodaira and Terada, 2019), they have seldom been utilized in practice.
2.5 Model coordinates
We primarily use $\bm{q}_{i}$ as the feature vector of model $p_{i}$ and refer to it as the model coordinates. As shown in (3), the squared Euclidean distance in the $\bm{q}$-coordinate system approximates the KL divergence between language models, [^3] That is, $\|\bm{q}_{i}-\bm{q}_{j}\|^{2}\approx 2N\,\mathrm{KL}(p_{i},p_{j})$. For simplicity, we omit the constant scaling factor in expressions of this type. indicating that $\bm{q}_{i}$ represents the position of $p_{i}$ in the space of probability distributions. Since $\bm{\xi}_{i}$ differs from $\bm{q}_{i}$ only by an offset from the origin, $\bm{\xi}_{i}$ also serves as a model coordinate, and $\|\bm{q}_{i}-\bm{q}_{j}\|^{2}=\|\bm{\xi}_{i}-\bm{\xi}_{j}\|^{2}$. However, we prefer $\bm{q}_{i}$ for its more interpretable components and thus adopt it throughout this paper.
For visualization purposes, we mainly use $\bm{\ell}_{i}$ as the coordinates of the model map, as $\bm{\ell}_{i}$ can be intuitively interpreted as encoding $\sqrt{N}\,\bar{\ell}_{i}$ in the “height” dimension and $\bm{q}_{i}$ in the “horizontal” dimensions. As shown in Appendix D.6,
$\|\bm{\ell}_{i}-\bm{\ell}_{j}\|^{2}=\|\bm{q}_{i}-\bm{q}_{j}\|^{2}+N(\bar{\ell}_{i}-\bar{\ell}_{j})^{2},$ (4)
which means the squared Euclidean distance in the $\bm{\ell}$-coordinate system can be decomposed into the sum of $2N\,\mathrm{KL}(p_{i},p_{j})$ and $N\,(\bar{\ell}_{i}-\bar{\ell}_{j})^{2}$.
3 Experimental Setup
Figure 2: Text embeddings for 10,000 texts in dataset $D$, computed via simcse-roberta-large Gao et al. (2021b) and visualized with t-SNE. Colors indicate 17 text categories.
Figure 3: The double-centered log-likelihood matrix $\bm{Q}$, with rows and columns reordered by hierarchical clustering. Each row corresponds to one of the 1,018 models, color-coded by model type. Each column represents one of the 10,000 texts, color-coded by text category.
Figure 4: Hierarchical clustering of the top 100 most-downloaded models, based on their feature vectors $\bm{q}_{i}$. Model names are color-coded by model type.
We describe the key components of our experiment. In particular, Section 3.1 explains the procedure for selecting the 10,000 texts used to compute the model coordinates, and Section 3.2 discusses how we selected the 1,018 language models. Further details are given in Appendix G.
3.1 Selection of text data
The texts used for computing the language models’ coordinates were extracted from the Pile Gao et al. (2020), with five categories of copyrighted material removed. [^4] https://huggingface.co/datasets/monology/pile-uncopyrighted This yielded a dataset $D$ consisting of 10,000 texts, each tagged with a category label from the Pile. Figure 2 visualizes these texts.
To build the dataset, we began by dividing the first 1M texts from the Pile Uncopyrighted corpus into 1,024-byte chunks (UTF-8 encoded). In cases where decoding errors occurred, we truncated by one byte at a time. Chunks smaller than 256 bytes were discarded, resulting in about 5.7M valid chunks. From these, we randomly sampled 10,000 texts to create the final dataset used for computing model coordinates. The average length of these 10,000 texts was 972.3188 bytes.
3.2 Selection of language models
We used $K=1{,}018$ language models in total. Of these, 1,000 were selected from models listed on Open LLM Leaderboard v1. Specifically, we considered CausalLM models ranging from 1B to 13B parameters and ranked them by their number of downloads over the 30 days preceding February 1, 2025.555Hugging Face’s API provides the total number of downloads in the past 30 days. We initially selected the top 1,100 models by download count and attempted log-likelihood calculations. Among these, 1,011 successfully produced valid log-likelihood values, and we chose the 1,000 most frequently downloaded from that set. In addition, we included 18 models from the DeepSeek language model series. We obtained information on model parameter sizes and architectures from the Leaderboard. Appendix G provides basic information on the selected models and details on how model types were defined. A complete list of models used in this study is given in Appendix L.
3.3 Computation of the log-likelihood
The log-likelihood matrix $\bm{L}$ was computed in float16 precision, with the bottom 2% of values clipped. This clipping mitigates the large impact of extremely low likelihoods on (3). After computing $\bm{L}$, we applied row-wise and column-wise centering to obtain the double-centered log-likelihood matrix $\bm{Q}$. Figure 3 visualizes $\bm{Q}$. Each value in the matrix can be interpreted in two ways: as the relative probability of a text for each model or as the relative likelihood of a model for each text. Both models and texts exhibit clustering patterns. Figure 4 shows a dendrogram of the top 100 models. We examined the effective dimension from the perspective of feature vector dimensionality reduction and found that the cumulative contribution ratio, based on the sum of squared singular values of $\bm{Q}$, reached 90% at 42 dimensions and 95% at 82 dimensions.
3.4 Obtaining the leaderboard scores
We obtained benchmark scores for the language models used in our experiments from Open LLM Leaderboard v1.666https://huggingface.co/spaces/open-llm-leaderboard-old/open_llm_leaderboard Between April 2023 and June 2024, this leaderboard evaluated language models on six tasks: AI2 Reasoning Challenge (ARC) Clark et al. (2018), HellaSwag Zellers et al. (2019), MMLU Hendrycks et al. (2021a), TruthfulQA Lin et al. (2022a), Winogrande Sakaguchi et al. (2019), and GSM8K Cobbe et al. (2021). Along with individual task scores, we also use the average score across all six tasks, referred to as 6-TaskMean.
3.5 Byte-normalized KL divergence for cross-experiment comparison
The KL divergence of text generation, $\textrm{KL}(p_{i},p_{j})$, generally increases with the number of tokens in the text. As a result, it cannot be directly compared with the KL divergence from other experiments using different text data. For models that use the same tokenizer, one can normalize the KL divergence by the average number of tokens in the text to obtain the KL divergence per token. However, when comparing KL divergence between models with different tokenizers, as in this study, it is more appropriate to normalize by the average text length in bytes and use the KL divergence per byte. For instance, if $\textrm{KL}(p_{i},p_{j})=1{,}000$, dividing by the average text length of 972.3188 bytes yields a KL divergence per byte of $1{,}000/972.3=1.028$ nats $=1.484$ bits [^7] To convert the code length unit from nats to bits, multiply by $1.44271/\log 2=$ 1.4427..
KL divergence can be understood from the perspective of coding theory as a measure of how much longer a message becomes when encoded using an incorrect probability distribution. $\textrm{KL}(p_{i},p_{j})$ represents the extra code length required when encoding text generated from probability distribution $p_{i}$ using a different distribution $p_{j}$. Dividing this value by the average text length in bytes gives the additional code length per byte. In the example above, this means that, due to differences between the models, an extra 1.484 bits are needed to encode each byte of text.
Figure 5: Model maps illustrating model performance. From left to right, the panels show each model’s mean log-likelihood, 6-TaskMean score, and the “primary task,” meaning the task for which each model achieves its highest standardized score among the six tasks (Appendix G.4). The color bar is clipped at the 10th percentile for mean log-likelihood and 6-TaskMean, with darker colors indicating better performance. In the primary task panel, models with standardized scores below zero on all six tasks are labeled “All Under 0.”
Figure 6: A model map color-coded by (Left) model size and (Right) model creation date.
4 Map of Language Models
We applied t-SNE van der Maaten and Hinton (2008) to the log-likelihood matrix $\bm{L}$ for dimensionality reduction [^8] The perplexity value was set to 30, and we used the scikit-learn implementation (Varoquaux et al., 2015).. Using this visualization, we analyze the insights gained from the model map in this section. While this paper presents model maps using $\bm{L}$, alternative maps using the double-centered log-likelihood matrix $\bm{Q}$, as well as a model map with labels for all language models, are available in Appendix H.
4.1 Visualizing attributes on the model map
Model type.
The top panel of Fig. 1 visualizes the distribution of model types, with each model color-coded according to its type. We observe that models belonging to the same type tend to cluster together, forming distinct regions on the map (e.g., llama-2, mistral, and gemma). In particular, models optimized for coding tasks [^9] For example, starcoder2-7b (Lozhkov et al., 2024), deepseek-coder-1.3b-base (Guo et al., 2024), codegemma-2b (CodeGemma Team et al., 2024) and CodeLlama-7b-Instruct-hf (Rozière et al., 2024). appear in a relatively compact region, suggesting that these models share notable similarities in their probability distributions.
Figure 7: (Left) Models tagged with “the Pile.” (Right) Difference between the standardized mean log-likelihood and the standardized 6-TaskMean score.
Text category.
The bottom panel of Fig. 1 shows the model map with each model color-coded according to the text category in which it achieves the highest standardized log-likelihood score (Appendix G.4). From this figure, we see that models exhibiting high likelihoods for the same text category are grouped together. Notably, the cluster containing coding-specialized models in the top panel aligns with the GitHub/StackExchange region in the bottom panel, suggesting that these models have relatively high likelihoods for text originating from GitHub and StackExchange.
Model performance.
Figure 5 visualizes two evaluation metrics: mean log-likelihood and benchmark task performance. From the left and central panels, we see that both metrics exhibit similar trends on the map, where models that lie close together tend to show similar metric values. Additionally, in the right panel, the GSM8K/MMLU region corresponds to the ArXiv/PubMed Central region in the bottom panel of Fig. 1, suggesting that models with high likelihoods on academic and scientific texts also tend to perform well on mathematical reasoning and academic knowledge-intensive tasks.
Model size and creation date.
Figure 6 shows the distribution of models by size and creation date. Compared to Fig. 5, newer models generally perform better, but model size does not always correlate with performance, as some smaller models perform comparably to larger ones.
4.2 Detection of data leakage
Since the text data we used was extracted from the Pile corpus, models that were pre-trained on the Pile are likely to exhibit higher log-likelihood values than their actual capabilities measured by benchmarks. We analyze this effect using the model map in Fig. 7. The left panel highlights models that used the Pile for pre-training. The right panel shows models with high mean log-likelihood relative to their 6-TaskMean score. The alignment between these two distributions suggests that models pre-trained on the Pile tend to achieve higher likelihoods on our text data, while their benchmark performance remains comparatively lower.
ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K 6-TaskMean mean log-likelihood
Pearson’s $r$ 0.972 0.948 0.963 0.958 0.965 0.935 0.976 0.994
Spearman’s $\rho$ 0.976 0.974 0.969 0.937 0.973 0.899 0.978 0.990
Table 2: Results of ridge regression for predicting benchmark scores from model coordinates. Predictions for 6-TaskMean and mean log-likelihood are also included. High correlation coefficients are observed across all settings.
5 Predicting Model Performance from Model Coordinates
As shown in the central panel of Fig. 5, the positioning of models on the map suggests that a model’s benchmark performance may be inferred from its coordinates. In this section, we conduct a regression analysis using the $\bm{q}$-coordinates to predict benchmark scores and evaluate predictive performance.
5.1 Benchmark scores and models
We use six benchmark scores from Open LLM Leaderboard v1, as described in Section 3.4. Our experiments are conducted on 996 models for which these benchmark scores are available101010 From the 1,018 models, we excluded four that lacked TruthfulQA scores and 18 from deepseek-ai, as their benchmark scores were incomplete due to the models being relatively new, leaving 996 for analysis. For simplicity, we denote the number of models as $K$. .
5.2 Setting for regression analysis
For each benchmark task, the dataset is given as $\{(\bm{q}_{1},f_{1}),\dots,(\bm{q}_{K},f_{K})\}$, where $\bm{q}_{i}\in\mathbb{R}^{N}$ is the double-centered log-likelihood vector of the language model $p_{i}$, and $f_{i}\in[0,100]$ is its corresponding benchmark score. We use ridge regression to predict each benchmark score. Let $\bm{Q}\in\mathbb{R}^{K\times N}$ be the matrix of explanatory variables, and let $\bm{f}=(f_{1},\ldots,f_{K})\in\mathbb{R}^{K}$ be the vector for the target variable. The objective function with parameter $\bm{w}\in\mathbb{R}^{N}$ is given by:
$L(\bm{w})=\|\bm{f}-\bm{Q}\bm{w}\|^{2}+\alpha\|\bm{w}\|^{2},$
where $\alpha\in\mathbb{R}_{>0}$ is a hyperparameter that controls the strength of regularization. Since the number of variables $N$ is much larger than the sample size $K$ ($N\gg K$), making this a high-dimensional regression setting, we carefully set $\alpha$ using cross-validation to avoid overfitting.
We split the set of models into five folds and perform parameter training [^11] We used scikit-learn Varoquaux et al. (2015) RidgeCV. and benchmark score prediction. To mitigate the effect of randomness, we repeat the data splitting with five different seeds and take the average of the predictions as the final predicted score. As evaluation metrics, we compute Pearson’s $r$ and Spearman’s $\rho$ to measure the correlation between the predicted and benchmark scores. Additionally, we conduct experiments by replacing the target variable with 6-TaskMean and mean log-likelihood, leading to a total of eight experimental settings. See Appendix K.1 for details.
5.3 Results
Table 2 presents the results of the regression analysis. For all benchmark tasks, the correlation coefficients are high, demonstrating that ridge regression achieves strong predictive performance. Similarly, for 6-TaskMean, the correlation remains high, as also illustrated by the scatter plot in Fig. 8. Even for mean log-likelihood, the regression model attains a high correlation, despite the explanatory variables being double-centered.
Figure 8: Scatter plot of predicted scores versus 6-TaskMean scores for test sets. Pearson’s $r$ is 0.976, and Spearman’s $\rho$ is 0.978. Each point is color-coded by the mean log-likelihood, and the color bar is clipped at the 10th percentile. Higher mean log-likelihood values generally correspond to higher 6-TaskMean scores. However, some models with extremely high mean log-likelihood due to data leakage (Section 4.2) appear in the lower score region, deviating from the general trend. Despite this, the regression-based predictions remain highly accurate, as reflected in the strong correlation coefficients. Scatter plots for individual benchmark tasks are provided in Fig. 18 in Appendix K.2.
6 Empirical Validation of Theory
Figure 9: Relationship between the squared Euclidean distance of model coordinates and KL divergence. (Left) In the token-level experiment (Section 6.1), each point represents a text. (Right) In the text-level experiment (Section 6.2), each point represents a pair of models.
Figure 10: Visualization of 36 language models obtained by linearly interpolating pretrained model weights. Each point is color-coded according to its mean log-likelihood. (Left) Models in the weight parameter space. (Right) Models in the probability distribution space, represented by the $\bm{q}$-coordinate system.
In Section 2, we discussed model coordinates for text probability models. Appendix E extends this framework to token-sequence probability models. These theoretical results state that the squared Euclidean distance in the model-coordinate space approximates the KL divergence between models. To validate this, we first evaluate token-level conditional probability models (Section 6.1), since token-level experiments are generally easier to conduct than text-level experiments. We then extend our analysis to text probability models (Section 6.2). Additionally, in Section 6.3, we explore the relationship between model weight parameters and model coordinates.
6.1 Validation of (31) for token-level models
Settings.
Two models with a shared tokenizer Llama-2-7b-hf and Llama-2-7b-chat-hf (Touvron et al., 2023b), denoted as $p_{1}$ and $p_{2}$, were used. Following the method described in Appendix E.1, we computed the model coordinates $\bm{\zeta}_{1}(x)$ and $\bm{\zeta}_{2}(x)$ for each text $x=(y_{1},\cdots,y_{n})\in D$, where these coordinates are centered vectors with elements $\log p_{i}(y_{t}|y^{t-1})$ as defined in (27). We then calculated the squared Euclidean distance between these coordinates, $\|\bm{\zeta}_{1}(x)-\bm{\zeta}_{2}(x)\|^{2}$. To obtain the exact KL divergence between models, we used the outputs of the softmax function in the language models. We computed the sum of per-token KL divergences: $\sum_{t=1}^{n}\mathrm{KL}(p_{1}(y_{t}|y^{t-1}),p_{2}(y_{t}|y^{t-1})),$ which is also used in Lv et al. (2023).
Results and discussion.
The left panel of Fig. 9 shows a scatter plot of the squared Euclidean distance and KL divergence for $x\in D$, with a Pearson’s correlation coefficient of $r=0.893$. This result indicates that (31) provides a good approximation in actual language models.
6.2 Validation of (3) for text-level models
Settings.
Among the 292 language models sharing the tokenizer with Llama-2-7b-hf (Touvron et al., 2023b), we excluded the five models with the largest values of $\sum_{x\in D}\|\bm{\zeta}_{i}(x)\|^{2}$, leaving 287 models for our experiment. Using the approach described in Section 2, we computed the model coordinates for each model and calculated the squared Euclidean distance $\|\bm{q}_{i}-\bm{q}_{j}\|^{2}$ between every pair of models. Because it is extremely difficult to directly compute $\mathrm{KL}(p_{i},p_{j})$, we instead used $\frac{1}{N}\sum_{x\in D}\bigl{\|}\bm{\zeta}_{i}(x)-\bm{\zeta}_{j}(x)\bigr{\|}^{2}$ as a proxy. For a theoretical justification that this quantity approximates $\mathrm{KL}(p_{i},p_{j})$, see (32) in Appendix E.
Results and discussion.
As shown in the right panel of Fig. 9, the scatter plot of squared Euclidean distance versus KL divergence exhibits a Pearson’s correlation coefficient of $r=0.904$. This finding confirms that the relationship in (3) holds approximately in practical language models.
6.3 Relationship between model weights and model coordinates
For language models with the same architecture, comparison can also be conducted via weight parameters. We investigated how the structure of the model-coordinate space aligns with the structure of the weight-parameter space. We generated 36 new language models by linearly interpolating the weights of Llama-2-7b-hf, Llama-2-7b-chat-hf (Touvron et al., 2023b), and vicuna-7b-v1.5 (Zheng et al., 2023). For these 36 models, we computed text-generation log-likelihoods and visualized the resulting model coordinates in two dimensions using Principal Component Analysis (PCA). Figure 10 shows these 36 models in both weight space and model coordinate space. We observe that the two-dimensional grid structure in the weight space is mapped continuously into the model coordinate space. This finding supports the validity of (6) and (14) in Appendix D, especially for models that are close to one another. Further details on this experiment can be found in Appendix J.
7 Conclusion
We propose a method to compare autoregressive language models using log-likelihoods from a predefined text set. By treating these as model coordinates, we show that the squared Euclidean distance approximates KL divergence, enabling efficient large-scale comparisons. Experiments with over 1,000 models confirm its effectiveness in structuring model relationships and predicting benchmark task performance, while also validating the theoretical foundations.
Limitations
  • Changing the text data will alter the analysis results of the model map. This is both a limitation and an advantage of the proposed method, because it allows us to choose text data according to the analysis objective. For example, if we want to investigate code-focused language models in more detail, we can increase the proportion of code data from GitHub, thereby increasing the resolution of code-focused models on the model map.
  • The proof (Section 2 and Appendix D) that the squared Euclidean distance in the model coordinate system approximates the KL divergence assumes that the language model’s text-generation probabilities closely match the distribution of the text data. When this assumption does not hold, the approximation accuracy decreases. However, even under such circumstances, the model coordinates should still function sufficiently as model features.
  • If the text data used for the model coordinates is contained in a language model’s pre-training corpus, that model’s mean log-likelihood may be overestimated. This data leakage, or data contamination, is generally non-negligible, as shown in Fig. 7 in Section 4.2, which illustrates the effect of using the Pile corpus. However, comparing it against benchmark scores makes it possible to detect such data leakage, and one can remove models that are affected. Furthermore, the model map based on the $\bm{q}$-coordinate system is robust to contamination in the data, as the estimation of KL divergence using the squared Euclidean distance in the $\bm{q}$-coordinate system remains valid even if the true generative model $p_{0}$ that produces the dataset $D$ varies, as long as all compared models remain sufficiently close to $p_{0}$. For instance, even if $D$ is generated from one of the $K$ models, such as $p_{i}$, the estimation formula (3) remains correct (Appendix D.7).
  • Beyond the data leakage mentioned above, other systematic errors introduced into the model coordinates can also affect the model map. As noted in Appendix C, $\bm{\ell}_{i}$ and ¯$\bar{\ell}_{i}$ are susceptible to systematic biases. However, thanks to double centering, $\bm{q}_{i}$ is less influenced by bias terms.
  • Although the calculation of model coordinates is linear time $O(KN)$ in the number of models $K$ and the number of texts $N$, it still requires a non-negligible amount of computation. In our experiments, it took about 10 minutes on a single GPU (RTX 6000 Ada) to compute coordinates ($N=10^{4}$, float16) for a single 7B model.
  • Computing the model map visualization from the model coordinates is generally not linear in $K$. For example, t-SNE requires a distance matrix, incurring $O(K^{2}N)$ computational cost. However, in modern computing environments, as long as $K$ is not extremely large (e.g., in the millions), the cost of visualization is negligible compared to the cost of calculating the model coordinates.
  • A sufficiently large number of text samples, $N$, is desirable for the model coordinates. We used $N=10^{4}$. Since the error in the KL divergence estimate due to randomness decreases proportionally to $N^{-1/2}$, $N$ must be increased according to the desired resolution of the model map.
  • When using model coordinates as feature vectors, $N=10^{4}$ can be unwieldy. According to the experiment in Section 3.3, applying PCA to reduce the dimensionality of $\bm{q}$-coordinates to 82 dimensions still retains 95% of the information. However, the predictive performance of such dimension-reduced features has not yet been tested.
  • Since the language models used in our experiments were obtained from the Open LLM Leaderboard v1 (which ran from April 2023 to June 2024), our discussion of models released after June 2024 is limited.
  • The results of the task-performance prediction in Section 5 should be interpreted conservatively. While we use a proper cross-validation setup to prevent data leakage (splitting the training set, validation set, and test set), if very similar models appear in both training and test sets (for example, models that share a base model and differ only slightly by fine-tuning), it might become easier to predict task performance. Also, we have not evaluated models that are not listed on the leaderboard.
  • The theoretical validation experiment in Section 6 is limited. Currently, it is difficult to directly compute exact KL divergence values among text-generation probability models, so conducting more precise validation experiments remains a future challenge.
  • In the method of computing model coordinates from token-sequence conditional probabilities (Appendix E and Appendix F), the proof that the squared Euclidean distance in the model coordinate system approximates the KL divergence requires additional assumptions (Appendix F.3). In practice, Assumption 2 does not hold, and due to the variations in (38), $\|\bm{\zeta}_{i}-\bm{\zeta}_{j}\|^{2}$ in (31) tends to overestimate the KL divergence. Nonetheless, even in such situations, token-level model coordinates will still likely function sufficiently as features.
Acknowledgments
This study was partially supported by JSPS KAKENHI 22H05106, 23H03355, JST CREST JPMJCR21N3, JST SPRING JPMJSP2110, JST BOOST JPMJBS2407.
Appendix A Related Work
In recent years, research on comparing large language models (LLMs) has gained attention. This section provides an overview of existing studies from three perspectives: model parameters, activations [^12] In general, “activations” refer to the intermediate outputs of Transformer models (e.g., the residual stream or neurons) Bereska and Gavves (2024)., and probability distributions.
Comparison of model parameters.
One approach to comparing LLMs is to analyze their parameters. Zhu et al. (2025) proposed a statistical framework for evaluating parameter similarity between different models and introduced a method for determining whether these models were trained independently. Additionally, Yadav et al. (2023b) focused on parameter changes due to task adaptation, specifically analyzing task vectors [^13] The difference in parameters before and after fine-tuning is referred to as a task vector, and arithmetic operations on these vectors are effective Ilharco et al. (2023).. They proposed a method to mitigate interference when integrating task vectors from different models. Specifically, by reducing redundant numerical components and adjusting for conflicting signs, their approach enables effective model merging.
Comparison of activations.
Comparisons of LLMs based on activations have also been studied. Zhou et al. (2025) quantified the similarity between LLMs by measuring the cosine similarity of activation differences for linguistic minimal pairs. In particular, they used datasets such as BLiMP Warstadt et al. (2020) and showed that model similarity is significantly influenced by the pre-training dataset.
Comparison of probability distributions.
Several approaches compare LLMs using probability distributions. Lv et al. (2023) proposed a method for computing coefficients in parameter ensembling by providing the same input text to two models and comparing the softmax probability distributions at each token. Specifically, they used KL divergence and summed the results to derive appropriate coefficients. Furthermore, Yax et al. (2025) proposed a similarity metric based on the conditional probabilities of LLMs and introduced a method for calculating the phylogenetic distance between different models. Additionally, a method has been proposed for measuring differences in conditional probabilities based on prompts using KL divergence Melamed et al. (2024).
Appendix B Double Centering
We confirm the notation and computational operations. The matrices $\bm{L}$, $\bm{\Xi}$, and $\bm{Q}$ are all of size $K\times N$, and their elements are denoted by $\ell_{is},\xi_{is},q_{is}$, respectively. In particular, we have $\ell_{is}=\ell_{i}(x_{s})$. First, row-wise centering of $\bm{L}$ is performed by subtracting the mean log-likelihood ¯$\bar{\ell}_{i}$ of each model from each row $(\ell_{i1},\ldots,\ell_{iN})$, resulting in $\bm{\Xi}=(\bm{\xi}_{1},\ldots,\bm{\xi}_{K})^{\top}$. Next, column-wise centering of $\bm{\Xi}$ is performed by subtracting the coordinate component $\bar{\xi}_{s}$ of the mean vector $\bar{\bm{\xi}}$ from each column $(\xi_{1s},\ldots,\xi_{Ks})^{\top}$, yielding $\bm{Q}=(\bm{q}_{1},\ldots,\bm{q}_{K})^{\top}$. Thus, this process involves double centering, where column-wise centering follows row-wise centering. Notably, even after column-wise centering, the row-wise mean of $\bm{Q}$ remains zero:
$\frac{1}{N}\sum_{s=1}^{N}q_{is}$ $=\frac{1}{N}\sum_{s=1}^{N}(\xi_{is}-\bar{\xi}_{s})$
$=\frac{1}{N}\sum_{s=1}^{N}\xi_{is}-\frac{1}{NK}\sum_{i=1}^{K}\sum_{s=1}^{N}\xi_{is}$
$=0-0=0.$ (5)
The column-wise centering can be interpreted as follows. In the $\bm{\xi}$-coordinate system, the mean vector $\bar{\bm{\xi}}$ of the $K$ model coordinates $\bm{\xi}_{1},\ldots,\bm{\xi}_{K}$ can be regarded as representing an “average model.” By redefining this average model as the new origin, we obtain the $\bm{q}$-coordinate system. From the definition $\bm{q}_{i}=\bm{\xi}_{i}-\bar{\bm{\xi}}$, its mean satisfies
$\sum_{i=1}^{K}\bm{q}_{i}/K=\bm{0}.$
The row-wise centering can be interpreted as follows. Let $\bm{1}_{N}=(1,\ldots,1)^{\top}\in\mathbb{R}^{N}$. Since ¯$\bar{\ell}_{i}=\bm{1}_{N}^{\top}\bm{\ell}_{i}/N$, we have $\bm{\xi}_{i}=\bm{\ell}_{i}-\bar{\ell}_{i}\bm{1}_{N}$. Thus,
$\bm{1}_{N}^{\top}\bm{\xi}_{i}=\bm{1}_{N}^{\top}\bm{\ell}_{i}-\bar{\ell}_{i}N=0,$
and furthermore, $\bm{1}_{N}^{\top}\bar{\bm{\xi}}=0$. From this, equation (5) is actually trivial, as
$\bm{1}_{N}^{\top}\bm{q}_{i}=\bm{1}^{\top}(\bm{\xi}_{i}-\bar{\bm{\xi}})=0-0=0.$
The row-wise centering implies that $\bm{\xi}_{1},\ldots,\bm{\xi}_{K}$ and $\bm{q}_{1},\ldots,\bm{q}_{K}$ lie in the subspace orthogonal to $\bm{1}_{N}$.
Appendix C Effect of Errors in Model Coordinates
We analyze the impact of additive errors $\epsilon_{is}$ in the log-likelihood vector components $\ell_{is}$ for $i=1,\ldots,K$ and $s=1,\ldots,N$. Denoting the true values with an asterisk as $\ell_{is}^{*}$, the observed values can be expressed as
$\ell_{is}=\ell_{is}^{*}+\epsilon_{is}.$
We decompose the error as follows:
$\epsilon_{is}=a+b_{i}+c_{s}+d_{is},$
where we assume, without loss of generality, the constraints
$\sum_{i=1}^{K}b_{i}=\sum_{s=1}^{N}c_{s}=\sum_{i=1}^{K}d_{is}=\sum_{s=1}^{N}d_{is}=0.$
Here, $a$, $b_{i}$, and $c_{s}$ are bias terms, while $d_{is}$ represents interaction terms. Using simple calculations, we obtain:
¯$\bar{\ell}_{i}=\frac{1}{N}\sum_{s=1}^{N}\ell_{is}=\bar{\ell}_{i}^{*}+a+b_{i},$
$\xi_{is}=\ell_{is}-\bar{\ell}_{i}=\xi_{is}^{*}+c_{s}+d_{is},$
$\bar{\xi}_{s}=\frac{1}{K}\sum_{i=1}^{K}\xi_{is}=\bar{\xi}_{s}^{*}+c_{s},$
$q_{is}=\xi_{is}-\bar{\xi}_{s}=q_{is}^{*}+d_{is}.$
Additionally, for the differences between two models, which are crucial for the model map, we obtain:
$\ell_{is}-\ell_{js}=\ell_{is}^{*}-\ell_{js}^{*}+b_{i}-b_{j}+d_{is}-d_{js},$
$\xi_{is}-\xi_{js}=\xi_{is}^{*}-\xi_{js}^{*}+d_{is}-d_{js},$
$q_{is}-q_{js}=q_{is}^{*}-q_{js}^{*}+d_{is}-d_{js}.$
Thus, the terms affected by the error are ¯$\bar{\ell}_{i}$, which is influenced by $a+b_{i}$, and $\ell_{is}-\ell_{js}$, which is affected by $b_{i}+d_{is}$, meaning it is influenced by the bias terms. However, in the centered values $\xi_{is}-\xi_{js}$ and $q_{is}-q_{js}$, only the interaction term $d_{is}$ contributes to the error.
Appendix D Theory of Model Coordinates for Text Probability Distributions
In this section, we prove the main results of Section 2, namely (2) and (3). Our discussion applies not only to text probability distributions but also more generally to any setting where i.i.d. observations $x_{1},\dots,x_{N}\sim p_{i}(x)$ are available. Compared to the previous study that proposed model maps (Shimodaira, 1993; Shimodaira and Cao, 1998), we conduct a more precise analysis in this paper. Specifically, while the previous study provides only a brief evaluation of the approximation, we present a more transparent discussion based on the properties of the exponential family of distributions. In the next section, as the starting point of our discussion, we construct a super model that includes the $K$ models $p_{i}$, $i=1,\dots,K$, as submodels. This model is introduced as a mathematical tool to rigorously prove the main theorem of this paper, and we do not compute it numerically in practice.
D.1 Exponential family of distributions
We first consider a model in the exponential family of distributions parameterized by a $K$-dimensional parameter $\bm{\theta}\in\mathbb{R}^{K}$:
$p(x;\bm{\theta})=p_{0}(x)\exp(\bm{\theta}^{\top}\bm{b}(x)-\psi(\bm{\theta})).$ (6)
Here, the function $\bm{b}(x)=(b_{1}(x),\dots,b_{K}(x))^{\top}$ will be defined later using the $K$ models. The normalization constant is given by
$Z(\bm{\theta})$ $=\sum_{x\in\mathcal{X}}p_{0}(x)\exp(\bm{\theta}^{\top}\bm{b}(x)),$
$\psi(\bm{\theta})$ $=\log Z(\bm{\theta}),$
[…]
$p(x;\bm{0})=p_{0}(x).$
To associate the $K$ models with (6), we define $\bm{b}(x)$. For a constant $\lambda>0$, we set
$\lambda b_{i}(x):=\ell_{i}(x)-\ell_{0}(x).$ (7)
The constant $\lambda$ is an order parameter introduced for theoretical convenience, and in our theoretical framework, we assume that $b_{i}(x)$ is of constant order and that $\lambda$ is sufficiently small [^14] If numerical computation were to be performed, $\lambda$ could be set to any arbitrary value (e.g., $\lambda=1$).. Thus, we essentially assume that $|\ell_{i}(x)-\ell_{0}(x)|=O_{p}(\lambda)$ is sufficiently small, implying that each model $p_{i}$ provides a good approximation of the true generative model $p_{0}$. In the proof of the main theorem, we consider the asymptotic theory as $\lambda\to 0$, retaining terms up to $O(\lambda^{2})$ while ignoring those of $O(\lambda^{3})$.
A one-hot vector is defined as $\bm{e}_{i}=(0,\ldots,0,1,0,\ldots,0)^{\top}\in\mathbb{R}^{K}$ for $i=1,\ldots,K$, where only the $i$-th element is 1. Then, setting $\bm{\theta}=\lambda\bm{e}_{i}$ gives
$p(x;\lambda\bm{e}_{i})=p_{i}(x).$ (8)
Indeed, substituting (7) into (6) yields
$p(x;\lambda\bm{e}_{i})$
$=$ $p_{0}(x)\exp(\lambda\bm{e}_{i}^{\top}\bm{b}(x)-\psi(\lambda\bm{e}_{i}))$
$=$ $p_{0}(x)\exp(\ell_{i}(x)-\ell_{0}(x)-\psi(\lambda\bm{e}_{i}))$
$=$ $p_{0}(x)(p_{i}(x)/p_{0}(x))\exp(-\psi(\lambda\bm{e}_{i}))$
$=$ $p_{i}(x)\exp(-\psi(\lambda\bm{e}_{i}))$
$=$ $p_{i}(x),$
where $\psi(\lambda\bm{e}_{i})=0$.
D.2 Properties of the exponential family of distributions
This section outlines some well-known basic properties of the exponential family of distributions, which have been established in the literature (Barndorff-Nielsen, 2014; Efron, 1978, 2022; Amari, 1982). We define the expectation and covariance matrix of $\bm{b}(x)$ as follows:
$\bm{\eta}(\bm{\theta}):=$ $\operatorname*{\mathbb{E}}_{x\sim p(\bm{\theta})}\bigl{(}\bm{b}(x)\bigr{)}=\sum_{x\in\mathcal{X}}\bm{b}(x)p(x;\bm{\theta}),$
$G(\bm{\theta}):=$ $\operatorname*{\mathbb{E}}_{x\sim p(\bm{\theta})}\bigl{\{}(\bm{b}(x)-\bm{\eta}(\bm{\theta}))(\bm{b}(x)-\bm{\eta}(\bm{\theta}))^{\top}\bigr{\}}$
$=$ $\operatorname*{Var}_{x\sim p(\bm{\theta})}\bigl{(}\bm{b}(x)\bigr{)}.$
Here, the elements of $\bm{\eta}(\bm{\theta})$ are expectations given by $\operatorname*{\mathbb{E}}_{x\sim p(\bm{\theta})}(b_{i}(x))$, and the elements of $G(\bm{\theta})$ are covariances given by $G_{ij}(\bm{\theta})=\operatorname*{Cov}_{x\sim p(\bm{\theta})}(b_{i}(x),b_{j}(x))$. These quantities can be expressed in terms of $\psi(\bm{\theta})$ as follows:
$\bm{\eta}(\bm{\theta})=\frac{\partial\psi(\bm{\theta})}{\partial\bm{\theta}},$ (9)
$G(\bm{\theta})=\frac{\partial^{2}\psi(\bm{\theta})}{\partial\bm{\theta}\partial\bm{\theta}^{\top}}.$ (10)
We now derive these two equations. First, from
$\frac{\partial Z(\bm{\theta})}{\partial\bm{\theta}}=\sum_{x\in\mathcal{X}}\bm{b}(x)p_{0}(x)e^{\bm{\theta}^{\top}\bm{b}(x)},$
we obtain
$\frac{\partial\psi(\bm{\theta})}{\partial\bm{\theta}}=$ $\frac{\partial\log Z}{\partial\bm{\theta}}=\frac{1}{Z(\bm{\theta})}\frac{\partial Z}{\partial\bm{\theta}}$
$=$ $\frac{1}{Z(\bm{\theta})}\sum_{x\in\mathcal{X}}\bm{b}(x)p_{0}(x)e^{\bm{\theta}^{\top}\bm{b}(x)}$
$=$ $\sum_{x\in\mathcal{X}}\bm{b}(x)p(x;\bm{\theta})=\bm{\eta}(\bm{\theta}).$
Thus, we have established (9). Next, using
$\frac{\partial p(x;\bm{\theta})}{\partial\bm{\theta}}=$ $\Bigl{(}\bm{b}(x)-\frac{\partial\psi}{\partial\bm{\theta}}\Bigr{)}p(x;\bm{\theta})$
$=$ $(\bm{b}(x)-\bm{\eta}(\bm{\theta}))p(x;\bm{\theta}),$
we obtain
$\frac{\partial^{2}\psi(\bm{\theta})}{\partial\bm{\theta}\partial\bm{\theta}^{\top}}=\frac{\partial\bm{\eta}(\bm{\theta})^{\top}}{\partial\bm{\theta}}$
$=$ $\frac{\partial}{\partial\bm{\theta}}\sum_{x\in\mathcal{X}}\bm{b}(x)^{\top}p(x;\bm{\theta})$
$=$ $\sum_{x\in\mathcal{X}}(\bm{b}(x)-\bm{\eta}(\bm{\theta}))\bm{b}(x)^{\top}p(x;\bm{\theta})$
$=$ $\sum_{x\in\mathcal{X}}(\bm{b}(x)-\bm{\eta}(\bm{\theta}))(\bm{b}(x)-\bm{\eta}(\bm{\theta}))^{\top}p(x;\bm{\theta})$
$=$ $G(\bm{\theta}).$
Thus, we have established (10).
D.3 Approximation of the KL divergence
The Kullback-Leibler (KL) divergence between the models $p(\bm{\theta})$ and $p(\bm{\theta}^{\prime})$ at parameter values $\bm{\theta},\bm{\theta}^{\prime}\in\mathbb{R}^{K}$ is given by
$\mathrm{KL}(p(\bm{\theta}),p(\bm{\theta}^{\prime}))$
$=$ $\sum_{x\in\mathcal{X}}p(x;\bm{\theta})\log\frac{p(x;\bm{\theta})}{p(x;\bm{\theta}^{\prime})}$
$=$ $\sum_{x\in\mathcal{X}}p(x;\bm{\theta})\bigl{\{}(\bm{\theta}-\bm{\theta}^{\prime})^{\top}\bm{b}(x)-\psi(\bm{\theta})+\psi(\bm{\theta}^{\prime})\bigr{\}}$
$=$ $(\bm{\theta}-\bm{\theta}^{\prime})^{\top}\bm{\eta}(\bm{\theta})-\psi(\bm{\theta})+\psi(\bm{\theta}^{\prime}).$ (11)
Here, we assume that the parameter values $\bm{\theta}$ and $\bm{\theta}^{\prime}$ are sufficiently close to $\bm{0}$. In particular, we assume $\|\bm{\theta}\|=O(\lambda)$ and $\|\bm{\theta}^{\prime}\|=O(\lambda)$. Substituting (9) and (10) into the Taylor expansion of $\psi(\bm{\theta})$ gives
$\psi(\bm{\theta}^{\prime})$
$=$ $\psi(\bm{\theta})+\frac{\partial\psi}{\partial\bm{\theta}^{\top}}(\bm{\theta}^{\prime}-\bm{\theta})$
$+\frac{1}{2}(\bm{\theta}^{\prime}-\bm{\theta})^{\top}\frac{\partial^{2}\psi(\bm{\theta})}{\partial\bm{\theta}\partial\bm{\theta}^{\top}}(\bm{\theta}^{\prime}-\bm{\theta})$
$+O(\|\bm{\theta}^{\prime}-\bm{\theta}\|^{3})$
$=$ $\psi(\bm{\theta})+\bm{\eta}(\bm{\theta})^{\top}(\bm{\theta}^{\prime}-\bm{\theta})$
$+\frac{1}{2}(\bm{\theta}^{\prime}-\bm{\theta})^{\top}G(\bm{\theta})(\bm{\theta}^{\prime}-\bm{\theta})+O(\lambda^{3}).$ (12)
Substituting (12) into (11) gives
$\mathrm{KL}(p(\bm{\theta}),p(\bm{\theta}^{\prime}))$
$=\frac{1}{2}(\bm{\theta}^{\prime}-\bm{\theta})^{\top}G(\bm{\theta})(\bm{\theta}^{\prime}-\bm{\theta})+O(\lambda^{3}).$ (13)
This corresponds to eq. (9) of Oyama et al. (2023). Here, the equation holds approximately by ignoring higher-order terms of $O(\lambda^{3})$. For more details, refer to Amari (1982, p. 369) and Efron (2022, p. 35). More generally, $G(\bm{\theta})$ represents the Fisher information metric, and (13) holds for a wide class of probability models Amari (1998) with $\|\bm{\theta}^{\prime}-\bm{\theta}\|=O(\lambda)$. Furthermore, since $G(\bm{\theta})=G(\bm{0})+O(\|\bm{\theta}\|)=G(\bm{0})+O(\lambda)$, we obtain
$\mathrm{KL}(p(\bm{\theta}),p(\bm{\theta}^{\prime}))$
$=\frac{1}{2}(\bm{\theta}^{\prime}-\bm{\theta})^{\top}G(\bm{0})(\bm{\theta}^{\prime}-\bm{\theta})+O(\lambda^{3}).$ (14)
D.4 The variance representation of the KL divergence
Substituting $p_{i}=p(\lambda\bm{e}_{i})$ into (14) gives
$2\mathrm{KL}(p_{i},p_{j})=2\mathrm{KL}(p(\lambda\bm{e}_{i}),p(\lambda\bm{e}_{j}))$
$=\lambda^{2}(\bm{e}_{i}-\bm{e}_{j})^{\top}G(\bm{0})(\bm{e}_{i}-\bm{e}_{j})+O(\lambda^{3}).$ (15)
Here, we have
$\bm{e}_{i}^{\top}G(\bm{0})\bm{e}_{j}=G_{ij}(\bm{0})$
$=\operatorname*{\mathbb{E}}_{x\sim p_{0}}\Bigl{\{}(b_{i}(x)-\eta_{i}(\bm{0}))(b_{j}(x)-\eta_{j}(\bm{0}))\Bigr{\}}.$ (16)
Next, we substitute (16) and (7) into the right-hand side of (15) to derive an alternative expression for the KL divergence:
$\lambda^{2}(\bm{e}_{i}-\bm{e}_{j})^{\top}G(\bm{0})(\bm{e}_{i}-\bm{e}_{j})$
$=$ $\lambda^{2}\operatorname*{\mathbb{E}}_{x\sim p_{0}}\Bigl{[}\Bigl{\{}(b_{i}(x)-\eta_{i}(\bm{0}))$
$\qquad-(b_{j}(x)-\eta_{j}(\bm{0}))\Bigr{\}}^{2}\Bigr{]}$
$=$ $\operatorname*{\mathbb{E}}_{x\sim p_{0}}\Bigl{[}\Bigl{\{}(\ell_{i}(x)-\ell_{0}(x))-(\ell_{j}(x)-\ell_{0}(x))$
$\qquad-\operatorname*{\mathbb{E}}_{x^{\prime}\sim p_{0}}\Bigl{(}(\ell_{i}(x^{\prime})-\ell_{0}(x^{\prime}))$
$\qquad\qquad-(\ell_{j}(x^{\prime})-\ell_{0}(x^{\prime}))\Bigr{)}\Bigr{\}}^{2}\Bigr{]}$
$=$ $\operatorname*{\mathbb{E}}_{x\sim p_{0}}\Bigl{[}\Bigl{\{}\ell_{i}(x)-\ell_{j}(x)$
$\qquad-\operatorname*{\mathbb{E}}_{x^{\prime}\sim p_{0}}\Bigl{(}\ell_{i}(x^{\prime})-\ell_{j}(x^{\prime})\Bigr{)}\Bigr{\}}^{2}\Bigr{]}$
$=$ $\operatorname*{Var}_{x\sim p_{0}}\Bigl{(}\ell_{i}(x)-\ell_{j}(x)\Bigr{)}.$
Finally, substituting this result into (15) yields
$2\mathrm{KL}(p_{i},p_{j})=\operatorname*{Var}_{x\sim p_{0}}\Bigl{(}\ell_{i}(x)-\ell_{j}(x)\Bigr{)}+O(\lambda^{3}).$ (17)
This establishes (2). Furthermore, since $|\ell_{i}(x)-\ell_{j}(x)|=O_{p}(\lambda)$, the magnitude of (17) is $O(\lambda^{2})$.
D.5 Estimation of the KL divergence
[…]
$\operatorname*{\mathbb{E}}_{x\sim D}(f(x))=\operatorname*{\mathbb{E}}_{x\sim p_{0}}(f(x))+O_{p}(N^{-1/2}).$
Applying this to (17) and ignoring the terms of order $O_{p}(\lambda^{3}+\lambda^{2}N^{-1/2})$, we obtain the following approximation:
$2\mathrm{KL}(p_{i},p_{j})\approx\operatorname*{Var}_{x\sim D}\Bigl{(}\ell_{i}(x)-\ell_{j}(x)\Bigr{)}.$ (18)
We define the coordinates $\bm{\xi}_{i}\in\mathbb{R}^{N}$ of model $p_{i}$ as
$\bm{\xi}_{i}=(\xi_{i1},\ldots,\xi_{iN})^{\top}$
with
$\xi_{is}:=\ell_{i}(x_{s})-\operatorname*{\mathbb{E}}_{x\sim D}(\ell_{i}(x))$
for $s=1,\ldots,N$. From (18), we obtain
$2\mathrm{KL}(p_{i},p_{j})\approx$ $\frac{1}{N}\sum_{s=1}^{N}(\xi_{is}-\xi_{js})^{2}$
$=$ $\frac{1}{N}\|\bm{\xi}_{i}-\bm{\xi}_{j}\|^{2}.$ (19)
Since
$\|\bm{\xi}_{i}-\bm{\xi}_{j}\|^{2}=\|(\bm{q}_{i}+\bar{\bm{\xi}})-(\bm{q}_{j}+\bar{\bm{\xi}})\|^{2}=\|\bm{q}_{i}-\bm{q}_{j}\|^{2},$
this establishes (3).
D.6 Relationships among the three types of model coordinates
Let $\bm{1}_{N}=(1,\ldots,1)^{\top}\in\mathbb{R}^{N}$. From the definitions of the $\bm{\xi}$-coordinate system and the $\bm{q}$-coordinate system, we have
$\bm{\xi}_{i}$ $=\bm{q}_{i}+\bar{\bm{\xi}},$
$\bm{\ell}_{i}$ $=\bm{\xi}_{i}+\bar{\ell}_{i}\bm{1}_{N}$
$=\bm{q}_{i}+\bar{\ell}_{i}\bm{1}_{N}+\bar{\bm{\xi}}.$ (20)
Additionally, equation (5) in Appendix B can be rewritten as
$\bm{1}_{N}^{\top}\bm{q}_{i}=0.$ (21)
Thus, we obtain
$\|\bm{\ell}_{i}-\bm{\ell}_{j}\|^{2}$
$=$ $\|(\bm{q}_{i}-\bm{q}_{j})+(\bar{\ell}_{i}-\bar{\ell}_{j})\bm{1}_{N}\|^{2}$
$=$ $\|\bm{q}_{i}-\bm{q}_{j}\|^{2}+N(\bar{\ell}_{i}-\bar{\ell}_{j})^{2}$
$\quad+2(\bar{\ell}_{i}-\bar{\ell}_{j})\bm{1}_{N}^{\top}(\bm{q}_{i}-\bm{q}_{j})$
$=$ $\|\bm{q}_{i}-\bm{q}_{j}\|^{2}+N(\bar{\ell}_{i}-\bar{\ell}_{j})^{2},$
where (20) and (21) are used in the first and last equations, respectively. This establishes (4).
Moreover, since ¯$\bar{\ell}_{i}=\bm{1}_{N}^{\top}\bm{\ell}_{i}/N$, it is straightforward that the component of the $\bm{\ell}$-coordinate system in the $\bm{1}_{N}$ direction is given by
$(\bm{1}_{N}/\sqrt{N})^{\top}\bm{\ell}_{i}=\sqrt{N}\bar{\ell}_{i}.$
D.7 Additional notes on modifying the underlying generative model
We examine the effect of changing the true generative model $p_{0}=p(\bm{0})$ that produces the data $D$. For clarity, we continue to use the same $p_{0}$ as before in constructing $p(x;\bm{\theta})$ as described in Section D.1. We then introduce a new parameter value $\bm{\theta}^{*}$. We assume that each element $x_{s}$ of the dataset $D$ is generated independently from
$x_{1},\ldots,x_{N}\sim p(x;\bm{\theta}^{*}),$ (22)
and that $\|\bm{\theta}^{*}\|=O(\lambda)$. In other words, the true generative model remains sufficiently close to the $K$ models, satisfying $\|\bm{\theta}_{i}-\bm{\theta}^{*}\|=O(\lambda)$ for $i=1,\ldots,K$.
First, consider replacing $G(\bm{\theta})$ in (13) of Section D.3 with $G(\bm{\theta}^{*})$. Because $G(\bm{\theta}^{*})=G(\bm{\theta})+O(\lambda)$, we obtain
$\mathrm{KL}(p(\bm{\theta}),p(\bm{\theta}^{\prime}))$
$=\frac{1}{2}(\bm{\theta}^{\prime}-\bm{\theta})^{\top}G(\bm{\theta}^{*})(\bm{\theta}^{\prime}-\bm{\theta})+O(\lambda^{3}).$ (23)
We are generalizing the discussion in the previous sections, and indeed, if we set $\bm{\theta}^{*}=\bm{0}$ in (23), then (14) is recovered.
Now substitute $p_{i}=p(\lambda\bm{e}_{i})$ into (23), yielding
$2\,\mathrm{KL}(p_{i},p_{j})=2\,\mathrm{KL}\bigl{(}p(\lambda\bm{e}_{i}),p(\lambda\bm{e}_{j})\bigr{)}$
$=\lambda^{2}(\bm{e}_{i}-\bm{e}_{j})^{\top}G(\bm{\theta}^{*})(\bm{e}_{i}-\bm{e}_{j})+O(\lambda^{3}),$ (24)
which is a generalization of (15) in Section D.4. Using the definition of $\bm{G}(\bm{\theta})$, we have
$\bm{e}_{i}^{\top}G(\bm{\theta}^{*})\bm{e}_{j}=G_{ij}(\bm{\theta}^{*})=\operatorname*{Cov}_{x\sim p(\bm{\theta}^{*})}\bigl{(}b_{i}(x),b_{j}(x)\bigr{)}.$ (25)
Substituting (25) and (7) into the right-hand side of (24) yields
$\lambda^{2}(\bm{e}_{i}-\bm{e}_{j})^{\top}G(\bm{\theta}^{*})(\bm{e}_{i}-\bm{e}_{j})$
$=$ $\,\lambda^{2}\Bigl{\{}\operatorname*{Var}_{x\sim p(\bm{\theta}^{*})}\bigl{(}b_{i}(x)\bigr{)}+\operatorname*{Var}_{x\sim p(\bm{\theta}^{*})}\bigl{(}b_{j}(x)\bigr{)}$
$\qquad-2\,\operatorname*{Cov}_{x\sim p(\bm{\theta}^{*})}\bigl{(}b_{i}(x),b_{j}(x)\bigr{)}\Bigr{\}}$
$=$ $\,\lambda^{2}\operatorname*{Var}_{x\sim p(\bm{\theta}^{*})}\bigl{(}b_{i}(x)-b_{j}(x)\bigr{)}$
$=$ $\,\operatorname*{Var}_{x\sim p(\bm{\theta}^{*})}\bigl{(}\ell_{i}(x)-\ell_{j}(x)\bigr{)}.$
Finally, substituting this back into (24) gives
$2\,\mathrm{KL}(p_{i},p_{j})=\operatorname*{Var}_{x\sim p(\bm{\theta}^{*})}\bigl{(}\ell_{i}(x)-\ell_{j}(x)\bigr{)}+O(\lambda^{3}),$ (26)
which is a generalization of (2).
Hence, the estimators for KL divergence in Section D.5, specifically (18) and (19), and also (3) in Section 2, remain valid even when the texts are generated by (22). Since (26) holds for any $\bm{\theta}^{*}$ with $\|\bm{\theta}^{*}\|=O(\lambda)$, this result also applies when the true data-generating model is $p(\bm{\theta}^{*})=p_{0}$, or, for instance, one of the models $p(\bm{\theta}^{*})=p_{i}$, or a mixture of the $K$ models, $\bm{\theta}^{*}=\sum_{i=1}^{K}\alpha_{i}\lambda\bm{e}_{i}$ with $\alpha_{i}=O(1)$. Therefore, this method is robust to contamination in the dataset $D$ (e.g., when the text corpus used for pre-training a model is included in $D$), as the estimation of KL divergence via the squared Euclidean distance in the $\bm{\xi}$-coordinate system or the $\bm{q}$-coordinate system remains relatively unaffected.
Appendix E Mapping Language Models into the Space of Token Probability Distributions
In Section 2, we discussed model maps based on the probability distributions $p_{i}(x)$ of texts generated by language models. This approach requires computing probabilities for a large number of texts in the dataset $D=(x_{1},\ldots,x_{N})$, leading to high computational costs. To mitigate this issue, we focus on the fact that a text $x=(y_{1},\ldots,y_{n})$ is a sequence of tokens. Instead of using text probabilities, we discuss model maps based on the conditional probability distributions of token generation, $p_{i}(y_{t}|y^{t-1})$. In this approach, model coordinates are computed using only a single text $x$. A limitation of this approach is that it can only be used for comparing models that share the same tokenizer. Furthermore, the current estimation method ignores the variance of the expected log-likelihood ratio of conditional probabilities, resulting in a rough approximation. Thus, the estimated values should be regarded only as reference values rather than precise measurements.
E.1 Model coordinates
For a text $x=(y_{1},\ldots,y_{n})$, the coordinates of model $p_{i}$
$\bm{\zeta}_{i}=(\zeta_{i1},\ldots,\zeta_{in})^{\top}\in\mathbb{R}^{n}$
are defined as
$\zeta_{it}:=\log p_{i}(y_{t}|y^{t-1})-\ell_{i}(x)/n$ (27)
for $t=1,\ldots,n$. This is centered for each $i$ and for each text, satisfying $\sum_{t=1}^{n}\zeta_{it}=0$.
E.2 Kullback-Leibler divergence
The KL divergence for next-token generation in language models, where $y_{t}\sim p_{i}(y_{t}|y^{t-1})$, is given by
$\mathrm{KL}(p_{i}(y_{t}|y^{t-1}),p_{j}(y_{t}|y^{t-1}))=$ (28)
$\qquad\sum_{y_{t}\in\mathcal{V}}p_{i}(y_{t}|y^{t-1})\log\frac{p_{i}(y_{t}|y^{t-1})}{p_{j}(y_{t}|y^{t-1})}.$ (29)
We apply the results for text probability distributions from Section 2 and Appendix D to the conditional probability distributions of token generation. The equation corresponding to (2) is
$2\mathrm{KL}(p_{i}(y_{t}|y^{t-1}),p_{j}(y_{t}|y^{t-1}))\approx$
$\operatorname*{Var}_{y_{t}\sim p_{0}(y_{t}|y^{t-1})}\biggl{\{}\log\frac{p_{i}(y_{t}|y^{t-1})}{p_{j}(y_{t}|y^{t-1})}\biggr{\}}.$ (30)
The squared Euclidean distance in the $\bm{\zeta}$-coordinate system provides an estimate of the sum of (30) over all tokens in the text $x$:
$\|\bm{\zeta}_{i}-\bm{\zeta}_{j}\|^{2}$
$\approx$ $2\sum_{t=1}^{n}\mathrm{KL}(p_{i}(y_{t}|y^{t-1}),p_{j}(y_{t}|y^{t-1}))$ (31)
$\approx$ $2\mathrm{KL}(p_{i},p_{j}).$ (32)
The proof is provided in Appendix F. To justify the estimation in (31), we assume the following:
$\operatorname*{\mathbb{E}}_{y_{t}\sim p_{0}(y_{t}|y^{t-1})}\biggl{\{}\log\frac{p_{i}(y_{t}|y^{t-1})}{p_{j}(y_{t}|y^{t-1})}\biggr{\}}$ (33)
takes a constant value independent of $t$. In reality, this assumption is not entirely correct, and the degree of variation affects the accuracy of the approximation in (31) [^16] Since this assumption does not affect (3), there is no concern regarding the use of model maps based on text probabilities.. On the other hand, the approximation in (32) holds more generally and is demonstrated in Appendix F.5.
Appendix F Theory of Model Coordinates for Token Probability Distributions
In this section, we provide a more detailed explanation of the content discussed in Appendix E. We extend the discussion of text probability distributions in Appendix D to the case of conditional probability distributions for token generation.
F.1 Exponential family of distributions
We apply the same setting as for $p_{i}(x)$ in Section D to the conditional probability distributions of tokens:
$y_{t}\sim p_{i}(y_{t}|y^{t-1}),\quad t=1,\ldots,n.$
The exponential family of distributions incorporating $K$ models, corresponding to (6), is given here as
$p(y_{t}|y^{t-1};\bm{\theta}):=p_{0}(y_{t}|y^{t-1})$
$\qquad\exp(\bm{\theta}^{\top}\bm{b}(y_{t}|y^{t-1})-\psi(\bm{\theta}|y^{t-1})).$ (34)
The setting (7), which associates the $K$ models with (F.1), is given here as
$\lambda b_{i}(y_{t}|y^{t-1}):=$
$\quad\log p_{i}(y_{t}|y^{t-1})-\log p_{0}(y_{t}|y^{t-1}).$ (35)
Thus, we have
$p_{i}(y_{t}|y^{t-1})=p(y_{t}|y^{t-1};\lambda\bm{e}_{i})$
for $i=1,\ldots,K$.
F.2 The variance representation of the KL divergence
The KL divergence is given by (29). Applying the result for the model $p(x;\bm{\theta})$ in (17) to the token-level conditional distribution model $p(y_{t}|y^{t-1};\bm{\theta})$, we obtain
$2\mathrm{KL}(p_{i}(y_{t}|y^{t-1}),p_{j}(y_{t}|y^{t-1}))=$
$\operatorname*{Var}_{y_{t}\sim p_{0}(y_{t}|y^{t-1})}\biggl{\{}\log\frac{p_{i}(y_{t}|y^{t-1})}{p_{j}(y_{t}|y^{t-1})}\biggr{\}}+O(\lambda^{3}).$ (36)
F.3 Two additional assumptions
To estimate the KL divergence from a single text $x$, two additional assumptions are required, as described below. Such assumptions were not necessary when estimating the KL divergence from the dataset $D$ in Appendix D. In reality, these two assumptions are not strictly satisfied, and the discrepancy between these assumptions and reality affects the accuracy of the KL divergence approximation.
Assumption 1:
We assume that the probability distribution of $y_{t}$ depends only on the past $k$ tokens, denoted as $y^{t-1}_{t-k}=(y_{t-k},y_{t-k+1},\ldots,y_{t-1})$. That is,
$p_{i}(y_{t}|y^{t-1})=p_{i}(y_{t}|y^{t-1}_{t-k}),$
which allows us to regard $y^{t}_{t-k}$ as the state of a Markov chain. More generally, we use the notation $y^{k}$ to represent a state variable. We consider a function $f$ of the state variable $y^{k}$. Furthermore, we assume that this Markov chain is positive Harris recurrent, has a stationary distribution $\pi$, and that $f$ is absolutely integrable, i.e.,
$\operatorname*{\mathbb{E}}_{y^{k}\sim\pi}(|f(y^{k})|)<\infty.$
Then, by the strong law of large numbers for Markov chains (Meyn and Tweedie, 2009, Theorem 17.0.1 (i)), in the limit $n\to\infty$,
[…]
For simplicity in notation and discussion, we assume that $y_{-k+1},\ldots,y_{0}$ are appropriately defined. Since the Markov chain converges to $\pi$, we also have
$\frac{1}{n}\sum_{t=1}^{n}f(y^{t}_{t-k})\to\frac{1}{n}\sum_{t=1}^{n}\operatorname*{\mathbb{E}}_{y^{t}\sim p_{0}}(f(y^{t}_{t-k}))$ (37)
almost surely as $n\to\infty$.
Assumption 2:
$\operatorname*{\mathbb{E}}_{y_{t}\sim p_{0}(y_{t}|y^{t-1})}\log\frac{p_{i}(y_{t}|y^{t-1})}{p_{j}(y_{t}|y^{t-1})}=c$ (38)
for some $c\in\mathbb{R}$ that can depend on the indices $i$ and $j$ but not on $t$. In other words, (38) takes a constant value independent of $t$.
F.4 Estimation of the KL divergence
Define
$h(y^{t}):=\log\frac{p_{i}(y_{t}|y^{t-1})}{p_{j}(y_{t}|y^{t-1})}.$
From Assumption 1, $h(y^{t})$ can be written in the form $h(y^{t})=f_{1}(y^{t}_{t-k})$ for some $f_{1}$, so applying (37), for sufficiently large $n$, we obtain
$\frac{1}{n}\sum_{t=1}^{n}h(y^{t})\approx\frac{1}{n}\sum_{t=1}^{n}\operatorname*{\mathbb{E}}_{y^{t}\sim p_{0}}(h(y^{t})).$
Applying (38) to the right-hand side gives
$\frac{1}{n}\sum_{t=1}^{n}h(y^{t})\approx c.$
Next, since $(h(y^{t})-c)^{2}$ can be written in the form of $f_{2}(y^{t}_{t-k})$ for some function $f_{2}$, applying (37) again yields
$\frac{1}{n}\sum_{t=1}^{n}(h(y^{t})-c)^{2}$
$\approx$ $\frac{1}{n}\sum_{t=1}^{n}\operatorname*{\mathbb{E}}_{y^{t-1}\sim p_{0}}\Bigl{\{}\operatorname*{Var}_{y_{t}\sim p_{0}(y_{t}|y^{t-1})}(h(y^{t}))\Bigr{\}}$
$\approx$ $\frac{1}{n}\sum_{t=1}^{n}\operatorname*{Var}_{y_{t}\sim p_{0}(y_{t}|y^{t-1})}(h(y^{t})).$
In the final equation, we applied (37) using the fact that $\operatorname*{Var}_{y_{t}\sim p_{0}(y_{t}|y^{t-1})}(h(y^{t}))=f_{3}(y^{t}_{t-k})$ for some $f_{3}$. Using (36), we obtain
$\sum_{t=1}^{n}(h(y^{t})-c)^{2}\approx$
$\qquad 2\sum_{t=1}^{n}\mathrm{KL}(p_{i}(y_{t}|y^{t-1}),p_{j}(y_{t}|y^{t-1})).$ (39)
Meanwhile, the components of the model coordinate $\bm{\zeta}_{i}$ are given by
$\zeta_{it}=\log p_{i}(y_{t}|y^{t-1})-c_{i}$
where
$c_{i}=\frac{1}{n}\sum_{t=1}^{n}\log p_{i}(y_{t}|y^{t-1}).$
Since
$\zeta_{it}-\zeta_{jt}=h(y^{t})-(c_{i}-c_{j})$
with $c_{i}-c_{j}\approx c$, equation (39) can be rewritten as
$\|\bm{\zeta}_{i}-\bm{\zeta}_{j}\|^{2}\approx$
$\qquad 2\sum_{t=1}^{n}\mathrm{KL}(p_{i}(y_{t}|y^{t-1}),p_{j}(y_{t}|y^{t-1})).$
Thus, (31) is established.
F.5 Connecting the KL divergence of token and text probability distributions
Here, we fix the sequence length of the text $x=(y_{1},\ldots,y_{n})$ as $n$, i.e., we set $\mathcal{X}=\mathcal{V}^{n}$. For notational simplicity, we define
$g_{i}(y^{t})=\log p_{i}(y_{t}|y^{t-1}).$
Noting that
$p_{i}(x)=\prod_{t=1}^{n}p_{i}(y_{t}|y^{t-1})=\prod_{t=1}^{n}e^{g_{i}(y^{t})},$
we obtain
$\textrm{KL}(p_{i},p_{j})$
$=$ $\sum_{x\in\mathcal{X}}\prod_{t^{\prime}=1}^{n}e^{g_{i}(y^{t^{\prime}})}\sum_{t=1}^{n}(g_{i}(y^{t})-g_{j}(y^{t}))$
$=$ $\sum_{t=1}^{n}\sum_{y^{t}\in\mathcal{V}^{t}}\prod_{t^{\prime}=1}^{t}e^{g_{i}(y^{t^{\prime}})}(g_{i}(y^{t})-g_{j}(y^{t}))$
$=$ $\sum_{t=1}^{n}\sum_{y^{t-1}\in\mathcal{V}^{t-1}}\prod_{t^{\prime}=1}^{t-1}e^{g_{i}(y^{t^{\prime}})}$
$\qquad\sum_{y_{t}\in\mathcal{V}}e^{g_{i}(y^{t})}(g_{i}(y^{t})-g_{j}(y^{t}))$
$=$ $\sum_{t=1}^{n}\sum_{y^{t-1}\in\mathcal{V}^{t-1}}p_{i}(y^{t-1})$
$\qquad\textrm{KL}(p_{i}(y_{t}|y^{t-1}),p_{j}(y_{t}|y^{t-1}))$
$=$ $\sum_{t=1}^{n}\operatorname*{\mathbb{E}}_{y^{t-1}\sim p_{i}}\Bigl{\{}\textrm{KL}(p_{i}(y_{t}|y^{t-1}),p_{j}(y_{t}|y^{t-1}))\Bigr{\}}$
$=$ $\operatorname*{\mathbb{E}}_{x\sim p_{i}}\Bigl{\{}\sum_{t=1}^{n}\textrm{KL}(p_{i}(y_{t}|y^{t-1}),p_{j}(y_{t}|y^{t-1}))\Bigr{\}}.$
Thus, for sufficiently large $n$, by the strong law of large numbers for Markov chains, we obtain
$\textrm{KL}(p_{i},p_{j})\approx\sum_{t=1}^{n}\textrm{KL}(p_{i}(y_{t}|y^{t-1}),p_{j}(y_{t}|y^{t-1})).$ (40)
This corresponds to (32). Assumption 1 from Appendix F.3 is used in (40), but Assumption 2 is not needed in the discussion of this subsection.
Appendix G Details of Experiments
G.1 Information obtained via the Hugging Face Hub API
We used the Hugging Face Hub API to retrieve information about each language model’s tags, the date the model was created, the number of downloads over the past 30 days, and the model’s configuration details. All of this information is current as of February 1, 2025.
Among the model tags, we specifically used llama2, llama-2, license:llama2, llama3, llama-3, and license:llama3 to determine the model type (llama-1, llama-2, or llama-3). Furthermore, to identify language models that were pre-trained on the Pile in Section 4, we employed tags such as dataset:eleutherai/pile, dataset:eleutherai/the_pile, dataset:eleutherai/the_pile_deduplicated, and arxiv:2101.00027.
G.2 How the model type was determined
Model type Models
llama-1 69
llama-2 223
llama-3 62
llama 217
mistral 232
gpt_neox 54
deepseek 26
gptj 19
gemma 18
opt 15
bloom 12
falcon 11
qwen2 10
mixtral 9
mpt 6
stablelm 6
gpt_neo 3
phi 3
gpt_bigcode 3
phi3 3
xglm 3
rwkv 3
starcoder2 2
olmo 2
camelidae 2
codegen 2
deci 1
recurrent_gemma 1
stablelm_alpha 1
Total 1,018
Table 3: Number of models by model type.
In principle, we used the value of model_type in the config retrieved from the Hugging Face Hub API as the model type. However, out of the 1,018 language models we examined, there were 587 whose config model_type was listed as llama. For these, we used the following procedure to determine whether they were the original Llama (llama-1), Llama-2, or Llama-3; if we were able to identify which version they were, we reclassified them as llama-1, llama-2, or llama-3 accordingly.
  1. 1.
    We checked the tags assigned to each model. Of these, 136 models that included any of llama2, llama-2, or license:llama2 were classified as llama-2. Similarly, 39 models that included any of llama3, llama-3, or license:llama3 were classified as llama-3.
  2. 2.
    For the remaining 412 models whose classification was not determined by tags alone, we used the creation date and the model name (converted to lowercase) to make a decision. First, 69 models that were created prior to July 18, 2023 (the Llama-2 release date) were classified as llama-1. Next, 88 models whose lowercase model name contained either llama2 or llama-2 were classified as llama-2. Among those whose lowercase model name contained llama3 or llama-3, 22 models whose creation date was after April 18, 2024 (the Llama-3 release date) were classified as llama-3.
  3. 3.
    After following the steps above, the 217 models that could not be classified were left as llama.
Furthermore, any model whose name prior to the slash (/) was deepseek-ai was defined as deepseek. In addition, even though abacusai/Llama-3-Smaug-8B was tagged with license:llama2, we manually reclassified it as llama-3.
Table 3 shows the number of models classified into each model type.
G.3 Basic information on the dataset
Text category Texts
Pile-CC 2,353
PubMed Central 1,763
ArXiv 1,172
Github 925
FreeLaw 837
StackExchange 712
Wikipedia (en) 567
USPTO Backgrounds 487
PubMed Abstracts 464
Gutenberg (PG-19) 251
DM Mathematics 151
EuroParl 83
HackerNews 67
Ubuntu IRC 54
PhilPapers 51
NIH ExPorter 41
Enron Emails 22
Total 10,000
Table 4: Number of texts in each text category.
The dataset used in our experiments consists of a total of 10,000 texts, which are divided into 17 text categories. Table 4 shows the number of texts in each category.
To assign colors to the text categories, we first compute the average text embedding for each category in the Pile using simcse-roberta-large Gao et al. (2021b). Next, we calculate a tour over the 17 average embedding vectors by solving the traveling salesman problem [^17] Inspired by Sato (2022); Yamagiwa et al. (2024).. The TSP is solved using the nearest neighbor method to generate an initial tour, which is then refined using a 2-opt improvement procedure, and Euclidean distance is used as the metric. Based on the adjacency relationships along this tour, we segment the hue circle at equal intervals and color each category accordingly.
G.4 Standard scores
In the three experiments described below, we use standardized values [^18] By subtracting the mean from each value and dividing by the standard deviation, the data is transformed to have a mean of 0 and a variance of 1., or $Z$-score normalization, of both the log-likelihood and the benchmark scores calculated for the $K$ language models.
  • In Figures 1 and 11, where we define each language model’s primary text category, we use the average log-likelihood for each category, standardized across all models.
  • In Section 4, to determine each language model’s primary task, we standardize each task’s score across the $K$ models.
  • Furthermore, in the data leakage detection described in Section 4, we use the difference between the standardized mean log-likelihood and the standardized 6-TaskMean score as the indicator.
Figure 11: Model maps obtained by double centered log-likelihood matrix $\bm{Q}$. These maps correspond to Figure 1. (Top) Colors indicate model types. (Bottom) Colors indicate the text category in which each model attains the highest standardized log-likelihood score among 17 categories.
Figure 12: Model maps obtained by double centered log-likelihood matrix $\bm{Q}$. These maps correspond to Figure 5. These maps are illustrating model performance. From left to right, the panels show each model’s mean log-likelihood, 6-TaskMean score, and the “primary task,” which refers to the task where each model achieves the highest standardized score among the six tasks, color-coded accordingly. The color bar is clipped at the 10th percentile for mean log-likelihood and 6-TaskMean, with darker colors indicating better performance. In the primary task panel, models with standardized scores below zero across all six tasks are labeled as “All Under 0.”
Figure 13: Model maps obtained by double centered log-likelihood matrix $\bm{Q}$ color-coded by (Left) model size and (Right) model creation date.
G.5 Hierarchical clustering settings
Figure 3 displays the double-centered log-likelihood matrix $\bm{Q}$, with hierarchical clustering applied to both its rows and columns. We implemented the clustering using SciPy (Virtanen et al., 2020). Distance matrices were computed using scipy.spatial.distance.pdist, and clustering was performed using scipy.cluster.hierarchy.linkage. For clustering models, we used sqeuclidean as the metric and median as the linkage method. For clustering texts, we used correlation as the metric and average as the linkage method. For the hierarchical clustering shown in Fig. 4, which presents a dendrogram of 100 language models, we used sqeuclidean as the metric and median as the linkage method. The vertical axis of the dendrogram uses the symmetric logarithmic scale (with a linear threshold of 250) implemented in matplotlib. To ensure that the values on the vertical axis correspond to the Kullback-Leibler divergence, we used the $\bm{q}$-coordinates divided by $\sqrt{2N}$.
Appendix H Additional Model Maps
In this section, we present additional model maps, including a figure that lists all the model names and a map obtained through dimensionality reduction of the double-centered log-likelihood matrix $\bm{Q}$.
H.1 Model map via the double centered log-likelihood matrix
In the main text, we use a model map generated by dimensionality reduction of the log-likelihood matrix $\bm{L}$. Here, in Figs. 11, 12 and 13, we present model maps obtained by dimensionality reduction of the double-centered log-likelihood matrix $\bm{Q}$.
H.2 Model map with model names
Figure 14: Model map obtained by dimensionality reduction of the log-likelihood matrix $\bm{L}$. Each point on the model map is labeled with the corresponding model name.
Figure 15: Model map obtained by dimensionality reduction of the log-likelihood matrix $\bm{L}$. Each point on the model map is labeled with the corresponding model name. Colors indicate the model’s “primary text category,” the text category where the model achieves the highest standardized log-likelihood score among 17 categories.
Figure 16: Model map obtained by dimensionality reduction of the double-centered log-likelihood matrix $\bm{Q}$. Each point on the model map is labeled with the corresponding model name.
Figure 17: Model map obtained by dimensionality reduction of the double-centered log-likelihood matrix $\bm{Q}$. Each point on the model map is labeled with the corresponding model name. Colors indicate the model’s “primary text category,” the text category where the model achieves the highest standardized log-likelihood score among 17 categories.
We present figures that display the model names corresponding to each point on the model maps defined in Section 4. For both the log-likelihood matrix $\bm{L}$ and the double-centered log-likelihood matrix $\bm{Q}$, we provide two types of maps: one colored by model type and another colored by primary text category. Figures 14 and 15 show the maps obtained by applying dimensionality reduction to $\bm{L}$, while Figures 16 and 17 show the maps obtained using $\bm{Q}$.
Appendix I Table of Nearest Neighbor Models
Table 5 presents the top 10 nearest neighbors among the 1,018 language models for each of the models highlighted in the top panel of Figure 1. Each sub-table corresponds to a specific model of interest, listing its nearest neighbors in descending order of the KL divergence. From this table, we can see that similar models tend to cluster together, exhibiting relatively small KL divergence values. Additionally, the values in parentheses denote the KL divergence measured in bits per byte of text. This conversion makes it easier to interpret the information cost per byte when comparing predictive distributions of different models.
bigcode/starcoder2-7b KL (/byte) codellama/CodeLlama-7b-Instruct-hf KL (/byte) bigcode/starcoder2-3b 769.780 (1.142) NousResearch/CodeLlama-7b-hf 48.192 (0.072) stabilityai/stable-code-3b 1606.483 (2.384) codellama/CodeLlama-7b-hf 48.192 (0.072) deepseek-ai/deepseek-coder-6.7b-base 1687.186 (2.503) codellama/CodeLlama-13b-Instruct-hf 138.606 (0.206) EleutherAI/llemma_7b 1892.812 (2.808) TheBloke/CodeLlama-13B-Instruct-fp16 138.627 (0.206) deepseek-ai/deepseek-coder-7b-base-v1.5 2031.482 (3.014) Nexusflow/NexusRaven-V2-13B 147.114 (0.218) deepseek-ai/DeepSeek-Coder-V2-Lite-Base 2042.158 (3.030) codellama/CodeLlama-13b-hf 158.786 (0.236) deepseek-ai/deepseek-coder-7b-instruct-v1.5 2045.330 (3.035) NousResearch/CodeLlama-13b-hf 158.949 (0.236) deepseek-ai/deepseek-coder-6.7b-instruct 2140.788 (3.176) OpenAssistant/codellama-13b-oasst-sft-v10 219.381 (0.326) deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct 2149.008 (3.189) HiTZ/GoLLIE-7B 225.804 (0.335) meta-math/MetaMath-Llemma-7B 2184.274 (3.241) WhiteRabbitNeo/WhiteRabbitNeo-13B-v1 260.047 (0.386) deepseek-ai/deepseek-coder-1.3b-base KL (/byte) deepseek-ai/deepseek-llm-7b-base KL (/byte) deepseek-ai/deepseek-coder-1.3b-instruct 410.897 (0.610) deepseek-ai/deepseek-moe-16b-base 130.715 (0.194) deepseek-ai/deepseek-coder-6.7b-instruct 512.997 (0.761) deepseek-ai/deepseek-llm-7b-chat 245.408 (0.364) deepseek-ai/deepseek-coder-6.7b-base 601.002 (0.892) deepseek-ai/deepseek-moe-16b-chat 247.889 (0.368) bigcode/starcoderbase-1b 798.322 (1.185) deepseek-ai/DeepSeek-V2-Lite 306.530 (0.455) bigcode/starcoderbase-7b 1249.899 (1.855) deepseek-ai/ESFT-vanilla-lite 307.487 (0.456) NTQAI/Nxcode-CQ-7B-orpo 1307.618 (1.940) deepseek-ai/DeepSeek-V2-Lite-Chat 585.090 (0.868) Qwen/CodeQwen1.5-7B-Chat 1317.166 (1.954) mistralai/Mistral-7B-Instruct-v0.1 913.708 (1.356) bigcode/starcoder2-3b 1858.329 (2.757) statking/zephyr-7b-sft-full-orpo 1089.127 (1.616) Salesforce/codegen-6B-multi 1974.095 (2.929) Severian/ANIMA-Phi-Neptune-Mistral-7B 1140.161 (1.692) google/codegemma-2b 2115.462 (3.139) sethuiyer/Medichat-Llama3-8B 1157.599 (1.718) EleutherAI/gpt-neo-1.3B KL (/byte) EleutherAI/pythia-12b KL (/byte) EleutherAI/gpt-neo-2.7B 218.723 (0.325) matsuo-lab/weblab-10b 139.270 (0.207) EleutherAI/pythia-1.4b 289.377 (0.429) h2oai/h2ogpt-oig-oasst1-256-6_9b 159.560 (0.237) EleutherAI/pythia-1b-deduped 413.204 (0.613) Salesforce/codegen-6B-nl 175.546 (0.260) HWERI/pythia-1.4b-deduped-sharegpt 421.551 (0.625) EleutherAI/gpt-j-6b 186.613 (0.277) beaugogh/pythia-1.4b-deduped-sharegpt 421.551 (0.625) TehVenom/Dolly_Malion-6b 231.701 (0.344) EleutherAI/pythia-1.4b-deduped 447.122 (0.663) TehVenom/PPO_Shygmalion-6b 233.159 (0.346) PygmalionAI/metharme-1.3b 448.970 (0.666) TehVenom/Dolly_Shygmalion-6b-Dev_V8P2 236.753 (0.351) RWKV/rwkv-raven-1b5 512.789 (0.761) TehVenom/PPO_Pygway-V8p4_Dev-6b 237.270 (0.352) databricks/dolly-v2-3b 525.574 (0.780) TehVenom/GPT-J-Pyg_PPO-6B-Dev-V8p4 245.015 (0.364) EleutherAI/pythia-2.8b-deduped 587.699 (0.872) TehVenom/PPO_Shygmalion-V8p4_Dev-6b 245.364 (0.364) facebook/opt-6.7b KL (/byte) google/codegemma-2b KL (/byte) KoboldAI/OPT-6.7B-Erebus 0.685 (0.001) deepseek-ai/deepseek-coder-1.3b-instruct 1654.767 (2.455) KoboldAI/OPT-6.7B-Nerybus-Mix 78.604 (0.117) bigcode/starcoderbase-1b 1812.768 (2.690) KoboldAI/OPT-6B-nerys-v2 124.520 (0.185) deepseek-ai/deepseek-coder-1.3b-base 2115.462 (3.139) facebook/opt-13b 196.634 (0.292) bigcode/gpt_bigcode-santacoder 2644.348 (3.924) KoboldAI/OPT-13B-Nerybus-Mix 230.046 (0.341) deepseek-ai/deepseek-coder-6.7b-instruct 2675.004 (3.969) KoboldAI/OPT-13B-Erebus 237.845 (0.353) Qwen/CodeQwen1.5-7B-Chat 2759.243 (4.094) KoboldAI/OPT-13B-Nerys-v2 273.410 (0.406) NTQAI/Nxcode-CQ-7B-orpo 2767.578 (4.106) facebook/opt-2.7b 327.851 (0.486) Salesforce/codegen-6B-multi 2859.701 (4.243) KoboldAI/OPT-2.7B-Nerybus-Mix 427.279 (0.634) bigcode/starcoderbase-7b 2990.326 (4.437) KoboldAI/OPT-2.7B-Erebus 446.712 (0.663) deepseek-ai/deepseek-coder-6.7b-base 3282.352 (4.870) google/gemma-7b KL (/byte) lmsys/vicuna-13b-v1.3 KL (/byte) SeaLLMs/SeaLLM-7B-v2.5 247.533 (0.367) TheBloke/stable-vicuna-13B-HF 195.152 (0.290) VAGOsolutions/SauerkrautLM-Gemma-7b 257.986 (0.383) Yhyu13/chimera-inst-chat-13b-hf 197.130 (0.292) lemon-mint/gemma-ko-7b-instruct-v0.62 344.582 (0.511) junelee/wizard-vicuna-13b 219.148 (0.325) google/gemma-2b 503.918 (0.748) TheBloke/wizard-vicuna-13B-HF 219.148 (0.325) google/codegemma-7b 596.383 (0.885) TheBloke/UltraLM-13B-fp16 219.689 (0.326) PathFinderKR/Waktaverse-Llama-3-KO-8B-Instruct 814.680 (1.209) NousResearch/Nous-Hermes-13b 221.324 (0.328) FlagAlpha/Llama3-Chinese-8B-Instruct 817.954 (1.214) project-baize/baize-v2-13b 233.813 (0.347) FairMind/Llama-3-8B-4bit-UltraChat-Ita 832.782 (1.236) TheBloke/guanaco-13B-HF 234.090 (0.347) migtissera/Llama-3-8B-Synthia-v3.5 842.861 (1.251) openaccess-ai-collective/minotaur-13b-fixed 235.470 (0.349) Orenguteng/Llama-3-8B-Lexi-Uncensored 844.748 (1.253) openaccess-ai-collective/wizard-mega-13b 238.073 (0.353) medalpaca/medalpaca-7b KL (/byte) meta-llama/Llama-2-13b-hf KL (/byte) TheBloke/guanaco-7B-HF 285.806 (0.424) TaylorAI/Flash-Llama-13B 0.000 (0.000) eachadea/vicuna-7b-1.1 336.518 (0.499) TheBloke/Llama-2-13B-fp16 0.002 (0.000) TehVenom/Pygmalion-Vicuna-1.1-7b 369.003 (0.548) StudentLLM/Alpagasus-2-13b-QLoRA-merged 4.502 (0.007) lmsys/vicuna-7b-v1.3 374.469 (0.556) CHIH-HUNG/llama-2-13b-FINETUNE2_TEST_2.2w 7.610 (0.011) jphme/orca_mini_v2_ger_7b 383.382 (0.569) garage-bAInd/Platypus2-13B 10.373 (0.015) ajibawa-2023/Uncensored-Jordan-7B 384.396 (0.570) CHIH-HUNG/llama-2-13b-dolphin_5w 14.386 (0.021) bofenghuang/vigogne-7b-instruct 390.797 (0.580) CHIH-HUNG/llama-2-13b-OpenOrca_5w 14.982 (0.022) TheBloke/airoboros-7b-gpt4-fp16 391.991 (0.582) CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r16-gate_up_down-test1 17.256 (0.026) TheBloke/tulu-7B-fp16 423.476 (0.628) CHIH-HUNG/llama-2-13b-FINETUNE4_addto15k_4.5w-r16-gate_up_down 17.405 (0.026) TehVenom/Pygmalion_AlpacaLora-7b 431.540 (0.640) CHIH-HUNG/llama-2-13b-FINETUNE4_compare15k_4.5w-r16-gate_up_down 19.427 (0.029) meta-llama/Llama-2-7b-hf KL (/byte) meta-llama/Meta-Llama-3-8B KL (/byte) meta-llama/Llama-2-7b-hf 0.000 (0.000) Undi95/Meta-Llama-3-8B-hf 0.003 (0.000) TheTravellingEngineer/llama2-7b-chat-hf-v4 0.000 (0.000) dfurman/Llama-3-8B-Orpo-v0.1 6.994 (0.010) ibranze/araproje-llama2-7b-hf 0.000 (0.000) migtissera/Tess-2.0-Llama-3-8B 28.829 (0.043) yeen214/test_llama2_7b 0.000 (0.000) freewheelin/free-llama3-dpo-v0.2 67.783 (0.101) TheTravellingEngineer/llama2-7b-chat-hf-v2 0.000 (0.000) jondurbin/bagel-8b-v1.0 110.365 (0.164) NewstaR/Starlight-7B 0.003 (0.000) migtissera/Llama-3-8B-Synthia-v3.5 128.141 (0.190) Delcos/Mistral-Pygmalion-7b 14.809 (0.022) nvidia/Llama3-ChatQA-1.5-8B 128.776 (0.191) elliotthwang/elliott_Llama-2-7b-hf 16.612 (0.025) ruslanmv/Medical-Llama3-8B 138.566 (0.206) garage-bAInd/Platypus2-7B 22.811 (0.034) FairMind/Llama-3-8B-4bit-UltraChat-Ita 199.748 (0.296) Lazycuber/L2-7b-Base-Guanaco-Uncensored 23.345 (0.035) NousResearch/Hermes-2-Theta-Llama-3-8B 222.357 (0.330) meta-math/MetaMath-Mistral-7B KL (/byte) mistralai/Mistral-7B-v0.3 KL (/byte) Weyaxi/MetaMath-NeuralHermes-2.5-Mistral-7B-Linear 43.095 (0.064) MaziyarPanahi/Mistral-7B-v0.3 0.000 (0.000) Weyaxi/MetaMath-Tulpar-7b-v2-Slerp 81.882 (0.121) mistral-community/Mistral-7B-v0.2 7.221 (0.011) Weyaxi/MetaMath-OpenHermes-2.5-neural-chat-v3-3-Slerp 101.097 (0.150) unsloth/mistral-7b-v0.2 7.222 (0.011) Q-bert/Bumblebee-7B 106.376 (0.158) mistralai/Mistral-7B-v0.1 22.038 (0.033) OpenPipe/mistral-ft-optimized-1227 110.089 (0.163) Cartinoe5930/Llama2_init_Mistral 29.819 (0.044) Toten5/Marcoroni-neural-chat-7B-v2 114.004 (0.169) Locutusque/Hercules-3.1-Mistral-7B 32.099 (0.048) ignos/Mistral-T5-7B-v1 114.320 (0.170) migtissera/Synthia-7B-v3.0 33.418 (0.050) Weyaxi/MetaMath-Chupacabra-7B-v2.01-Slerp 114.414 (0.170) uukuguy/speechless-zephyr-code-functionary-7b 35.723 (0.053) Q-bert/Optimus-7B 117.829 (0.175) uukuguy/zephyr-7b-alpha-dare-0.85 35.745 (0.053) Weyaxi/MetaMath-NeuralHermes-2.5-Mistral-7B-Ties 127.457 (0.189) crumb/apricot-wildflower-20 40.907 (0.061)
Table 5: Top 10 nearest neighbors among the 1,018 language models for the models labeled in the top panel of Fig. 1. The KL divergence is computed using formula (3) in Section 2. The values in parentheses represent the KL divergence measured in bits per byte of text (Section 3.5), obtained by multiplying the KL divergence by 0.001484.
Figure 18: Scatter plots of predicted scores and benchmark scores for six benchmark tasks. Additionally, results for predicting 6-TaskMean (identical to Fig. 8) and the mean log-likelihood are also shown. Each point is color-coded by the mean log-likelihood, with higher mean log-likelihood values generally corresponding to higher task scores. For better visualization, the color bar range is clipped to the 10th–100th percentile.
Appendix J Details of Weight Interpolation
In this section, we describe the experimental details concerning the relationship between model coordinates and model weights, as introduced in Section 6.3.
Constructing the weight grid.
Let $p_{0}$ denote the base model, and $p_{1}$ and $p_{2}$ denote the fine-tuned models derived from $p_{0}$. We denote the weight parameter vectors of these models as $W_{0}$, $W_{1}$, and $W_{2}$, respectively. To construct the weight grid, we merged the model weights using the following linear operation:
$W_{\alpha,\beta}=W_{0}+\alpha(W_{1}-W_{0})+\beta(W_{2}-W_{0}),$ (41)
where the merge ratios $\alpha,\beta\in\mathbb{R}$ were chosen from 36 evenly spaced combinations within the interval $[0,1]$: $\{0.0,0.2,0.6,0.8,1.0\}$. The original models $W_{0}$, $W_{1}$, and $W_{2}$ correspond to $W_{0,0}$, $W_{1,0}$, and $W_{0,1}$, respectively. When $\alpha+\beta\leq 1$, the operation corresponds to linear interpolation between models. Even among models with the same architecture, the sizes of the embedding/unembedding matrices may differ. In such cases, we truncated or reshaped the weight parameters to match the base model.
Computing model coordinates.
For each composed model $p_{\alpha,\beta}$ with weight $W_{\alpha,\beta}$, we computed the model coordinates $\bm{q}_{\alpha,\beta}$ following the method described in Section 2. The tokenizer of the base model was used to ensure consistency. The text data consists of 1,000 randomly sampled sentences from the dataset of 10,000 sentences created in Section 3.1.
Selection of models.
We set the base model $p_{0}$ as Llama-2-7b-hf. Among the fine-tuned models based on Llama-2-7b-hf, we selected the two most downloaded models: vicuna-7b-v1.5 and Llama-2-7b-chat-hf, denoted as $p_{1}$ and $p_{2}$, respectively.
Visualization.
Figure 10 shows the linearly merged models, visualized in both weight space and log-likelihood space. The corners of the grid are labeled with their corresponding $(\alpha,\beta)$ values.
In the left panel, we visualized $W_{\alpha,\beta}$ in the weight space. Since the dimensionality of $W_{\alpha,\beta}$, i.e., the number of model parameters, is extremely high, we employed a 2D projection method using the norms of the difference vectors $r_{1}=\left\|W_{1}-W_{0}\right\|_{2}$, $r_{2}=\left\|W_{2}-W_{0}\right\|_{2}$, and the angle between them, $\phi=\arccos((W_{1}-W_{0})^{\top}(W_{2}-W_{0})/r_{1}r_{2})$. Each point was placed at $(\alpha r_{1}+\beta r_{2}\cos\phi,\beta r_{2}\sin\phi)$.
In the right panel, we visualized the model coordinates $\bm{q}_{\alpha,\beta}$ by Principal Component Analysis (PCA). Each model $p_{\alpha,\beta}$ was mapped onto the $\bm{q}$-coordinate system to analyze the structure of the interpolated models.
ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K Average mean log-likelihood
     Pearson’s $r$ 0.969$\pm$0.002 0.943$\pm$0.004 0.959$\pm$0.003 0.954$\pm$0.001 0.961$\pm$0.002 0.930$\pm$0.002 0.973$\pm$0.001 0.993$\pm$0.001
     Spearman’s $\rho$ 0.975$\pm$0.001 0.971$\pm$0.002 0.966$\pm$0.003 0.933$\pm$0.001 0.971$\pm$0.002 0.892$\pm$0.001 0.976$\pm$0.001 0.989$\pm$0.001
Table 6: This table presents the mean and standard deviation of correlation coefficients between predicted and actual benchmark scores. These coefficients were computed using ridge regression across five different data splits. Results for predicting 6-TaskMean and the mean log-likelihood are also included.
Appendix K Details of Model Performance Prediction
This section provides additional details on the prediction of benchmark scores using model coordinates, as discussed in Section 5.
K.1 Details of ridge regression
As described in Section 5.2, ridge regression requires setting a regularization strength parameter, $\alpha$. To determine $\alpha$ from $\{10^{1},\ldots,10^{9}\}$, we performed a five-fold cross-validation within each training dataset [^19] The training dataset consisted of four out of the five folds obtained by splitting the dataset for each benchmark task.. As a post-processing step, we clipped the predicted scores to the range $[0,100]$.
For the setting where the target variable $\mathbf{f}$ was replaced with the mean log-likelihood (¯$(\bar{\ell}_{1},\ldots,\bar{\ell}_{K})\in\mathbb{R}^{K}$, we searched for $\alpha$ within $\{10^{-4},\ldots,10^{4}\}$ and did not apply clipping as a post-processing step.
K.2 Details of prediction results
Figure 18 shows scatter plots of predicted scores and actual benchmark scores for each benchmark task, as well as for 6-TaskMean and the mean log-likelihood. As in Fig. 8, the scatter plots show strong correlations for all six benchmark tasks and for the mean log-likelihood.
To account for randomness, we ran five different data splits when predicting each benchmark score. The final predicted score was the average of these five runs. For each split, we computed the correlation coefficients between the predicted scores and the actual benchmark scores. Table 6 presents their mean and standard deviation, and shows a similar trend as Table 2.
Appendix L Model List
Table LABEL:tab:model_table lists the 1,018 models used in this study, sorted alphabetically by their names. The BibTeX entries cited for each model were determined through the following procedure.
First, we extracted the BibTeX entries available in each model’s Hugging Face model card [^20] https://github.com/huggingface/huggingface_hub.. If the BibTeX entry was missing a year of publication, we filled it in with the model’s creation date [^21] https://github.com/sciunto-org/python-bibtexparser.. Additionally, we generated BibTeX entries using the arXiv IDs found in the model card tags by querying the arXiv API [^22] https://github.com/lukasschwab/arxiv.py.. This process resulted in a set of BibTeX entries for each model.
Next, we manually checked pairs of different BibTeX entries where the title similarity [^23] We used Python’s difflib.SequenceMatcher. was high, or the authors matched, to determine whether they corresponded to the same source. This step allowed us to create groups of BibTeX entries that were considered identical.
Then, for each BibTeX group, we selected a representative entry as follows. Within each group, the entry most frequently cited by the models was chosen as the representative. If multiple candidates met this criterion, we prioritized BibTeX entries generated from arXiv IDs when available. If no such entry existed, we selected the one with the longest string.
Finally, we replaced each model’s BibTeX entry with the representative entry from its corresponding group. Any selected BibTeX entry that contained typos or formatting errors was manually corrected based on compilation errors. If the author information was incomplete, we corrected it manually by checking the source.
Note that for google/codegemma-2b and deepseek-ai/deepseek-llm-7b-base in Table 1, as well as for deepseek-ai/deepseek-coder-1.3b-base and mistralai/Mistral-7B-v0.3 in Table 5, we manually prepared the BibTeX entries for citation based on their respective sources. We also used the same BibTeX entries for all other models that were considered to be of the same type.
Table 7: List of 1018 models. “ID” denotes the alphabetical index; “Model Name” denotes the name of the model; “Model Type” denotes the classification defined in this paper; “Size” denotes the size of the model (B: billion); “Date” denotes the date of model creation; “DLs” denotes the total number of downloads; ¯$\bar{\ell}_{i}$ denotes the mean log-likelihood; “Task” denotes the mean of the 6 benchmark scores (i.e., 6-TaskMean).
ID Model Name Model Type Size Date DLs ¯$\bar{\ell}_{i}$ Task
1 aaditya/Llama3-OpenBioLLM-8B Singhal et al. (2022); Nori et al. (2023); Singhal et al. (2023); Pal and Sankarasubbu (2024a, b); Rafailov et al. (2024) llama-3 8B 2024-04-20 9308 -590.14 54.06
2 aari1995/germeo-7b-laser mistral 7B 2024-01-09 2351 -686.18 62.82
3 abacusai/bigstral-12b-32k mistral 12B 2024-03-06 5216 -632.41 62.17
4 abacusai/Giraffe-13b-32k-v3 llama-2 13B 2023-12-06 1214 -546.12 57.24
5 abacusai/Llama-3-Smaug-8B Pal et al. (2024) llama-3 8B 2024-04-19 12873 -565.82 64.61
6 abacusai/Slerp-CM-mist-dpo mistral 7B 2024-01-03 4030 -553.48 73.10
7 abhinand/Llama-3-OpenBioMed-8B-slerp-v0.3 llama-3 8B 2024-05-02 2713 -558.51 62.08
8 abhinand/tamil-llama-7b-base-v0.1 Balachandran (2023) llama-2 7B 2023-11-08 1382 -877.00 44.52
9 abhinand/tamil-llama-7b-instruct-v0.1 Balachandran (2023) llama-2 7B 2023-11-08 3507 -708.97 45.52
10 abhishekchohan/mistral-7B-forest-dpo mistral 7B 2024-01-21 1961 -561.51 63.28
11 abhishekchohan/Yi-9B-Forest-DPO-v1.0 llama 9B 2024-03-18 2755 -528.97 64.11
12 acrastt/Bean-3B llama 3B 2023-09-02 1191 -592.51 40.18
13 acrastt/Griffin-3B llama 3B 2023-08-18 1198 -584.63 41.13
14 acrastt/Marx-3B-V2 llama 3B 2023-08-22 1234 -609.79 42.08
15 acrastt/Marx-3B llama 3B 2023-08-15 2006 -603.32 41.71
16 acrastt/OmegLLaMA-3B llama 3B 2023-08-26 1201 -640.98 38.28
17 acrastt/Puma-3B llama 3B 2023-08-16 1198 -584.54 41.02
18 acrastt/RedPajama-INCITE-Chat-Instruct-3B-V1 gpt_neox 2B 2023-07-27 1202 -565.14 39.23
19 adamo1139/Mistral-7B-AEZAKMI-v1 mistral 7B 2023-11-27 1199 -596.56 54.92
20 adonlee/LLaMA_2_13B_SFT_v0 llama 13B 2023-10-03 1268 -556.59 57.31
21 adonlee/LLaMA_2_13B_SFT_v1 llama 13B 2023-11-06 1265 -546.75 63.04
22 aerdincdal/CBDDO-LLM-8B-Instruct-v1 llama 8B 2024-05-02 4114 -613.44 56.94
23 ahnyeonchan/OpenOrca-AYT-13B llama-2 13B 2023-09-07 1184 -569.10
24 AIChenKai/TinyLlama-1.1B-Chat-v1.0-x2-MoE mixtral 1B 2024-01-03 1217 -619.71 36.98
25 AIJUUD/juud-Mistral-7B-dpo Lacoste et al. (2019) mistral 7B 2024-02-07 3217 -567.00 60.89
26 AIJUUD/juud-Mistral-7B mistral 7B 2024-01-31 1312 -564.68 61.72
27 ajibawa-2023/carl-7b llama 7B 2023-07-22 1212 -599.09 46.16
28 ajibawa-2023/Code-13B llama 13B 2023-12-08 1181 -604.28 54.81
29 ajibawa-2023/Python-Code-13B llama 13B 2023-11-11 1201 -582.71 53.61
30 ajibawa-2023/SlimOrca-13B llama 13B 2023-11-27 1177 -608.08 60.39
31 ajibawa-2023/Uncensored-Frank-13B llama 13B 2023-09-14 1192 -577.46 55.64
32 ajibawa-2023/Uncensored-Frank-7B llama 7B 2023-09-14 1178 -654.45 47.90
33 ajibawa-2023/Uncensored-Jordan-13B llama 13B 2023-10-23 1174 -578.48 56.27
34 ajibawa-2023/Uncensored-Jordan-7B llama 7B 2023-10-23 1174 -636.79 49.95
35 akjindal53244/Mistral-7B-v0.1-Open-Platypus mistral 7B 2023-10-05 1286 -543.69 58.92
36 AlekseyKorshuk/pygmalion-6b-vicuna-chatml gptj 6B 2023-06-22 1169 -564.90 42.08
37 alignment-handbook/zephyr-7b-sft-full mistral 7B 2023-11-09 11605 -569.88 57.52
38 allbyai/ToRoLaMa-7b-v1.0 Do et al. (2023) llama-2 7B 2023-12-19 1185 -732.57 47.87
39 allenai/OLMo-1B-hf Touvron et al. (2023a); Groeneveld et al. (2024) olmo 1B 2024-04-12 15221 -619.51 36.78
40 allenai/OLMo-7B-hf Touvron et al. (2023a); Groeneveld et al. (2024) olmo 6B 2024-04-12 5407 -554.02 43.36
41 allknowingroger/MultiverseEx26-7B-slerp mistral 7B 2024-03-30 3034 -599.87 76.80
42 alnrg2arg/blockchainlabs_7B_merged_test2_4_prune mistral 7B 2024-01-18 1939 -673.45 57.91
43 alnrg2arg/blockchainlabs_7B_merged_test2_4 mistral 7B 2024-01-17 1455 -577.39 75.28
44 aloobun/bun_mistral_7b_v2 mistral 7B 2023-12-20 1305 -558.93 59.76
45 aloobun/falcon-1b-cot-t2 falcon 1B 2024-01-07 2161 -735.39 28.56
46 aloobun/open-llama-3b-v2-elmv3 llama 3B 2023-12-08 1207 -604.18 41.14
47 andreaskoepf/llama2-13b-megacode2_min100 llama-2 13B 2023-08-14 1162 -571.36 56.92
48 Andron00e/YetAnother_Open-Llama-3B-LoRA llama 3B 2023-07-21 1163 -646.46
49 argilla/notus-7b-v1 mistral 7B 2023-11-16 7660 -564.14 60.22
50 Artples/L-MChat-Small phi 2B 2024-04-11 2793 -640.49 63.14
51 Artples/L-MChat-7b mistral 7B 2024-04-02 9634 -557.41 69.57
52 Aspik101/StableBeluga-13B-instruct-PL-lora_unload llama-2 13B 2023-08-04 1244 -544.64 56.24
53 Aspik101/vicuna-13b-v1.5-PL-lora_unload llama-2 13B 2023-08-03 1263 -566.80 55.24
54 Aspik101/WizardVicuna-Uncensored-3B-instruct-PL-lora_unload llama-2 3B 2023-08-07 1161 -626.59 39.95
55 AtAndDev/CapybaraMarcoroni-7B mistral 7B 2024-01-03 1162 -541.33 70.32
56 athirdpath/NSFW_DPO_Noromaid-7b mistral 7B 2023-12-12 1280 -546.50 61.59
57 augmxnt/shisa-base-7b-v1 mistral 7B 2023-11-19 1188 -648.62 51.64
58 augmxnt/shisa-gamma-7b-v1 mistral 7B 2023-12-23 151426 -590.25 55.50
59 augmxnt/shisa-7b-v1 Jain et al. (2023); Rafailov et al. (2024) mistral 7B 2023-11-27 1195 -643.26 55.01
60 Austism/chronos-hermes-13b-v2 llama-2 13B 2023-08-02 1225 -585.56 56.10
61 automerger/YamshadowExperiment28-7B mistral 7B 2024-03-18 3152 -595.65 76.86
62 Azazelle/Argetsu mistral 7B 2023-12-30 1208 -560.37 69.64
63 Azazelle/Dumb-Maidlet mistral 7B 2023-12-30 1201 -548.86 68.34
64 Azazelle/Half-NSFW_Noromaid-7b mistral 7B 2023-12-29 1208 -547.63 62.32
65 Azazelle/Maylin-7b mistral 7B 2024-01-04 1198 -575.94 70.26
66 Azazelle/Silicon-Medley mistral 7B 2023-12-29 1200 -566.66 69.49
67 Azazelle/SlimMelodicMaid mistral 7B 2023-12-30 1205 -577.28 69.70
68 Azazelle/smol_bruin-7b mistral 7B 2023-12-29 1207 -572.80 71.05
69 Azazelle/Tippy-Toppy-7b mistral 7B 2024-01-03 1197 -560.94 69.58
70 Azazelle/xDAN-SlimOrca mistral 7B 2023-12-29 1206 -575.20 68.04
71 Azazelle/Yuna-7b-Merge mistral 7B 2024-01-05 1201 -580.65 71.46
72 Azure99/blossom-v1-3b bloom 3B 2023-07-29 1228 -641.52 36.90
73 Azure99/blossom-v2-llama2-7b llama-2 7B 2023-09-06 1241 -572.17 51.71
74 Azure99/blossom-v2-3b bloom 3B 2023-08-08 1237 -654.44 35.98
75 Azure99/blossom-v3-mistral-7b mistral 7B 2023-11-20 1319 -565.31 62.95
76 Azure99/blossom-v3_1-mistral-7b mistral 7B 2023-11-27 1320 -567.17 62.53
77 Azure99/blossom-v4-mistral-7b mistral 7B 2023-12-26 1315 -550.37 63.61
78 BarryFutureman/NeuralTurdusVariant1-7B mistral 7B 2024-01-22 1170 -592.94 74.83
79 beaugogh/Llama2-7b-openorca-mc-v1 llama-2 7B 2023-08-20 1177 -613.56 52.24
80 beaugogh/Llama2-7b-openorca-mc-v2-dpo llama-2 7B 2023-10-06 1173 -608.74 52.32
81 beaugogh/Llama2-7b-sharegpt4 llama-2 7B 2023-08-12 1183 -660.22 51.05
82 beaugogh/pythia-1.4b-deduped-sharegpt gpt_neox 1B 2023-07-25 1293 -557.93 35.11
83 BEE-spoke-data/Mixtral-GQA-400m-v2 mixtral 2B 2023-12-20 1182 -794.25 28.45
84 beomi/KoAlpaca-KoRWKV-6B rwkv 6B 2023-06-02 2288 -1099.15 28.57
85 beomi/KoAlpaca-Polyglot-5.8B gpt_neox 6B 2023-03-16 4198 -1309.35 29.46
86 beomi/KoRWKV-6B rwkv 6B 2023-05-26 2127 -1148.78 28.19
87 beomi/llama-2-ko-7b Lee (2023) llama-2 6B 2023-07-20 5226 -907.04 45.32
88 beomi/Yi-Ko-6B Lee (2024b) llama 6B 2023-11-30 4261 -589.43 50.27
89 beowolx/CodeNinja-1.0-OpenChat-7B mistral 7B 2023-12-20 6300 -555.93 67.40
90 berkeley-nest/Starling-LM-7B-alpha Zhu et al. (2023b, a) mistral 7B 2023-11-25 18366 -571.63 67.05
91 bhavinjawade/SOLAR-10B-OrcaDPO-Jawade llama 10B 2024-01-06 1224 -563.19 74.27
92 bigcode/gpt_bigcode-santacoder gpt_bigcode 1B 2023-04-06 40198 -861.42 28.49
93 bigcode/starcoderbase-1b Shazeer (2019); Dao et al. (2022); Bavarian et al. (2022); Li et al. (2023a) gpt_bigcode 1B 2023-07-03 5247 -734.99 30.06
94 bigcode/starcoderbase-7b Shazeer (2019); Dao et al. (2022); Bavarian et al. (2022); Li et al. (2023a) gpt_bigcode 7B 2023-07-26 3566 -652.23 33.75
95 bigcode/starcoder2-3b Beltagy et al. (2020); Bavarian et al. (2022); Dao et al. (2022); Ainslie et al. (2023); Lozhkov et al. (2024) starcoder2 3B 2023-11-29 445316 -620.70 39.25
96 bigcode/starcoder2-7b Beltagy et al. (2020); Bavarian et al. (2022); Dao et al. (2022); Ainslie et al. (2023); Lozhkov et al. (2024) starcoder2 7B 2024-02-20 11834 -584.03 42.95
97 bigscience/bloomz-3b Muennighoff et al. (2023) bloom 3B 2022-10-08 7975 -701.42 37.03
98 bigscience/bloomz-7b1-mt Muennighoff et al. (2023) bloom 7B 2022-09-28 2427 -653.08 42.14
99 bigscience/bloomz-7b1 Muennighoff et al. (2023) bloom 7B 2022-09-27 12152 -652.34 42.21
100 bigscience/bloom-1b1 Shoeybi et al. (2020); Press et al. (2022); Dettmers et al. (2022) bloom 1B 2022-05-19 11054 -687.97 32.47
101 bigscience/bloom-1b7 Shoeybi et al. (2020); Press et al. (2022); Dettmers et al. (2022) bloom 1B 2022-05-19 40840 -657.05 33.98
102 bigscience/bloom-3b Shoeybi et al. (2020); Press et al. (2022); Dettmers et al. (2022) bloom 3B 2022-05-19 13914 -633.79 36.07
103 bigscience/bloom-7b1 Shoeybi et al. (2020); Press et al. (2022); Dettmers et al. (2022) bloom 7B 2022-05-19 21428 -604.20 39.18
104 BioMistral/BioMistral-7B-DARE Yadav et al. (2023a); Labrak et al. (2024); Yu et al. (2024a) mistral 7B 2024-02-05 1209 -586.37 57.03
105 BioMistral/BioMistral-7B Labrak et al. (2024) mistral 7B 2024-02-14 11349 -652.04 52.33
106 BlueNipples/TimeCrystal-l2-13B llama-2 13B 2023-11-11 1282 -567.72 59.26
107 bofenghuang/vigogne-2-13b-instruct llama-2 13B 2023-07-26 1203 -534.88 55.14
108 bofenghuang/vigogne-2-7b-chat llama-2 7B 2023-07-29 1170 -558.60 52.45
109 bofenghuang/vigogne-2-7b-instruct llama-2 7B 2023-07-20 1243 -555.11 52.02
110 bofenghuang/vigogne-7b-instruct llama-1 7B 2023-03-22 1167 -570.61 47.76
111 bofenghuang/vigostral-7b-chat mistral 7B 2023-09-29 4118 -535.15 59.18
112 BramVanroy/GEITje-7B-ultra Vanroy (2024) mistral 7B 2024-01-27 1597 -674.95 52.61
113 Brouz/Slerpeno llama 13B 2023-09-08 1201 -545.39 56.59
114 budecosystem/boomer-1b llama 1B 2023-10-03 1267 -846.13 28.44
115 CalderaAI/13B-BlueMethod llama-1 13B 2023-07-07 1193 -569.51 54.12
116 CalderaAI/13B-Legerdemain-L2 llama-2 13B 2023-08-03 1200 -551.69 55.13
117 CalderaAI/13B-Ouroboros llama 13B 2023-07-20 1197 -1133.56 49.54
118 CalderaAI/13B-Thorns-l2 llama 13B 2023-09-06 1207 -580.08 54.72
119 Cartinoe5930/Llama2_init_Mistral llama-2 7B 2024-01-16 1174 -528.82 60.98
120 castorini/rank_vicuna_7b_v1_fp16 Touvron et al. (2023b); Pradeep et al. (2023) llama-2 7B 2023-09-27 1165 -720.69 44.36
121 ceadar-ie/FinanceConnect-13B CeADAR (2023) llama 13B 2023-11-28 1270 -661.62 49.34
122 Changgil/K2S3-SOLAR-11b-v1.0 llama 10B 2024-03-03 2290 -811.23 36.67
123 Changgil/k2s3_test_24001 llama-2 13B 2024-02-14 2346 -561.61 56.67
124 chargoddard/storytime-13b llama-2 13B 2023-09-22 1166 -599.45 56.64
125 chickencaesar/llama2-platypus-llama2-chat-13B-hf llama-2 13B 2023-09-28 1216 -536.92 54.11
126 CHIH-HUNG/llama-2-13b-dolphin_20w llama-2 13B 2023-08-29 1187 -533.73 55.06
127 CHIH-HUNG/llama-2-13b-dolphin_5w llama-2 13B 2023-08-25 1175 -533.03 55.53
128 CHIH-HUNG/llama-2-13b-FINETUNE2_TEST_2.2w llama-2 13B 2023-09-04 1174 -530.90 53.20
129 CHIH-HUNG/llama-2-13b-FINETUNE3_3.3w-r16-gate_up_down llama-2 13B 2023-09-20 1171 -533.48 54.32
130 CHIH-HUNG/llama-2-13b-FINETUNE3_3.3w-r4-gate_up_down llama-2 13B 2023-09-19 1166 -535.81 53.35
131 CHIH-HUNG/llama-2-13b-FINETUNE3_3.3w-r4-q_k_v_o llama-2 13B 2023-09-19 1170 -536.50 53.62
132 CHIH-HUNG/llama-2-13b-FINETUNE3_3.3w-r8-gate_up_down llama-2 13B 2023-09-20 1165 -533.51 53.71
133 CHIH-HUNG/llama-2-13b-FINETUNE3_3.3w-r8-q_k_v_o_gate_up_down llama-2 13B 2023-09-24 1165 -536.24 53.43
134 CHIH-HUNG/llama-2-13b-FINETUNE3_3.3w-r8-q_k_v_o llama-2 13B 2023-09-20 1170 -535.00 53.99
135 CHIH-HUNG/llama-2-13b-FINETUNE4_addto15k_4.5w-r16-gate_up_down llama-2 13B 2023-10-08 1161 -531.46 54.88
136 CHIH-HUNG/llama-2-13b-FINETUNE4_compare15k_4.5w-r16-gate_up_down llama-2 13B 2023-10-08 1169 -532.00 53.94
137 CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r16-gate_up_down-test1 llama-2 13B 2023-10-07 1165 -531.27 53.66
138 CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r16-gate_up_down llama-2 13B 2023-09-22 1165 -533.56 53.52
139 CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r16-q_k_v_o llama-2 13B 2023-09-22 1164 -534.36 53.68
140 CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r4-gate_up_down llama-2 13B 2023-09-21 1174 -537.05 53.48
141 CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r4-q_k_v_o_gate_up_down llama-2 13B 2023-09-25 1163 -540.47 53.23
142 CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r8-gate_up_down llama-2 13B 2023-09-21 1171 -533.65 53.58
143 CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r8-q_k_v_o_gate_up_down llama-2 13B 2023-09-25 1165 -537.12 52.88
144 CHIH-HUNG/llama-2-13b-FINETUNE4_3.8w-r8-q_k_v_o llama-2 13B 2023-09-21 1171 -533.89 54.06
145 CHIH-HUNG/llama-2-13b-FINETUNE5_4w-r16-gate_up_down llama-2 13B 2023-10-04 1167 -534.88 53.44
146 CHIH-HUNG/llama-2-13b-FINETUNE5_4w-r16-q_k_v_o llama-2 13B 2023-10-04 1161 -533.84 54.63
147 CHIH-HUNG/llama-2-13b-FINETUNE5_4w-r4-q_k_v_o_gate_up_down llama-2 13B 2023-10-02 1163 -541.64 53.38
148 CHIH-HUNG/llama-2-13b-FINETUNE5_4w-r4-q_k_v_o llama-2 13B 2023-10-01 1172 -536.02 53.32
149 CHIH-HUNG/llama-2-13b-FINETUNE5_4w-r8-gate_up_down llama-2 13B 2023-10-03 1163 -534.71 54.02
150 CHIH-HUNG/llama-2-13b-FINETUNE5_4w-r8-q_k_v_o_gate_up_down llama-2 13B 2023-10-04 1164 -540.68 53.69
151 CHIH-HUNG/llama-2-13b-FINETUNE5_4w-r8-q_k_v_o llama-2 13B 2023-10-02 1167 -534.88 54.64
152 CHIH-HUNG/llama-2-13b-OpenOrca_5w llama-2 13B 2023-08-24 1178 -533.15 55.80
153 CHIH-HUNG/llama-2-13b-Open_Platypus_and_ccp_2.6w-3_epoch llama-2 13B 2023-09-05 1165 -533.19 53.80
154 chinoll/Yi-6b-200k-dpo llama 6B 2023-12-01 1187 -573.79 51.93
155 circulus/Llama-2-13b-orca-v1 llama-2 13B 2023-08-01 1228 -540.82 57.05
156 circulus/Llama-2-7b-orca-v1 llama-2 7B 2023-08-01 1235 -565.19 53.56
157 clibrain/Llama-2-7b-ft-instruct-es llama-2 7B 2023-08-09 2062 -559.81 49.63
158 CobraMamba/mamba-gpt-3b-v4 chiliu (2023) llama 3B 2023-09-05 1183 -605.83 41.24
159 codellama/CodeLlama-13b-hf Rozière et al. (2024) llama-2 13B 2023-08-24 8432 -582.07 43.35
160 codellama/CodeLlama-13b-Instruct-hf Rozière et al. (2024) llama-2 13B 2023-08-24 25248 -579.86 45.82
161 codellama/CodeLlama-13b-Python-hf Rozière et al. (2024) llama-2 13B 2023-08-24 2537 -587.68 37.00
162 codellama/CodeLlama-7b-hf Rozière et al. (2024) llama-2 6B 2023-08-24 66040 -597.05 39.81
163 codellama/CodeLlama-7b-Instruct-hf Rozière et al. (2024) llama-2 6B 2023-08-24 133523 -598.06 40.05
164 codellama/CodeLlama-7b-Python-hf Rozière et al. (2024) llama-2 6B 2023-08-24 5138 -605.60 36.42
165 cognitivecomputations/dolphin-2.2.1-mistral-7b mistral 7B 2023-10-30 7292 -546.22 65.01
166 cognitivecomputations/dolphin-2.6-mistral-7b-dpo-laser Sharma et al. (2023) mistral 7B 2024-01-01 2070 -558.72 67.28
167 cognitivecomputations/dolphin-2.6-mistral-7b-dpo mistral 7B 2023-12-31 1236 -565.34 67.20
168 cognitivecomputations/dolphin-2.6-mistral-7b mistral 7B 2023-12-27 1418 -559.17 64.92
169 cognitivecomputations/dolphin-2.9-llama3-8b llama-3 8B 2024-04-20 130977 -640.39 65.92
170 cognitivecomputations/dolphin-2.9.1-llama-3-8b llama-3 8B 2024-05-10 7575 -647.01 66.23
171 cognitivecomputations/dolphin-2.9.1-yi-1.5-9b llama 8B 2024-05-18 4880 -598.79 68.92
172 cognitivecomputations/Llama-3-8B-Instruct-abliterated-v2 llama-3 8B 2024-05-09 8728 -581.93 66.00
173 cognitivecomputations/openchat-3.5-0106-laser mistral 7B 2024-01-27 6155 -554.46 69.46
174 cognitivecomputations/TinyDolphin-2.8-1.1b llama 1B 2024-01-21 1678 -683.06 36.34
175 cognitivecomputations/WestLake-7B-v2-laser mistral 7B 2024-01-26 3980 -591.19 74.78
176 ContextualAI/archangel_sft-kto_llama13b Ethayarajh et al. (2023) llama 13B 2023-12-03 1214 -542.49 52.87
177 cookinai/BruinHermes mistral 7B 2023-12-17 1193 -551.19 73.42
178 cookinai/CatMacaroni14 mistral 7B 2023-12-31 1186 -552.22 72.68
179 Corianas/gpt-j-6B-Dolly gptj 6B 2023-03-28 1173 -502.52 39.60
180 crumb/apricot-wildflower-20 mistral 7B 2023-12-19 1179 -531.99 59.74
181 cyberagent/open-calm-7b Andonian et al. (2021) gpt_neox 7B 2023-05-15 24168 -1002.50 28.21
182 cypienai/cymist2-v01-SFT Lacoste et al. (2019) mistral 7B 2024-05-12 2754 -628.53 51.71
183 cypienai/cymist-2-v02-SFT Lacoste et al. (2019) mistral 7B 2024-05-22 2717 -537.73 62.57
184 Dampish/Dante-2.8B gpt_neox 2B 2023-05-09 1237 -671.96
185 Dampish/StellarX-4B-V0.2 Black et al. (2022) gpt_neox 4B 2023-06-03 1245 -866.96 36.15
186 Dampish/StellarX-4B-V0 Black et al. (2022) gpt_neox 4B 2023-05-27 1268 -722.64 37.31
187 Danielbrdz/Barcenas-Llama3-8b-ORPO llama-3 8B 2024-04-29 12814 -557.50 72.50
188 Danielbrdz/Barcenas-13b llama-2 13B 2023-09-09 1178 -558.81 55.83
189 Danielbrdz/Barcenas-3b llama-2 3B 2023-11-15 1167 -554.05 41.74
190 Danielbrdz/Barcenas-7b llama 7B 2023-08-25 1175 -603.96 50.87
191 Danielbrdz/CodeBarcenas-7b llama-2 7B 2023-09-03 1174 -656.68 40.09
192 danielpark/gorani-100k-llama2-13b-instruct llama-2 13B 2023-10-04 1176 -1410.79 29.69
193 databricks/dolly-v2-12b Conover et al. (2023) gpt_neox 12B 2023-04-11 3860 -566.95 39.46
194 databricks/dolly-v2-3b Conover et al. (2023) gpt_neox 3B 2023-04-13 21443 -556.97
195 databricks/dolly-v2-7b Conover et al. (2023) gpt_neox 7B 2023-04-13 10069 -555.60 39.24
196 DBCMLAB/Llama-3-instruction-constructionsafety-layertuning AI@Meta (2024); Lee (2024a); Lee and Ahn (2024) llama-3 8B 2024-05-22 2712 -658.58 56.32
197 Deathsquad10/TinyLlama-Remix llama 1B 2023-11-26 1213 -812.12 34.00
198 Deathsquad10/TinyLlama-repeat llama 1B 2024-01-06 1191 -622.61 37.09
199 Deathsquad10/TinyLlama-1.1B-Remix-V.2 llama 1B 2024-01-05 1192 -670.53 34.91
200 Deci/DeciLM-7B-instruct DeciAI Research Team (2023) deci 7B 2023-12-10 9769 -570.53 63.19
201 DeepMount00/Llama-3-8b-Ita llama-3 8B 2024-05-01 179401 -578.66 73.65
202 DeepMount00/Mistral-Ita-7b mistral 7B 2023-11-08 2927 -579.30 58.26
203 deepseek-ai/DeepSeek-Coder-V2-Lite-Base Dai et al. (2024) deepseek 16B 2024-06-14 13642 -527.73
204 deepseek-ai/DeepSeek-Coder-V2-Lite-Instruct Dai et al. (2024) deepseek 16B 2024-06-14 142758 -569.70
205 deepseek-ai/deepseek-coder-1.3b-base Guo et al. (2024) deepseek 1B 2023-10-28 48960 -718.82
206 deepseek-ai/deepseek-coder-1.3b-instruct Guo et al. (2024) deepseek 1B 2023-10-29 28984 -785.33 32.40
207 deepseek-ai/deepseek-coder-6.7b-base Guo et al. (2024) deepseek 6B 2023-10-23 38348 -647.51 40.87
208 deepseek-ai/deepseek-coder-6.7b-instruct Guo et al. (2024) deepseek 6B 2023-10-29 28828 -686.56 43.57
209 deepseek-ai/deepseek-coder-7b-base-v1.5 Guo et al. (2024) deepseek 7B 2024-01-25 895 -570.98
210 deepseek-ai/deepseek-coder-7b-instruct-v1.5 Guo et al. (2024) deepseek 6B 2024-01-25 28928 -604.47 50.89
211 deepseek-ai/deepseek-llm-7b-base DeepSeek-AI et al. (2024a) deepseek 7B 2023-11-29 19180 -552.39
212 deepseek-ai/deepseek-llm-7b-chat DeepSeek-AI et al. (2024a) deepseek 7B 2023-11-29 34271 -594.23 59.38
213 deepseek-ai/deepseek-math-7b-base Shao et al. (2024) deepseek 7B 2024-02-05 37614 -566.93 57.61
214 deepseek-ai/deepseek-math-7b-instruct Shao et al. (2024) deepseek 7B 2024-02-05 8352 -593.77 51.48
215 deepseek-ai/deepseek-math-7b-rl Shao et al. (2024) deepseek 6B 2024-02-05 1633 -607.25 49.54
216 deepseek-ai/deepseek-moe-16b-base Dai et al. (2024) deepseek 16B 2024-01-08 13924 -544.49
217 deepseek-ai/deepseek-moe-16b-chat Dai et al. (2024) deepseek 16B 2024-01-09 8852 -578.23
218 deepseek-ai/DeepSeek-Prover-V1.5-Base Xin et al. (2024b) deepseek 7B 2024-08-15 230 -554.13
219 deepseek-ai/DeepSeek-Prover-V1.5-RL Xin et al. (2024b) deepseek 7B 2024-08-15 12223 -672.91
220 deepseek-ai/DeepSeek-Prover-V1.5-SFT Xin et al. (2024b) deepseek 7B 2024-08-15 6459 -670.56
221 deepseek-ai/DeepSeek-Prover-V1 Xin et al. (2024a) deepseek 7B 2024-08-16 114 -791.21
222 deepseek-ai/DeepSeek-R1-Distill-Llama-8B DeepSeek-AI et al. (2025) deepseek 8B 2025-01-20 199313 -693.69
223 deepseek-ai/DeepSeek-R1-Distill-Qwen-14B DeepSeek-AI et al. (2025) deepseek 14B 2025-01-20 139489 -605.82
224 deepseek-ai/DeepSeek-R1-Distill-Qwen-1.5B DeepSeek-AI et al. (2025) deepseek 2B 2025-01-20 290094 -854.97
225 deepseek-ai/DeepSeek-R1-Distill-Qwen-7B DeepSeek-AI et al. (2025) deepseek 7B 2025-01-20 201118 -758.13
226 deepseek-ai/DeepSeek-V2-Lite-Chat DeepSeek-AI et al. (2024b) deepseek 16B 2024-05-15 18056 -588.92
227 deepseek-ai/DeepSeek-V2-Lite DeepSeek-AI et al. (2024b) deepseek 16B 2024-05-15 25474 -533.54
228 deepseek-ai/ESFT-vanilla-lite Wang et al. (2024c) deepseek 16B 2024-07-04 270 -550.46
229 Delcos/Mistral-Pygmalion-7b llama-2 7B 2023-10-09 1185 -556.74 51.02
230 dfurman/Llama-3-8B-Orpo-v0.1 llama-3 8B 2024-04-26 5146 -515.13 64.67
231 dhmeltzer/Llama-2-13b-hf-ds_eli5_1024_r_64_alpha_16_merged llama-2 13B 2023-09-14 1188 -542.35 54.16
232 dhmeltzer/Llama-2-13b-hf-ds_wiki_1024_full_r_64_alpha_16_merged llama-2 13B 2023-09-14 1198 -536.93 52.94
233 dhmeltzer/Llama-2-13b-hf-eli5-wiki-1024_r_64_alpha_16_merged llama-2 13B 2023-09-14 1192 -538.52 53.57
234 dhmeltzer/llama-7b-SFT_eli5_wiki65k_1024_r_64_alpha_16_merged llama 6B 2023-08-25 1257 -568.38 50.00
235 diffnamehard/Psyfighter2-Noromaid-ties-13B llama 13B 2023-12-28 1215 -569.13 59.47
236 digitous/Alpacino13b llama-1 13B 2023-04-13 1182 -547.35 52.39
237 digitous/13B-Chimera llama-1 13B 2023-05-23 1359 -562.36 54.92
238 Doctor-Shotgun/CalliopeDS-L2-13B Yadav et al. (2023a) llama-2 13B 2023-09-16 1428 -568.72 56.34
239 Doctor-Shotgun/CalliopeDS-v2-L2-13B llama-2 13B 2023-09-28 1250 -572.65 57.12
240 DopeorNope/You_can_cry_Snowman-13B llama 13B 2023-12-27 1203 -577.68 69.46
241 dotvignesh/perry-7b llama 7B 2023-09-28 1171 -618.90 49.55
242 dreamgen/WizardLM-2-7B Luo et al. (2023); Xu et al. (2023a); Luo et al. (2025) mistral 7B 2024-04-16 2533 -641.93 63.75
243 dvruette/oasst-pythia-12b-flash-attn-5000-steps gpt_neox 12B 2023-03-12 1175 -602.37 40.73
244 dvruette/oasst-pythia-12b-pretrained-sft gpt_neox 12B 2023-04-03 1170 -552.58 41.48
245 dvruette/oasst-pythia-12b-reference gpt_neox 12B 2023-04-03 1174 -557.88 40.33
246 eachadea/vicuna-13b-1.1 llama-1 13B 2023-04-13 1336 -646.61 53.29
247 eachadea/vicuna-7b-1.1 llama-1 7B 2023-04-13 1388 -613.88 50.37
248 eldogbbhed/Peagle-9b mistral 8B 2024-03-10 10180 -581.92 73.30
249 EleutherAI/gpt-j-6b Gao et al. (2020); Wang (2021); Wang and Komatsuzaki (2021); Su et al. (2023a) gptj 6B 2022-03-02 241435 -479.14 40.10
250 EleutherAI/gpt-neo-1.3B Gao et al. (2020); Black et al. (2021) gpt_neo 1B 2022-03-02 243332 -545.00 33.58
251 EleutherAI/gpt-neo-2.7B Gao et al. (2020); Black et al. (2021) gpt_neo 2B 2022-03-02 205503 -520.30 36.20
252 EleutherAI/llemma_7b Azerbayev et al. (2024) llama-2 7B 2023-09-12 4671 -554.43 48.75
253 EleutherAI/polyglot-ko-12.8b Kim et al. (2022); Ko et al. (2023); Su et al. (2023a) gpt_neox 13B 2022-10-14 2941 -967.65 33.33
254 EleutherAI/pythia-1b-deduped Gao et al. (2020); Biderman et al. (2022, 2023) gpt_neox 1B 2023-02-14 17983 -551.07 32.78
255 EleutherAI/pythia-12b-deduped Gao et al. (2020); Biderman et al. (2022, 2023) gpt_neox 12B 2023-02-27 9737 -477.81 39.70
256 EleutherAI/pythia-12b Gao et al. (2020); Biderman et al. (2022, 2023) gpt_neox 12B 2023-02-28 9192 -475.42 38.82
257 EleutherAI/pythia-1.4b-deduped Gao et al. (2020); Biderman et al. (2022, 2023) gpt_neox 1B 2023-02-09 12730 -534.52 35.00
258 EleutherAI/pythia-1.4b Gao et al. (2020); Biderman et al. (2022, 2023) gpt_neox 1B 2023-02-09 22152 -531.68 34.75
259 EleutherAI/pythia-2.8b-deduped Gao et al. (2020); Biderman et al. (2022, 2023) gpt_neox 2B 2023-02-10 11649 -507.42 36.72
260 EleutherAI/pythia-6.9b-deduped Gao et al. (2020); Biderman et al. (2022, 2023) gpt_neox 6B 2023-02-25 9824 -490.70 39.30
261 elinas/chronos-13b-v2 llama-2 13B 2023-08-02 1944 -595.38 55.25
262 elliotthwang/elliott_Llama-2-7b-hf llama-2 6B 2023-10-09 1181 -553.66 50.20
263 elyza/ELYZA-japanese-Llama-2-13b-instruct Touvron et al. (2023b); Sasaki et al. (2023b) llama-2 13B 2023-12-25 1710 -594.98 54.72
264 elyza/ELYZA-japanese-Llama-2-13b Touvron et al. (2023b); Sasaki et al. (2023b) llama-2 13B 2023-12-25 1200 -587.25 56.14
265 elyza/ELYZA-japanese-Llama-2-7b-fast-instruct Touvron et al. (2023b); Sasaki et al. (2023a) llama-2 7B 2023-08-28 2387 -629.72 49.15
266 elyza/ELYZA-japanese-Llama-2-7b-fast Touvron et al. (2023b); Sasaki et al. (2023a) llama-2 7B 2023-08-28 1766 -640.71 47.67
267 elyza/ELYZA-japanese-Llama-2-7b-instruct Touvron et al. (2023b); Sasaki et al. (2023a) llama-2 7B 2023-08-28 6296 -625.45 49.78
268 elyza/ELYZA-japanese-Llama-2-7b Touvron et al. (2023b); Sasaki et al. (2023a) llama-2 7B 2023-08-28 2356 -630.66 48.70
269 Enno-Ai/ennodata-raw-pankajmathur-13b-peft llama 13B 2023-09-29 1215 -613.07 55.40
270 ericzzz/falcon-rw-1b-chat falcon 1B 2023-12-05 1319 -722.18 37.37
271 ericzzz/falcon-rw-1b-instruct-openorca falcon 1B 2023-11-24 1585 -752.98 37.63
272 euclaise/falcon_1b_stage1 falcon 1B 2023-09-15 2119 -734.28 37.25
273 euclaise/falcon_1b_stage2 falcon 1B 2023-09-17 4016 -717.40 37.59
274 euclaise/Ferret_7B mistral 7B 2023-10-28 1172 -634.24 53.87
275 Eurdem/Defne_llama3_2x8B mixtral 13B 2024-05-10 9076 -571.19 71.73
276 Expert68/llama2_13b_instructed_version2 llama-2 13B 2023-10-14 1184 -561.35 55.41
277 facebook/opt-iml-max-1.3b Iyer et al. (2023) opt 1B 2023-01-26 8604 -712.07 35.21
278 facebook/opt-13b Brown et al. (2020); Zhang et al. (2022b) opt 13B 2022-05-11 19130 -625.38 40.06
279 facebook/opt-1.3b Brown et al. (2020); Zhang et al. (2022b) opt 1B 2022-05-11 139758 -700.64 34.60
280 facebook/opt-2.7b Brown et al. (2020); Zhang et al. (2022b) opt 2B 2022-05-11 61450 -673.03 36.74
281 facebook/opt-6.7b Brown et al. (2020); Zhang et al. (2022b) opt 6B 2022-05-11 44909 -641.13 39.08
282 facebook/xglm-1.7B Lin et al. (2022b) xglm 1B 2022-03-02 2388 -723.91 31.42
283 facebook/xglm-4.5B Lin et al. (2022b) xglm 5B 2022-03-02 1436 -669.92 34.31
284 facebook/xglm-7.5B Lin et al. (2022b) xglm 7B 2022-03-02 4654 -663.87 36.38
285 failspy/Meta-Llama-3-8B-Instruct-abliterated-v3 llama-3 8B 2024-05-20 9975 -572.33 67.27
286 failspy/Phi-3-medium-4k-instruct-abliterated-v3 phi3 13B 2024-05-22 5430 -569.05 70.12
287 FairMind/Llama-3-8B-4bit-UltraChat-Ita llama-3 8B 2024-05-03 5014 -543.07 61.54
288 FairMind/Phi-3-mini-4k-instruct-bnb-4bit-Ita mistral 4B 2024-05-02 2746 -634.71 66.61
289 fblgit/LUNA-SOLARkrautLM-Instruct llama 10B 2023-12-22 1180 -582.98 73.79
290 fblgit/una-cybertron-7b-v2-bf16 Murias (2023) mistral 7B 2023-12-02 1440 -568.27 69.67
291 fblgit/UNA-POLAR-10.7B-InstructMath-v2 llama 10B 2024-01-02 1180 -560.34 74.07
292 feidfoe/Metamath-reproduce-7b llama-2 7B 2023-11-24 1198 -724.03 55.81
293 FelixChao/CodeLlama13B-Finetune-v1 llama 13B 2023-09-13 1188 -647.40 47.19
294 FelixChao/llama2-13b-math1.1 llama-2 13B 2023-08-12 1181 -602.88 54.18
295 FelixChao/llama2-13b-math1.2 llama-2 13B 2023-08-15 1197 -607.07 54.19
296 FinancialSupport/saiga-7b mistral 7B 2023-12-28 3902 -558.71 64.51
297 fireballoon/baichuan-vicuna-chinese-7b llama-1 7B 2023-06-18 1210 -702.08 46.06
298 FlagAlpha/Llama2-Chinese-7b-Chat llama-2 7B 2023-07-23 1383 -627.19 51.13
299 FlagAlpha/Llama3-Chinese-8B-Instruct llama-3 8B 2024-04-23 1978 -557.17 63.50
300 Fredithefish/Guanaco-3B-Uncensored-v2 gpt_neox 2B 2023-08-27 2189 -632.46 38.98
301 FreedomIntelligence/AceGPT-7B llama 7B 2023-09-14 2982 -565.76 49.47
302 FreedomIntelligence/phoenix-inst-chat-7b bloom 7B 2023-04-11 1296 -693.44 43.03
303 freewheelin/free-llama3-dpo-v0.2 Kim et al. (2024b) llama-3 8B 2024-05-09 3896 -499.17 62.69
304 gagan3012/MetaModelv2 llama 10B 2024-01-03 1203 -560.16 74.24
305 gagan3012/MetaModelv3 llama 10B 2024-01-05 1207 -561.14 74.39
306 garage-bAInd/Platypus2-13B Hu et al. (2022); Touvron et al. (2023b); Lee et al. (2024) llama 13B 2023-08-05 3716 -533.92 54.89
307 garage-bAInd/Platypus2-7B Hu et al. (2022); Touvron et al. (2023b); Lee et al. (2024) llama 6B 2023-08-22 6198 -554.08 49.97
308 GeneZC/MiniChat-1.5-3B Jain et al. (2023); Zhang et al. (2024a); Rafailov et al. (2024) llama 3B 2023-11-26 1601 -599.51 50.23
309 GeneZC/MiniChat-2-3B Jain et al. (2023); Zhang et al. (2024a); Rafailov et al. (2024) llama 3B 2023-12-27 4167 -607.10 51.49
310 GeneZC/MiniChat-3B Zhang et al. (2024a) llama 3B 2023-11-11 1633 -585.38 45.31
311 GeneZC/MiniMA-2-3B Zhang et al. (2024a) llama 3B 2023-12-27 1492 -526.37 44.75
312 GeneZC/MiniMA-3B Zhang et al. (2024a) llama 3B 2023-11-11 1494 -537.57 41.44
313 georgesung/llama2_7b_chat_uncensored llama-2 7B 2023-07-20 2602 -568.34 49.67
314 ghost-x/ghost-7b-alpha mistral 7B 2024-04-13 4854 -662.59 57.65
315 glaiveai/glaive-coder-7b llama-2 7B 2023-09-17 1222 -692.31 41.56
316 google/codegemma-2b CodeGemma Team et al. (2024) gemma 2B 2024-03-21 8798 -793.92 32.19
317 google/codegemma-7b-it CodeGemma Team et al. (2024) gemma 8B 2024-03-21 11227 -1040.98 58.28
318 google/codegemma-7b CodeGemma Team et al. (2024) gemma 8B 2024-03-21 3779 -556.79 56.73
319 google/gemma-1.1-7b-it Joshi et al. (2017); Zhao et al. (2018); Mihaylov et al. (2018); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Bisk et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Zellers et al. (2019); Austin et al. (2021); Cobbe et al. (2021); Hendrycks et al. (2021a); Chen et al. (2021); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-03-26 19835 -1355.60 60.09
320 google/gemma-2b-it Joshi et al. (2017); Mihaylov et al. (2018); Rudinger et al. (2018); Zhao et al. (2018); Sap et al. (2019); Bisk et al. (2019); Zellers et al. (2019); Clark et al. (2019); Talmor et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Cobbe et al. (2021); Hendrycks et al. (2021a); Dhamala et al. (2021); Chen et al. (2021); Austin et al. (2021); Hartvigsen et al. (2022); Parrish et al. (2022); Lin et al. (2022a); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 2B 2024-02-08 103851 -886.70 42.75
321 google/gemma-2b Joshi et al. (2017); Mihaylov et al. (2018); Rudinger et al. (2018); Zhao et al. (2018); Sap et al. (2019); Bisk et al. (2019); Zellers et al. (2019); Clark et al. (2019); Talmor et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Cobbe et al. (2021); Hendrycks et al. (2021a); Dhamala et al. (2021); Chen et al. (2021); Austin et al. (2021); Hartvigsen et al. (2022); Parrish et al. (2022); Lin et al. (2022a); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 2B 2024-02-08 256798 -581.35 46.51
322 google/gemma-7b-it Joshi et al. (2017); Mihaylov et al. (2018); Rudinger et al. (2018); Zhao et al. (2018); Sap et al. (2019); Bisk et al. (2019); Zellers et al. (2019); Clark et al. (2019); Talmor et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Cobbe et al. (2021); Hendrycks et al. (2021a); Dhamala et al. (2021); Chen et al. (2021); Austin et al. (2021); Hartvigsen et al. (2022); Parrish et al. (2022); Lin et al. (2022a); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-02-13 66776 -1327.43 53.56
323 google/gemma-7b Joshi et al. (2017); Mihaylov et al. (2018); Rudinger et al. (2018); Zhao et al. (2018); Sap et al. (2019); Bisk et al. (2019); Zellers et al. (2019); Clark et al. (2019); Talmor et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Cobbe et al. (2021); Hendrycks et al. (2021a); Dhamala et al. (2021); Chen et al. (2021); Austin et al. (2021); Hartvigsen et al. (2022); Parrish et al. (2022); Lin et al. (2022a); Srivastava et al. (2023); Zhong et al. (2023); Dettmers et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-02-08 72047 -539.12 63.75
324 google/recurrentgemma-2b-it Joshi et al. (2017); Mihaylov et al. (2018); Rudinger et al. (2018); Zhao et al. (2018); Sap et al. (2019); Bisk et al. (2019); Zellers et al. (2019); Clark et al. (2019); Talmor et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Gehman et al. (2020); Cobbe et al. (2021); Hendrycks et al. (2021a); Dhamala et al. (2021); Chen et al. (2021); Hendrycks et al. (2021b); Austin et al. (2021); Hartvigsen et al. (2022); Parrish et al. (2022); Lin et al. (2022a); Srivastava et al. (2023); Zhong et al. (2023); De et al. (2024); Griffin Team et al. (2024) recurrent_gemma 2B 2024-04-08 4046 -1033.09 40.86
325 gradientai/Llama-3-8B-Instruct-Gradient-1048k Peng et al. (2023b); Ding et al. (2023); Pekelis et al. (2024); AI@Meta (2024); Liu et al. (2025) llama-3 8B 2024-04-29 6854 -564.98 59.84
326 gradientai/Llama-3-8B-Instruct-262k Peng et al. (2023b); Ding et al. (2023); Pekelis et al. (2024); AI@Meta (2024); Liu et al. (2025) llama-3 8B 2024-04-25 12536 -555.17 60.26
327 GritLM/GritLM-7B Muennighoff et al. (2024) mistral 7B 2024-02-11 72991 -564.58 61.41
328 Gryphe/MythoLogic-L2-13b llama 13B 2023-08-03 1198 -570.07 56.19
329 Gryphe/MythoMax-L2-13b llama 13B 2023-08-10 14308 -583.80 56.00
330 Gryphe/MythoMix-L2-13b llama 13B 2023-08-08 1174 -566.44 56.31
331 guardrail/llama-2-7b-guanaco-instruct-sharded llama-2 6B 2023-07-21 1320 -647.09 50.58
332 gywy/llama2-13b-chinese-v2 llama-2 13B 2023-08-22 1163 -719.40 49.58
333 habanoz/tinyllama-oasst1-top1-instruct-full-lr1-5-v0.1 llama 1B 2023-11-19 1178 -679.96 35.58
334 habanoz/TinyLlama-1.1B-intermediate-step-715k-1.5T-lr-5-1epch-airoboros3.1-1k-instruct-V1 llama 1B 2023-11-22 1279 -658.65 34.98
335 habanoz/TinyLlama-1.1B-intermediate-step-715k-1.5T-lr-5-2.2epochs-oasst1-top1-instruct-V1 llama 1B 2023-11-20 1283 -648.88 35.45
336 habanoz/TinyLlama-1.1B-intermediate-step-715k-1.5T-lr-5-3epochs-oasst1-top1-instruct-V1 llama 1B 2023-11-21 1277 -651.34 35.42
337 habanoz/TinyLlama-1.1B-intermediate-step-715k-1.5T-lr-5-4epochs-oasst1-top1-instruct-V1 llama 1B 2023-11-21 1288 -646.55 35.28
338 hakurei/instruct-12b gpt_neox 12B 2023-04-09 1177 -602.56 38.63
339 hakurei/mommygpt-3B llama 3B 2023-11-12 1177 -598.84 41.36
340 haoranxu/ALMA-13B-Pretrain Xu et al. (2024b, a) llama 13B 2023-09-17 5510 -558.03 51.68
341 haoranxu/ALMA-13B Xu et al. (2024b, a) llama 13B 2023-09-17 2012 -563.34 50.16
342 haoranxu/ALMA-7B Xu et al. (2024b, a) llama 7B 2023-09-17 1292 -599.26 45.32
343 harborwater/open-llama-3b-claude-30k llama 3B 2023-11-21 1200 -603.17 40.93
344 health360/Healix-1.1B-V1-Chat-dDPO llama 1B 2023-11-05 2790 -796.10 33.00
345 heegyu/LIMA2-13b-hf Touvron et al. (2023b) llama-2 13B 2023-08-07 3330 -596.33 52.98
346 heegyu/LIMA2-7b-hf Touvron et al. (2023b) llama-2 7B 2023-08-04 3463 -651.31 49.27
347 heegyu/LIMA-13b-hf llama 13B 2023-08-01 3325 -544.21 52.61
348 heegyu/RedTulu-Uncensored-3B-0719 gpt_neox 3B 2023-07-23 1187 -703.96 39.19
349 heegyu/WizardVicuna2-13b-hf Touvron et al. (2023b) llama-2 13B 2023-08-07 6481 -612.15 51.05
350 heegyu/WizardVicuna-open-llama-3b-v2 llama 3B 2023-08-25 9364 -671.17 38.77
351 heegyu/WizardVicuna-Uncensored-3B-0719 llama 3B 2023-07-23 1168 -638.46 39.73
352 heegyu/WizardVicuna-3B-0719 llama 3B 2023-07-23 3356 -655.31 39.48
353 HenryJJ/Instruct_Mistral-7B-v0.1_Dolly15K mistral 7B 2024-01-02 1166 -577.98 60.45
354 HenryJJ/Instruct_Yi-6B_Dolly15K llama 6B 2024-01-06 1170 -524.87 56.85
355 HenryJJ/Instruct_Yi-6B_Dolly_CodeAlpaca llama 6B 2024-01-07 1177 -527.06 56.11
356 hfl/chinese-alpaca-2-13b llama 13B 2023-08-14 1308 -592.75 57.41
357 hfl/chinese-llama-2-13b llama-2 13B 2023-08-11 1176 -665.69 52.00
358 hfl/chinese-llama-2-1.3b llama-2 1B 2023-10-08 2269 -1132.44 28.59
359 HiTZ/GoLLIE-7B Sainz et al. (2024) llama-2 7B 2023-09-25 1394 -632.52 37.48
360 hiyouga/Baichuan2-7B-Base-LLaMAfied llama-2 7B 2023-09-08 1218 -530.80 48.99
361 hiyouga/Baichuan2-7B-Chat-LLaMAfied llama-2 7B 2023-09-09 1211 -609.32 51.42
362 hoskinson-center/proofGPT-v0.1-6.7B gpt_neox 6B 2023-02-04 1187 -974.98 29.72
363 hpcai-tech/Colossal-LLaMA-2-7b-base Li et al. (2023b); Touvron et al. (2023b); Dao (2023) llama-2 7B 2023-09-18 1192 -690.99 51.39
364 HuggingFaceFW/ablation-model-fineweb-v1 Lacoste et al. (2019) llama 1B 2024-04-20 2238 -771.42 36.76
365 HuggingFaceH4/mistral-7b-sft-beta mistral 7B 2023-10-26 7408 -553.64 59.78
366 HuggingFaceH4/zephyr-7b-alpha Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistral 7B 2023-10-09 12911 -559.73 59.50
367 HuggingFaceH4/zephyr-7b-beta Ding et al. (2023); Tunstall et al. (2023); Cui et al. (2024); Rafailov et al. (2024) mistral 7B 2023-10-26 294019 -570.40 59.08
368 huggyllama/llama-13b llama-1 13B 2023-04-03 6345 -541.90 51.33
369 huggyllama/llama-7b llama-1 6B 2023-04-03 192383 -562.04 46.37
370 HWERI/Llama2-7b-sharegpt4 llama-2 7B 2023-08-04 1200 -660.22 51.05
371 HWERI/pythia-1.4b-deduped-sharegpt gpt_neox 1B 2023-08-10 1200 -557.93 35.11
372 hyunseoki/ko-en-llama2-13b llama-2 13B 2023-10-02 3351 -550.60 51.27
373 hyunseoki/ko-ref-llama2-13b llama-2 13B 2023-10-04 3364 -775.26 43.62
374 hyunseoki/ko-ref-llama2-7b llama-2 7B 2023-10-04 3312 -851.38 40.75
375 hywu/Camelidae-8x13B Houlsby et al. (2019); Komatsuzaki et al. (2023); Dettmers et al. (2023); Wu et al. (2024) camelidae 13B 2024-01-10 1886 -535.46 59.40
376 hywu/Camelidae-8x7B Houlsby et al. (2019); Komatsuzaki et al. (2023); Dettmers et al. (2023); Wu et al. (2024) camelidae 7B 2024-01-10 1901 -565.23 54.47
377 h2oai/h2ogpt-gm-oasst1-en-1024-12b gpt_neox 12B 2023-05-02 1185 -495.39 40.65
378 h2oai/h2ogpt-gm-oasst1-en-2048-open-llama-7b-preview-300bt-v2 llama-1 7B 2023-05-10 1190 -746.41 37.55
379 h2oai/h2ogpt-gm-oasst1-en-2048-open-llama-7b-preview-300bt llama-1 7B 2023-05-04 1185 -1042.98 34.32
380 h2oai/h2ogpt-oasst1-512-12b gpt_neox 12B 2023-04-17 1319 -484.68 40.48
381 h2oai/h2ogpt-oig-oasst1-256-6_9b gpt_neox 9B 2023-04-17 1191 -492.94 38.62
382 h2oai/h2ogpt-oig-oasst1-512-6_9b gpt_neox 9B 2023-04-18 1749 -521.42 38.52
383 ibivibiv/llama-3-nectar-dpo-8B Lacoste et al. (2019) llama-3 8B 2024-05-14 6085 -576.43 67.92
384 ibm/merlinite-7b Sudalairaj et al. (2024) mistral 7B 2024-03-02 11156 -541.12 64.00
385 ibndias/NeuralHermes-MoE-2x7B mixtral 12B 2024-01-03 1172 -533.10 64.08
386 ibranze/araproje-llama2-7b-hf llama-2 7B 2023-10-06 1161 -549.87 49.73
387 IDEA-CCNL/Ziya-LLaMA-13B-Pretrain-v1 IDEA-CCNL (2021); Zhang et al. (2022a); Yang et al. (2022) llama-1 13B 2023-06-01 1171 -1397.30 29.96
388 IDEA-CCNL/Ziya-LLaMA-13B-v1 IDEA-CCNL (2021); Zhang et al. (2022a); Yang et al. (2022) llama-1 13B 2023-05-16 1244 -1397.34 29.82
389 ignos/LeoScorpius-GreenNode-Alpaca-7B-v1 mistral 7B 2023-12-15 1206 -558.75 74.74
390 ignos/LeoScorpius-GreenNode-Platypus-7B-v1 mistral 7B 2023-12-15 1188 -540.87 68.96
391 ignos/Mistral-T5-7B-v1 mistral 7B 2023-12-18 1266 -557.24 72.47
392 ikala/bloom-zh-3b-chat bloom 3B 2023-05-07 1352 -722.56 37.58
393 IkariDev/Athena-v1 llama 13B 2023-08-30 1225 -566.12 54.11
394 IkariDev/Athena-v4 llama 13B 2023-10-07 1214 -555.80 57.23
395 INSAIT-Institute/BgGPT-7B-Instruct-v0.2 mistral 7B 2024-03-03 3265 -594.56 63.08
396 Intel/neural-chat-7b-v3-1 Mukherjee et al. (2023) mistral 7B 2023-11-14 3976 -563.71 59.90
397 Intel/neural-chat-7b-v3-2 Yu et al. (2024b) mistral 7B 2023-11-21 2102 -554.92 68.29
398 Intel/neural-chat-7b-v3-3 Yu et al. (2024b) mistral 7B 2023-12-09 166226 -575.02 69.83
399 invalid-coder/Sakura-SOLAR-Instruct-CarbonVillain-en-10.7B-v2-slerp llama 10B 2024-01-10 12810 -560.73 74.45
400 itsliupeng/llama2_7b_code llama-2 7B 2023-09-28 1236 -538.58 49.05
401 itsliupeng/openllama-7b-base llama 7B 2023-12-08 1200 -536.60 47.09
402 itsliupeng/openllama-7b-icl Shi et al. (2024) llama 7B 2023-12-08 1186 -535.25 47.93
403 jae24/openhermes_dpo_norobot_0201 mistral 7B 2024-01-02 1161 -562.15 63.78
404 jan-hq/trinity-v1 mistral 7B 2023-12-14 1192 -559.84 74.80
405 Jayant9928/orpo_med_v3 llama 8B 2024-05-01 2705 -557.53 62.21
406 jeonsworld/CarbonVillain-en-10.7B-v1 llama 10B 2023-12-28 1164 -561.74 74.28
407 jeonsworld/CarbonVillain-en-10.7B-v4 llama 10B 2023-12-30 12898 -561.44 74.52
408 jerryjalapeno/nart-100k-7b llama-1 7B 2023-07-14 1194 -585.01 46.39
409 Jiayi-Pan/Tiny-Vicuna-1B llama 1B 2023-11-22 3094 -647.57 34.76
410 jingyeom/freeze_KoSoLAR-10.7B-v0.2_1.4_dedup llama 10B 2024-01-29 2287 -594.17 60.06
411 johnsnowlabs/BioLing-7B-Dare mistral 7B 2024-04-08 2682 -591.07 60.32
412 johnsnowlabs/JSL-MedLlama-3-8B-v2.0 llama-3 8B 2024-04-30 11813 -569.39 61.93
413 jondurbin/airoboros-gpt-3.5-turbo-100k-7b llama-1 7B 2023-05-12 1473 -614.18 47.05
414 jondurbin/airoboros-l2-13b-gpt4-m2.0 llama 13B 2023-07-28 1378 -614.41 52.66
415 jondurbin/airoboros-l2-13b-gpt4-2.0 llama 13B 2023-07-27 1395 -575.42 52.49
416 jondurbin/airoboros-l2-13b-2.1 llama-2 13B 2023-08-28 2322 -565.56 53.34
417 jondurbin/bagel-8b-v1.0 llama-3 8B 2024-04-24 7960 -512.36 67.84
418 Josephgflowers/TinyLlama-3T-Cinder-v1.2 llama 1B 2023-12-31 1413 -777.50 35.26
419 JosephusCheung/Qwen-LLaMAfied-7B-Chat llama-2 7B 2023-08-04 1221 -646.77 51.99
420 JosephusCheung/Qwen-VL-LLaMAfied-7B-Chat llama-2 7B 2023-08-30 1186 -788.30 45.00
421 jphme/em_german_leo_mistral mistral 7B 2023-10-07 1900 -625.44 51.69
422 jphme/Llama-2-13b-chat-german Touvron et al. (2023b) llama-2 13B 2023-07-21 1265 -572.28 55.07
423 jphme/orca_mini_v2_ger_7b Conover et al. (2023); Mathur (2023a); Su et al. (2023b); Harries (2023); Taori et al. (2023); Touvron et al. (2023a); Xu et al. (2023a) llama-1 7B 2023-07-04 1181 -603.44 47.65
424 junelee/wizard-vicuna-13b llama-1 13B 2023-05-03 2196 -610.98 52.73
425 Kabster/BioMistral-Zephyr-Beta-SLERP mistral 7B 2024-03-09 2758 -597.30 56.35
426 Kabster/Bio-Mistralv2-Squared mistral 7B 2024-03-09 2762 -603.34 57.73
427 kaist-ai/mistral-orpo-capybara-7k Hong et al. (2024) mistral 7B 2024-03-23 4821 -540.67 63.36
428 kekmodel/StopCarbon-10.7B-v5 llama 10B 2023-12-30 13910 -560.12 74.41
429 kevin009/llamaRAGdrama mistral 7B 2024-02-04 4248 -610.94 74.65
430 kfkas/Llama-2-ko-7b-Chat Touvron et al. (2023b) llama-2 7B 2023-07-25 3436 -767.03 40.27
431 kingbri/airolima-chronos-grad-l2-13B llama-2 13B 2023-08-04 1174 -587.96 55.50
432 kingbri/chronolima-airo-grad-l2-13B llama-2 13B 2023-08-04 1182 -582.73 55.50
433 klyang/MentaLLaMA-chat-7B Yang et al. (2024b) llama 7B 2023-09-26 2634 -640.26 51.17
434 KnutJaegersberg/deacon-13b llama 13B 2023-09-20 2018 -533.27 53.63
435 KnutJaegersberg/deacon-3b llama 3B 2023-09-18 2016 -621.85 39.05
436 KnutJaegersberg/falcon-1b-t-sft falcon 1B 2023-12-04 2029 -978.54 35.02
437 KnutJaegersberg/LLongMA-3b-LIMA llama 3B 2023-09-03 2010 -617.15 38.51
438 KnutJaegersberg/MistralInstructLongish mistral 7B 2023-11-15 1970 -543.74 53.62
439 KnutJaegersberg/Qwen-1_8B-Llamafied llama 1B 2024-01-03 2839 -760.12 44.75
440 KnutJaegersberg/Walter-Falcon-1B falcon 1B 2023-12-09 2014 -1010.10 34.07
441 KnutJaegersberg/webMistral-7B mistral 7B 2023-11-17 1972 -554.62 53.97
442 KoboldAI/GPT-J-6B-Adventure gptj 6B 2022-03-02 1297 -661.77 35.95
443 KoboldAI/GPT-J-6B-Janeway Gao et al. (2020); Wang and Komatsuzaki (2021) gptj 6B 2022-03-02 4382 -496.37 39.54
444 KoboldAI/GPT-J-6B-Shinen Gao et al. (2020); Wang and Komatsuzaki (2021) gptj 6B 2022-03-02 1498 -498.02 39.60
445 KoboldAI/GPT-J-6B-Skein Lacoste et al. (2019); Wang (2021) gptj 6B 2022-03-02 1236 -509.27 40.02
446 KoboldAI/LLaMA2-13B-Holomax llama-2 13B 2023-08-14 1255 -563.47 54.52
447 KoboldAI/LLaMA2-13B-Tiefighter llama-2 13B 2023-10-18 2321 -599.88 54.51
448 KoboldAI/OPT-13B-Erebus Zhang et al. (2022b) opt 13B 2022-09-09 6998 -652.81 39.61
449 KoboldAI/OPT-13B-Nerybus-Mix Zhang et al. (2022b) opt 13B 2023-02-13 1657 -646.08 39.61
450 KoboldAI/OPT-13B-Nerys-v2 Zhang et al. (2022b) opt 13B 2022-09-19 4259 -652.01 39.53
451 KoboldAI/OPT-2.7B-Erebus Zhang et al. (2022b) opt 2B 2022-09-19 5559 -692.73 36.96
452 KoboldAI/OPT-2.7B-Nerybus-Mix opt 2B 2023-02-09 1464 -686.59 36.88
453 KoboldAI/OPT-6B-nerys-v2 Zhang et al. (2022b) opt 6B 2022-06-26 4974 -656.62 38.72
454 KoboldAI/OPT-6.7B-Erebus Zhang et al. (2022b) opt 6B 2022-09-15 5587 -641.11 39.09
455 KoboldAI/OPT-6.7B-Nerybus-Mix opt 6B 2023-02-13 1585 -642.68 38.83
456 kodonho/SolarM-SakuraSolar-SLERP llama 10B 2024-01-12 3254 -562.65 74.29
457 kodonho/Solar-OrcaDPO-Solar-Instruct-SLERP llama 10B 2024-01-12 3266 -560.31 74.35
458 Korabbit/Llama-2-7b-chat-hf-afr-100step-flan-v2 llama-2 7B 2023-12-03 1208 -650.96 52.92
459 Korabbit/Llama-2-7b-chat-hf-afr-100step-flan llama-2 7B 2023-11-30 1208 -650.87 52.88
460 Korabbit/Llama-2-7b-chat-hf-afr-100step-v2 llama-2 7B 2023-11-22 1205 -655.71 50.89
461 Korabbit/Llama-2-7b-chat-hf-afr-200step-flan-v2 llama-2 7B 2023-12-03 1211 -648.67 52.75
462 Korabbit/Llama-2-7b-chat-hf-afr-200step-flan llama-2 7B 2023-11-30 1201 -647.72 52.62
463 Korabbit/Llama-2-7b-chat-hf-afr-200step-merged llama-2 7B 2023-11-21 1222 -653.05 52.26
464 Korabbit/Llama-2-7b-chat-hf-afr-200step-v2 llama-2 7B 2023-11-22 1203 -658.69 50.21
465 Korabbit/Llama-2-7b-chat-hf-afr-300step-flan-v2 llama-2 7B 2023-12-03 1208 -647.99 52.41
466 Korabbit/Llama-2-7b-chat-hf-afr-441step-flan-v2 llama-2 7B 2023-12-03 1209 -647.85 52.28
467 Kukedlc/NeuralExperiment-7b-MagicCoder-v7.5 mistral 7B 2024-03-07 4044 -569.72 74.28
468 Kukedlc/NeuralLLaMa-3-8b-DT-v0.1 llama-3 8B 2024-05-11 5707 -571.65 72.52
469 Kukedlc/NeuralLLaMa-3-8b-ORPO-v0.3 llama-3 8B 2024-05-14 7985 -562.41 72.66
470 Kukedlc/NeuralSynthesis-7B-v0.1 mistral 7B 2024-04-06 8910 -596.75 76.80
471 Kukedlc/NeuralSynthesis-7B-v0.3 mistral 7B 2024-04-07 3114 -597.52 76.70
472 Kukedlc/NeuralSynthesis-7b-v0.4-slerp mistral 7B 2024-04-12 3067 -597.78 76.76
473 kyujinpy/Sakura-SOLAR-Instruct-DPO-v2 llama 10B 2023-12-24 3298 -560.08 74.14
474 kyujinpy/Sakura-SOLAR-Instruct llama 10B 2023-12-24 4271 -559.44 74.40
475 kyujinpy/Sakura-SOLRCA-Math-Instruct-DPO-v1 llama 10B 2023-12-25 3284 -562.45 74.13
476 Lazycuber/L2-7b-Base-Guanaco-Uncensored llama 7B 2023-09-19 1172 -555.04 50.45
477 lcw99/llama-3-10b-it-kor-extented-chang-pro8 llama-3 9B 2024-05-21 2234 -561.88 63.76
478 lcw99/llama-3-10b-it-kor-extented-chang llama-3 9B 2024-05-15 2230 -558.58 54.76
479 lcw99/llama-3-10b-it-ko-2024-0527 llama-3 9B 2024-05-27 2236 -564.03 63.70
480 lcw99/llama-3-8b-it-kor-extented-chang llama-3 8B 2024-05-02 2259 -527.86 66.27
481 LDCC/LDCC-SOLAR-10.7B Kim et al. (2024b) llama 10B 2024-01-03 3267 -549.97 71.40
482 lemon-mint/gemma-ko-1.1-2b-it gemma 2B 2024-04-26 2337 -1058.64 30.92
483 lemon-mint/gemma-ko-7b-instruct-v0.62 gemma 8B 2024-04-03 7543 -589.67 69.25
484 lemon-mint/gemma-ko-7b-instruct-v0.71 gemma 8B 2024-04-09 2232 -711.63 59.23
485 lemon-mint/gemma-7b-openhermes-v0.80 gemma 8B 2024-04-09 4867 -691.34 56.91
486 LeoLM/leo-hessianai-7b-chat llama 7B 2023-09-10 3924 -735.02 49.29
487 LeoLM/leo-hessianai-7b llama 7B 2023-08-22 4502 -592.31 47.72
488 leveldevai/TurdusBeagle-7B mistral 7B 2024-01-18 1926 -577.38 75.15
489 lex-hue/Delexa-7b mistral 7B 2024-04-05 11797 -575.31 70.86
490 lgaalves/llama-2-13b-chat-platypus llama-2 13B 2023-09-06 1193 -570.00 53.92
491 lgaalves/llama-2-7b-hf_open-platypus llama-2 6B 2023-08-30 1204 -564.27 49.73
492 lgaalves/mistral-7b-platypus1k mistral 7B 2023-10-10 1183 -539.54 58.19
493 lgaalves/mistral-7b_open_platypus mistral 7B 2023-10-13 1259 -557.93 56.29
494 lgaalves/tinyllama-1.1b-chat-v0.3_platypus llama 1B 2023-10-09 1193 -665.44 34.50
495 lightblue/suzume-llama-3-8B-multilingual Devine (2024) llama-3 8B 2024-04-23 16442 -557.50 65.55
496 lighteternal/Llama3-merge-biomed-8b Yadav et al. (2023a); Yu et al. (2024a) llama-3 8B 2024-05-28 2731 -575.20 66.30
497 liminerity/M7-7b mistral 7B 2024-03-07 4186 -599.03 76.82
498 LinkSoul/Chinese-Llama-2-7b llama-2 7B 2023-07-20 40799 -584.42 52.59
499 llm-agents/tora-code-7b-v1.0 Gou et al. (2024) llama-2 7B 2023-10-08 1242 -682.23 40.21
500 llm-agents/tora-7b-v1.0 Gou et al. (2024) llama-2 7B 2023-10-08 1177 -628.41 48.50
501 lmsys/longchat-13b-16k llama-1 13B 2023-06-28 12228 -621.21 49.64
502 lmsys/longchat-7b-v1.5-32k llama 7B 2023-08-01 5114 -650.92 47.95
503 lmsys/vicuna-13b-delta-v1.1 Zheng et al. (2023); Touvron et al. (2023a) llama-1 13B 2023-04-12 1216 -1396.42 53.28
504 lmsys/vicuna-13b-v1.1 Zheng et al. (2023); Touvron et al. (2023a) llama-1 13B 2023-04-12 2662 -646.61 53.28
505 lmsys/vicuna-13b-v1.3 Zheng et al. (2023); Touvron et al. (2023a) llama-1 13B 2023-06-18 11951 -604.80 54.27
506 lmsys/vicuna-13b-v1.5-16k Touvron et al. (2023b); Zheng et al. (2023) llama-2 13B 2023-08-01 29580 -595.73 54.97
507 lmsys/vicuna-13b-v1.5 Touvron et al. (2023b); Zheng et al. (2023) llama-2 13B 2023-07-29 48653 -585.07 55.41
508 lmsys/vicuna-7b-delta-v1.1 Zheng et al. (2023); Touvron et al. (2023a) llama-1 7B 2023-04-12 1330 -1396.42 50.37
509 lmsys/vicuna-7b-v1.3 Zheng et al. (2023); Touvron et al. (2023a) llama-1 7B 2023-06-18 31436 -621.79 49.78
510 lmsys/vicuna-7b-v1.5-16k Touvron et al. (2023b); Zheng et al. (2023) llama-2 7B 2023-07-31 3362 -622.49 51.42
511 lmsys/vicuna-7b-v1.5 Touvron et al. (2023b); Zheng et al. (2023) llama-2 7B 2023-07-29 368410 -608.63 52.06
512 Locutusque/Hercules-2.5-Mistral-7B mistral 7B 2024-02-10 1953 -533.73 63.59
513 Locutusque/Hercules-3.1-Mistral-7B mistral 7B 2024-02-19 2780 -528.16 62.09
514 Locutusque/Orca-2-13b-SFT-v4 llama 13B 2023-11-25 2345 -617.05 59.75
515 Locutusque/Orca-2-13b-SFT-v6 llama 13B 2023-12-22 2319 -674.35 56.15
516 Locutusque/Orca-2-13b-SFT_v5 llama 13B 2023-12-13 2191 -602.83 56.77
517 LTC-AI-Labs/L2-7b-Base-WVG-Uncensored llama 7B 2023-09-23 1241 -556.11 50.63
518 LTC-AI-Labs/L2-7b-Beluga-WVG-Test llama 7B 2023-10-03 1222 -567.10 52.04
519 LTC-AI-Labs/L2-7b-Hermes-Synthia llama-2 7B 2023-11-23 1232 -562.42 52.21
520 LTC-AI-Labs/L2-7b-Hermes-WVG-Test llama 7B 2023-09-27 1225 -570.23 51.35
521 LTC-AI-Labs/L2-7b-Synthia-WVG-Test llama 7B 2023-09-28 1233 -576.22 51.25
522 luffycodes/nash-vicuna-13b-v1dot5-ep2-w-rag-w-simple Sonkar et al. (2023) llama-2 13B 2023-08-21 1165 -603.30 55.40
523 luffycodes/vicuna-class-shishya-13b-ep3 Sonkar et al. (2023) llama-2 13B 2023-12-21 2092 -633.43 48.52
524 luffycodes/vicuna-class-shishya-7b-ep3 Sonkar et al. (2023) llama-2 7B 2023-12-14 3501 -649.39 46.14
525 luffycodes/vicuna-class-tutor-13b-ep3 Sonkar et al. (2023) llama-2 13B 2023-12-21 1514 -605.75 55.88
526 luffycodes/vicuna-class-tutor-7b-ep3 Sonkar et al. (2023) llama-2 7B 2023-12-15 3735 -643.38 51.45
527 lu-vae/llama2-13B-sharegpt4-orca-openplatypus-8w llama-2 13B 2023-09-14 1190 -544.97 55.75
528 lu-vae/llama2-13b-sharegpt4-test llama-2 13B 2023-09-07 1195 -557.27 55.69
529 lyogavin/Anima-7B-100K llama-2 7B 2023-09-14 1207 -611.20 42.98
530 l3utterfly/llama2-7b-layla llama-2 7B 2023-08-07 1182 -561.12 52.05
531 l3utterfly/minima-3b-layla-v1 llama-2 3B 2023-12-12 1178 -548.62 43.21
532 l3utterfly/minima-3b-layla-v2 llama-2 3B 2023-12-19 1176 -547.03 43.39
533 l3utterfly/mistral-7b-v0.1-layla-v1 mistral 7B 2023-10-31 1206 -602.61 57.56
534 l3utterfly/mistral-7b-v0.1-layla-v2 mistral 7B 2023-12-16 1199 -567.42 57.60
535 l3utterfly/open-llama-3b-v2-layla llama 3B 2023-08-18 1172 -832.18 40.25
536 macadeliccc/laser-dolphin-mixtral-2x7b-dpo Gao et al. (2021a); Sharma et al. (2023) mixtral 12B 2024-01-08 1275 -594.20 67.16
537 macadeliccc/WestLake-7B-v2-laser-truthy-dpo mistral 7B 2024-01-27 5064 -593.71 75.37
538 malhajar/Mistral-7B-v0.2-meditron-turkish mistral 7B 2024-01-05 3865 -641.84 63.34
539 martyn/llama2-megamerge-dare-13b-v2 llama-2 13B 2023-12-17 1213 -594.59 57.94
540 martyn/mistral-megamerge-dare-7b mistral 7B 2023-12-14 1203 -638.88 48.93
541 martyn/solar-megamerge-dare-10.7b-v1 llama 10B 2023-12-31 1190 -533.45 68.79
542 matsuo-lab/weblab-10b-instruction-sft gpt_neox 10B 2023-08-04 1235 -515.77 39.13
543 matsuo-lab/weblab-10b gpt_neox 10B 2023-08-04 1596 -485.56 38.59
544 MayaPH/FinOPT-Franklin opt 1B 2023-05-26 1207 -1380.66 29.78
545 MayaPH/opt-flan-iml-6.7b Iyer et al. (2023) opt 6B 2023-08-15 1193 -729.78 35.84
546 maywell/PiVoT-0.1-early mistral 7B 2023-11-24 3267 -681.01 64.58
547 maywell/PiVoT-10.7B-Mistral-v0.2 mistral 10B 2023-12-13 3228 -578.51 64.25
548 maywell/Synatra-RP-Orca-2-7b-v0.1 llama 6B 2023-11-21 3278 -632.02 59.65
549 maywell/Synatra-V0.1-7B-Instruct mistral 7B 2023-10-09 3381 -623.00 55.86
550 maywell/Synatra-10.7B-v0.4 llama 10B 2023-12-27 3282 -532.85 65.48
551 maywell/Synatra-7B-v0.3-dpo mistral 7B 2023-11-08 4302 -575.15 60.55
552 maywell/Synatra-7B-v0.3-RP mistral 7B 2023-10-29 8360 -582.37 59.26
553 MaziyarPanahi/Llama-3-8B-Instruct-v0.4 llama-3 8B 2024-05-01 1407 -569.34 70.30
554 MaziyarPanahi/Llama-3-8B-Instruct-v0.8 llama-3 8B 2024-05-01 7281 -576.95 73.17
555 MaziyarPanahi/Llama-3-8B-Instruct-v0.9 llama-3 8B 2024-05-30 6241 -575.60 73.29
556 MaziyarPanahi/Mistral-7B-Instruct-v0.3 mistral 7B 2024-05-22 5230 -549.67 65.21
557 MaziyarPanahi/Mistral-7B-v0.3 mistral 7B 2024-05-22 5758 -531.44 60.40
558 medalpaca/medalpaca-7b Li et al. (2023d) llama-1 7B 2023-03-29 8136 -628.90 48.45
559 MediaTek-Research/Breeze-7B-Instruct-v1_0 Hsu et al. (2024) mistral 7B 2024-03-05 3065 -559.49 63.40
560 meta-llama/Llama-2-13b-chat-hf Touvron et al. (2023b) llama-2 13B 2023-07-13 266090 -616.08 54.91
561 meta-llama/Llama-2-13b-hf Touvron et al. (2023b) llama-2 13B 2023-07-13 120871 -530.82 55.69
562 meta-llama/Llama-2-7b-chat-hf Touvron et al. (2023b) llama-2 6B 2023-07-13 1402244 -656.64 50.74
563 meta-llama/Llama-2-7b-hf Touvron et al. (2023b) llama-2 6B 2023-07-13 1294737 -549.87 50.97
564 meta-llama/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 8B 2024-04-17 2058011 -573.95 66.87
565 meta-llama/Meta-Llama-3-8B AI@Meta (2024) llama-3 8B 2024-04-17 675850 -514.33 62.62
566 meta-math/MetaMath-Llemma-7B Azerbayev et al. (2024); Yu et al. (2024b) llama 7B 2023-11-19 1322 -612.21 53.19
567 meta-math/MetaMath-Mistral-7B Jiang et al. (2023); Yu et al. (2024b) mistral 7B 2023-10-22 3022 -571.33 65.78
568 microsoft/Orca-2-13b Mitra et al. (2023) llama 13B 2023-11-14 13616 -651.29 58.64
569 microsoft/Orca-2-7b Mitra et al. (2023) llama 7B 2023-11-14 11084 -684.33 54.55
570 microsoft/phi-1_5 Li et al. (2023c) phi 1B 2023-09-10 112476 -724.13 47.69
571 microsoft/phi-2 phi 2B 2023-12-13 237268 -621.44 61.33
572 microsoft/Phi-3-medium-128k-instruct phi3 13B 2024-05-07 34054 -544.42 73.00
573 microsoft/Phi-3-medium-4k-instruct phi3 13B 2024-05-07 35987 -552.43 73.45
574 migtissera/Llama-3-8B-Synthia-v3.5 llama-3 8B 2024-05-17 3310 -533.92 67.15
575 migtissera/SynthIA-7B-v1.3 Mukherjee et al. (2023); Tissera (2023) mistral 7B 2023-09-28 3336 -541.46 59.34
576 migtissera/SynthIA-7B-v1.5 mistral 7B 2023-10-07 1228 -537.68 59.59
577 migtissera/Synthia-7B-v3.0 mistral 7B 2023-12-08 1204 -531.17 61.99
578 migtissera/Tess-XS-v1.1 mistral 7B 2023-11-22 1213 -546.97 59.39
579 migtissera/Tess-2.0-Llama-3-8B llama-3 8B 2024-05-05 3314 -523.68 64.81
580 migtissera/Tess-7B-v1.4 mistral 7B 2023-12-04 1234 -599.60 62.19
581 Mihaiii/Metis-0.3 mistral 7B 2023-12-16 1279 -567.64 65.44
582 mindy-labs/mindy-7b-v2 mistral 7B 2023-12-14 1238 -550.38 72.11
583 Minirecord/Mini_DPO_test02 mistral 7B 2023-11-30 3470 -547.90 61.23
584 mistralai/Mistral-7B-Instruct-v0.1 Jiang et al. (2023) mistral 7B 2023-09-27 1370245 -593.06 54.96
585 mistralai/Mistral-7B-Instruct-v0.2 Jiang et al. (2023) mistral 7B 2023-12-11 3565248 -574.91 65.71
586 mistralai/Mistral-7B-v0.1 Jiang et al. (2023) mistral 7B 2023-09-20 1750089 -527.60 60.97
587 mistralai/Mistral-7B-v0.3 Jiang et al. (2023) mistral 7B 2024-05-22 1443065 -531.44 60.28
588 mistral-community/Mistral-7B-v0.2 mistral 7B 2024-03-23 30074 -532.63 60.41
589 mlabonne/AlphaMonarch-7B mistral 7B 2024-02-14 12804 -593.46 75.99
590 mlabonne/Beagle14-7B mistral 7B 2024-01-15 1986 -563.97 74.76
591 mlabonne/ChimeraLlama-3-8B-v2 llama-3 8B 2024-04-22 2806 -565.81 69.69
592 mlabonne/ChimeraLlama-3-8B-v3 llama-3 8B 2024-05-01 5718 -568.78 70.06
593 mlabonne/Daredevil-8B-abliterated llama 8B 2024-05-26 9010 -565.64 71.82
594 mlabonne/GML-Mistral-merged-v1 mistral 8B 2023-12-27 1166 -1384.76 48.54
595 mlabonne/Marcoro14-7B-slerp mistral 7B 2023-12-29 3792 -554.34 73.01
596 mlabonne/NeuralMarcoro14-7B mistral 7B 2024-01-06 2015 -566.69 73.57
597 mlabonne/NeuralMonarch-7B mistral 7B 2024-02-14 13486 -591.01 76.15
598 mncai/agiin-13.6B-v0.1 mistral 13B 2023-12-15 3317 -595.75 68.40
599 MoaData/Myrrh_solar_10.7b_3.0 llama 10B 2024-04-26 9855 -673.46 67.61
600 monology/openinstruct-mistral-7b mistral 7B 2023-11-20 1191 -537.83 63.64
601 mosaicml/mpt-7b-chat Henry et al. (2020); Dao et al. (2022); Press et al. (2022); MosaicML NLP Team (2023b) mpt 7B 2023-05-04 88069 -647.13 45.39
602 mosaicml/mpt-7b-instruct Henry et al. (2020); Dao et al. (2022); Press et al. (2022); MosaicML NLP Team (2023b) mpt 7B 2023-05-05 8599 -569.62 44.83
603 mosaicml/mpt-7b-storywriter Dao et al. (2022); Press et al. (2022); Chen et al. (2023a); MosaicML NLP Team (2023b) mpt 7B 2023-05-04 1920 -650.05 39.31
604 mosaicml/mpt-7b-8k-chat Henry et al. (2020); Dao et al. (2022); Press et al. (2022); MosaicML NLP Team (2023a) mpt 7B 2023-06-22 1304 -586.09 47.78
605 mosaicml/mpt-7b-8k Henry et al. (2020); Shoeybi et al. (2020); Press et al. (2022); Dao et al. (2022); Chen et al. (2023a); MosaicML NLP Team (2023b); Touvron et al. (2023a) mpt 7B 2023-06-30 2106 -546.57 47.24
606 mosaicml/mpt-7b Henry et al. (2020); Shoeybi et al. (2020); Press et al. (2022); Dao et al. (2022); Chen et al. (2023a); MosaicML NLP Team (2023b); Touvron et al. (2023a) mpt 7B 2023-05-05 31480 -548.53 44.28
607 mrm8488/llama-2-coder-7b Manuel Romero (2023) llama-2 7B 2023-07-26 1313 -572.92 49.95
608 MTSAIR/multi_verse_model mistral 7B 2024-03-07 6357 -598.96 76.74
609 mwitiderrick/open_llama_3b_code_instruct_0.1 llama 3B 2023-12-11 1252 -598.99 39.72
610 NekoPunchBBB/Llama-2-13b-hf_Open-Platypus-QLoRA-multigpu llama-2 13B 2023-09-15 1208 -536.60 54.40
611 Neko-Institute-of-Science/metharme-7b llama-1 6B 2023-04-30 1161 -562.36 47.48
612 Neko-Institute-of-Science/pygmalion-7b llama-1 6B 2023-04-30 1188 -562.13 46.04
613 NeverSleep/Llama-3-Lumimaid-8B-v0.1 llama-3 8B 2024-04-30 1499 -545.48 66.55
614 NeverSleep/Noromaid-7b-v0.2 mistral 7B 2023-12-21 1179 -552.34 61.78
615 NewstaR/Koss-7B-chat llama 7B 2023-10-01 1180 -660.75 50.37
616 NewstaR/Starlight-7B Clark et al. (2018); Zellers et al. (2019); Hendrycks et al. (2021a); Gao et al. (2021a); Lin et al. (2022a); Beeching et al. (2023) llama-2 7B 2023-09-11 1180 -549.87 49.73
617 NExtNewChattingAI/shark_tank_ai_7b_v2 mistral 7B 2023-12-23 1208 -618.14 66.54
618 NExtNewChattingAI/shark_tank_ai_7_b mistral 7B 2023-12-17 1224 -546.93 71.10
619 Nexusflow/NexusRaven-V2-13B Nexusflow.ai team (2023); Rozière et al. (2024) llama 13B 2023-12-04 3966 -593.55 48.21
620 Nexusflow/Starling-LM-7B-beta Ziegler et al. (2020); Zhu et al. (2023a) mistral 7B 2024-03-19 4847 -559.22 69.88
621 NickyNicky/Mistral-7B-OpenOrca-oasst_top1_2023-08-25-v2 Xiao et al. (2024) mistral 7B 2023-10-11 1236 -560.11 61.65
622 NickyNicky/Mistral-7B-OpenOrca-oasst_top1_2023-08-25-v3 Dao et al. (2022); Xiao et al. (2024) mistral 7B 2023-10-13 1227 -568.73 61.26
623 nlpguy/ColorShadow-7B-v2 mistral 7B 2023-12-30 1214 -569.63 66.88
624 nlpguy/ColorShadow-7B-v3 mistral 7B 2023-12-30 1222 -561.80 67.29
625 nlpguy/ColorShadow-7B mistral 7B 2023-12-30 1214 -552.87 68.34
626 Norquinal/llama-2-7b-claude-chat-rp llama-2 7B 2023-08-14 1218 -557.09 51.25
627 Norquinal/llama-2-7b-claude-chat llama-2 7B 2023-08-11 1221 -557.34 50.98
628 Norquinal/Mistral-7B-claude-instruct mistral 7B 2023-09-28 1248 -534.14 59.27
629 NousResearch/CodeLlama-13b-hf llama 13B 2023-08-24 6299 -582.02 43.35
630 NousResearch/CodeLlama-7b-hf llama 7B 2023-08-24 9428 -597.05 39.81
631 NousResearch/Hermes-2-Pro-Llama-3-8B "Teknium" et al. (2024b) llama-3 8B 2024-04-30 26797 -576.90 68.73
632 NousResearch/Hermes-2-Pro-Mistral-7B "interstellarninja" et al. (2024) mistral 7B 2024-03-11 14586 -620.93 67.35
633 NousResearch/Hermes-2-Theta-Llama-3-8B "Teknium" et al. (2024a) llama-3 8B 2024-05-05 12392 -565.47 69.21
634 NousResearch/Meta-Llama-3-8B-Instruct AI@Meta (2024) llama-3 8B 2024-04-18 104388 -573.95 67.10
635 NousResearch/Nous-Hermes-Llama2-13b llama-2 13B 2023-07-20 39412 -586.34 55.75
636 NousResearch/Nous-Hermes-llama-2-7b llama-2 6B 2023-07-25 12019 -569.93 51.87
637 NousResearch/Nous-Hermes-13b llama-1 13B 2023-06-03 1536 -605.76 54.04
638 NousResearch/Nous-Hermes-2-Mistral-7B-DPO "Teknium" et al. (2024c) mistral 7B 2024-02-18 8725 -565.26 68.10
639 NousResearch/Nous-Hermes-2-SOLAR-10.7B llama 10B 2024-01-01 9918 -542.37 71.00
640 NousResearch/Yarn-Mistral-7b-128k Peng et al. (2023b) mistral 7B 2023-10-31 20727 -532.22 59.42
641 NousResearch/Yarn-Mistral-7b-64k Peng et al. (2023b) mistral 7B 2023-10-31 10368 -532.16 59.63
642 NTQAI/Nxcode-CQ-7B-orpo Hong et al. (2024) qwen2 7B 2024-04-24 12700 -685.73 42.98
643 nvidia/Llama3-ChatQA-1.5-8B Liu et al. (2024) llama-3 8B 2024-04-28 10608 -515.82 56.71
644 occultml/CatMarcoro14-7B-slerp mistral 7B 2024-01-06 1234 -551.33 73.25
645 occultml/Helios-10.7B-v2 llama 7B 2023-12-31 1218 -1381.35 42.25
646 occultml/Helios-10.7B llama 7B 2023-12-31 1215 -1381.38 42.19
647 OEvortex/EMO-2B gemma 2B 2024-04-28 4137 -976.17 44.26
648 oh-yeontaek/llama-2-13B-LoRA-assemble llama-2 13B 2023-09-13 3303 -568.38 57.91
649 oh-yeontaek/llama-2-7B-LoRA-assemble llama-2 7B 2023-09-13 1337 -614.33 52.26
650 openaccess-ai-collective/DPOpenHermes-11B mistral 10B 2023-12-03 1185 -591.33 66.83
651 openaccess-ai-collective/jackalope-7b Mukherjee et al. (2023); Lian et al. (2023b); Longpre et al. (2023) mistral 7B 2023-10-07 1198 -550.33 61.16
652 openaccess-ai-collective/manticore-13b-chat-pyg llama-1 13B 2023-05-22 2148 -553.85 54.13
653 openaccess-ai-collective/manticore-13b llama-1 13B 2023-05-17 1224 -553.22 54.86
654 openaccess-ai-collective/minotaur-13b-fixed llama-1 13B 2023-06-12 1194 -568.52 55.19
655 openaccess-ai-collective/minotaur-13b llama-1 13B 2023-06-06 1202 -577.38 53.97
656 openaccess-ai-collective/mistral-7b-slimorcaboros mistral 7B 2023-10-13 1186 -569.20 61.18
657 openaccess-ai-collective/wizard-mega-13b llama-1 13B 2023-05-14 2168 -565.46 54.27
658 OpenAssistant/codellama-13b-oasst-sft-v10 llama-2 13B 2023-08-26 2122 -591.56 44.85
659 OpenAssistant/llama2-13b-orca-8k-3319 Mukherjee et al. (2023) llama-2 13B 2023-07-24 1270 -531.87 55.09
660 OpenAssistant/oasst-sft-1-pythia-12b gpt_neox 12B 2023-03-09 22359 -615.25 40.77
661 OpenAssistant/oasst-sft-4-pythia-12b-epoch-3.5 gpt_neox 12B 2023-04-03 458718 -572.27 41.31
662 OpenAssistant/pythia-12b-sft-v8-7k-steps gpt_neox 12B 2023-05-07 1323 -524.99 42.21
663 OpenAssistant/stablelm-7b-sft-v7-epoch-3 gpt_neox 7B 2023-04-20 1215 -764.93 34.85
664 openbmb/UltraLM-13b-v2.0 llama 13B 2023-09-22 1174 -554.24 58.72
665 OpenBuddy/openbuddy-atom-13b-v9-bf16 llama 13B 2023-08-05 1206 -625.44 52.31
666 OpenBuddy/openbuddy-llama2-13b-v11-bf16 llama-2 13B 2023-08-23 1199 -606.17 52.93
667 OpenBuddy/openbuddy-llama2-13b-v11.1-bf16 llama-2 13B 2023-08-24 1210 -595.93 55.28
668 OpenBuddy/openbuddy-llama2-13b-v8.1-fp16 llama-2 13B 2023-07-25 7339 -587.40 57.76
669 OpenBuddy/openbuddy-llama3-8b-v21.1-8k llama-3 8B 2024-04-20 2293 -578.14 65.31
670 OpenBuddy/openbuddy-mistral2-7b-v20.3-32k mistral 7B 2024-03-27 2302 -642.78 62.73
671 OpenBuddy/openbuddy-mistral-7b-v13-base mistral 7B 2023-10-11 1197 -677.88 51.99
672 OpenBuddy/openbuddy-mistral-7b-v13.1 mistral 7B 2023-10-11 1217 -638.68 52.62
673 OpenBuddy/openbuddy-mistral-7b-v13 mistral 7B 2023-10-10 1335 -643.47 53.50
674 OpenBuddy/openbuddy-openllama-13b-v7-fp16 llama-1 13B 2023-07-03 1209 -654.65 49.31
675 OpenBuddy/openbuddy-openllama-3b-v10-bf16 llama 3B 2023-08-10 1206 -794.41 36.87
676 OpenBuddy/openbuddy-openllama-7b-v12-bf16 llama 7B 2023-09-19 3153 -854.52 45.28
677 OpenBuddy/openbuddy-zephyr-7b-v14.1 mistral 7B 2023-11-06 3392 -657.03 51.86
678 openchat/openchat-3.5-0106-gemma Wang et al. (2024a) gemma 8B 2024-03-09 7817 -937.52 69.42
679 openchat/openchat-3.5-0106 Wang et al. (2024a); OpenAI et al. (2024) mistral 7B 2024-01-07 26858 -554.42 69.30
680 openchat/openchat-3.5-1210 Wang et al. (2024a); OpenAI et al. (2024) mistral 7B 2023-12-12 2123 -583.16 68.89
681 openchat/openchat-3.6-8b-20240522 Wang et al. (2024a) llama-3 8B 2024-05-07 10464 -561.80 68.14
682 openlm-research/open_llama_13b Geng and Liu (2023); Together Computer (2023); Touvron et al. (2023a) llama-1 13B 2023-06-15 2373 -582.20 47.26
683 openlm-research/open_llama_3b_v2 Geng and Liu (2023); Together Computer (2023); Touvron et al. (2023a) llama-1 3B 2023-07-16 21275 -578.08 40.28
684 openlm-research/open_llama_3b Geng and Liu (2023); Together Computer (2023); Touvron et al. (2023a) llama-1 3B 2023-06-07 157405 -612.46 38.26
685 openlm-research/open_llama_7b_v2 Geng and Liu (2023); Together Computer (2023); Touvron et al. (2023a) llama-1 7B 2023-07-06 2792 -556.22 44.26
686 openlm-research/open_llama_7b Geng and Liu (2023); Together Computer (2023); Touvron et al. (2023a) llama-1 7B 2023-06-07 44968 -610.20 42.31
687 OpenModels4all/gemma-1.1-7b-it Joshi et al. (2017); Zhao et al. (2018); Mihaylov et al. (2018); Clark et al. (2019); Talmor et al. (2019); Sap et al. (2019); Bisk et al. (2019); Chollet (2019); Sakaguchi et al. (2019); Zellers et al. (2019); Austin et al. (2021); Cobbe et al. (2021); Hendrycks et al. (2021a); Chen et al. (2021); Parrish et al. (2022); Srivastava et al. (2023); Zhong et al. (2023); Gemini Team et al. (2024) gemma 8B 2024-04-06 5128 -1355.60 59.78
688 OpenPipe/mistral-ft-optimized-1218 mistral 7B 2023-12-17 1405 -546.73 71.94
689 OpenPipe/mistral-ft-optimized-1227 mistral 7B 2023-12-27 6273 -564.28 70.54
690 openthaigpt/openthaigpt-1.0.0-beta-13b-chat-hf llama 13B 2023-12-18 1225 -712.61 50.45
691 openthaigpt/openthaigpt-1.0.0-beta-7b-chat-ckpt-hf llama 7B 2023-08-14 1291 -772.50 45.35
692 Open-Orca/LlongOrca-13B-16k Mukherjee et al. (2023); Wang et al. (2023a); Dale et al. (2023); Longpre et al. (2023); Touvron et al. (2023b) llama-2 13B 2023-08-16 1217 -551.07 56.59
693 Open-Orca/LlongOrca-7B-16k Mukherjee et al. (2023); Wang et al. (2023a); Longpre et al. (2023); Touvron et al. (2023b); Lian et al. (2023c) llama-2 7B 2023-08-05 1223 -583.06 53.02
694 Open-Orca/Mistral-7B-OpenOrca Mukherjee et al. (2023); Longpre et al. (2023); Lian et al. (2023d) mistral 7B 2023-09-29 18070 -575.75 60.17
695 Open-Orca/Mistral-7B-SlimOrca Mukherjee et al. (2023); Lian et al. (2023f, e); Longpre et al. (2023) mistral 7B 2023-10-08 3696 -569.42 60.37
696 Open-Orca/OpenOrcaxOpenChat-Preview2-13B Mukherjee et al. (2023); Wang et al. (2023a, b); Longpre et al. (2023); Touvron et al. (2023b) llama-2 13B 2023-07-31 2172 -547.85 56.70
697 Open-Orca/OpenOrca-Platypus2-13B Hu et al. (2022); Mukherjee et al. (2023); Wang et al. (2023a, b); Lee et al. (2023); Longpre et al. (2023); Touvron et al. (2023b); Lee et al. (2024) llama 13B 2023-08-11 6853 -557.72 57.28
698 Open-Orca/OpenOrca-Preview1-13B Mukherjee et al. (2023); Lian et al. (2023a); Longpre et al. (2023); Touvron et al. (2023a) llama-1 13B 2023-07-12 1240 -679.87 51.38
699 Orenguteng/Llama-3-8B-Lexi-Uncensored llama-3 8B 2024-04-23 9653 -564.93 66.18
700 pankajmathur/orca_mini_v3_13b Mukherjee et al. (2023); Touvron et al. (2023b); Mathur (2023b) llama 13B 2023-08-09 4717 -546.48 57.24
701 pankajmathur/orca_mini_v3_7b Mukherjee et al. (2023); Touvron et al. (2023b); Mathur (2023c); Touvron et al. (2023a) llama 7B 2023-08-07 2338 -572.18 53.47
702 PathFinderKR/Waktaverse-Llama-3-KO-8B-Instruct AI@Meta (2024); AI@Waktaverse (2024) llama-3 8B 2024-04-19 2305 -552.92 66.77
703 pe-nlp/llama-2-13b-vicuna-wizard llama-2 13B 2023-08-11 1176 -549.39 51.94
704 pillowtalks-ai/delta13b llama-1 13B 2023-04-14 1168 -646.61 53.29
705 Pirr/pythia-13b-deduped-green_devil gpt_neox 13B 2023-02-09 1351 -505.96 40.31
706 PistachioAlt/Synatra-MCS-7B-v0.3-RP-Slerp mistral 7B 2023-12-11 1162 -549.18 69.18
707 PracticeLLM/SOLAR-tail-10.7B-Merge-v1.0 llama 10B 2023-12-26 2264 -530.70 71.68
708 princeton-nlp/Sheared-LLaMA-1.3B-ShareGPT Xia et al. (2024) llama 1B 2023-11-22 1737 -721.70 37.14
709 princeton-nlp/Sheared-LLaMA-1.3B Xia et al. (2024) llama 1B 2023-10-10 28905 -646.81 35.95
710 princeton-nlp/Sheared-LLaMA-2.7B-ShareGPT Xia et al. (2024) llama-2 2B 2023-11-22 1866 -681.47 42.11
711 princeton-nlp/Sheared-LLaMA-2.7B Xia et al. (2024) llama-2 2B 2023-10-10 2590 -610.74 40.84
712 project-baize/baize-v2-13b Xu et al. (2023b) llama-1 13B 2023-05-23 2149 -578.69 52.94
713 project-baize/baize-v2-7b Xu et al. (2023b) llama-1 7B 2023-05-23 1212 -638.05 46.72
714 PygmalionAI/metharme-1.3b gpt_neox 1B 2023-06-02 1235 -541.84 35.04
715 PygmalionAI/mythalion-13b llama-2 13B 2023-09-05 2383 -552.46 56.48
716 PygmalionAI/pygmalion-1.3b gpt_neox 1B 2022-12-25 1519 -963.99 31.14
717 PygmalionAI/pygmalion-2-13b llama-2 13B 2023-09-04 2083 -540.27 55.12
718 PygmalionAI/pygmalion-2-7b llama-2 6B 2023-09-04 2063 -559.12 51.11
719 PygmalionAI/pygmalion-2.7b gpt_neo 2B 2023-01-05 1986 -659.52 33.98
720 PygmalionAI/pygmalion-6b gptj 6B 2023-01-07 4312 -555.36 38.47
721 qnguyen3/Master-Yi-9B llama 8B 2024-05-18 8810 -522.22 67.44
722 quantumaikr/llama-2-7b-hf-guanaco-1k llama-2 7B 2023-08-06 1186 -605.93 50.13
723 quantumaikr/QuantumLM-7B llama 7B 2023-07-22 1192 -625.18 49.51
724 quantumaikr/quantum-v0.01 mistral 7B 2023-12-17 1192 -559.84 74.68
725 Qwen/CodeQwen1.5-7B-Chat Bai et al. (2023) qwen2 7B 2024-04-15 77169 -691.05 43.26
726 Qwen/Qwen1.5-1.8B-Chat Bai et al. (2023) qwen2 1B 2024-01-30 10856 -673.92 43.99
727 Qwen/Qwen1.5-1.8B Bai et al. (2023) qwen2 1B 2024-01-22 137256 -610.58 46.55
728 Qwen/Qwen1.5-4B-Chat Bai et al. (2023) qwen2 3B 2024-01-30 5525 -601.85 46.79
729 Qwen/Qwen1.5-4B Bai et al. (2023) qwen2 3B 2024-01-22 6431 -553.86 57.05
730 Qwen/Qwen1.5-7B-Chat Bai et al. (2023) qwen2 7B 2024-01-30 26143 -608.59 55.15
731 Qwen/Qwen1.5-7B Bai et al. (2023) qwen2 7B 2024-01-22 122768 -533.48 61.76
732 Qwen/Qwen2-1.5B Yang et al. (2024a) qwen2 1B 2024-05-31 46510 -581.07 55.80
733 Qwen/Qwen2-7B Yang et al. (2024a) qwen2 7B 2024-06-04 37173 -507.30 68.40
734 Q-bert/Bumblebee-7B mistral 7B 2023-12-03 1226 -549.82 67.73
735 Q-bert/Optimus-7B mistral 7B 2023-12-03 1237 -550.18 69.09
736 Q-bert/Terminis-7B mistral 7B 2023-12-12 1233 -564.26 70.73
737 refuelai/Llama-3-Refueled llama-3 8B 2024-05-03 1246 -582.52 63.62
738 revolutionarybukhari/Llama-2-7b-chat-finetune-AUTOMATE llama-2 7B 2023-10-14 1166 -613.72 50.68
739 rinna/bilingual-gpt-neox-4b Zhao et al. (2023b); Sawada et al. (2024) gpt_neox 3B 2023-07-31 3267 -594.93 32.14
740 rinna/japanese-gpt-neox-3.6b Zhao and Sawada (2023); Sawada et al. (2024) gpt_neox 3B 2023-05-17 6042 -1117.39 29.28
741 rinna/llama-3-youko-8b Andonian et al. (2021); AI@Meta (2024); Sawada et al. (2024); Mitsuda et al. (2024) llama-3 8B 2024-05-01 1468 -502.57 57.55
742 rinna/youri-7b Andonian et al. (2021); Touvron et al. (2023b); Zhao et al. (2023a); Sawada et al. (2024) llama-2 7B 2023-10-30 2512 -545.03 47.11
743 rishiraj/CatPPT-base Acharya (2023) mistral 7B 2023-12-17 4392 -558.36 72.25
744 RoversX/llama-2-7b-hf-small-shards-Samantha-V1-SFT llama-2 7B 2023-08-11 1164 -558.46 49.96
745 RubielLabarta/LogoS-7Bx2-MoE-13B-v0.2 mixtral 12B 2024-01-21 3055 -589.76 77.15
746 ruslanmv/ai-medical-model-32bit llama 8B 2024-05-13 2810 -557.43 67.67
747 ruslanmv/Medical-Llama3-8B llama-3 8B 2024-04-21 5457 -514.00 60.61
748 rwitz2/go-bruins-v2.1.1 Murias (2023) mistral 7B 2023-12-14 1205 -562.22 74.95
749 rwitz2/go-bruins-v2.1 Murias (2023) mistral 7B 2023-12-14 1193 -561.68 74.50
750 RWKV/rwkv-raven-1b5 rwkv 1B 2023-05-04 1570 -526.30 33.56
751 saberai/Zro1.5_3B gpt_neox 2B 2023-12-25 1200 -695.87 38.02
752 Salesforce/codegen-6B-multi Nijkamp et al. (2023) codegen 6B 2022-04-13 1651 -733.14 32.43
753 Salesforce/codegen-6B-nl Nijkamp et al. (2023) codegen 6B 2022-04-13 1176 -478.78 40.00
754 samir-fama/FernandoGPT-v1 mistral 7B 2023-12-30 1209 -551.17 72.87
755 samir-fama/SamirGPT-v1 mistral 7B 2023-12-28 1215 -552.06 73.11
756 SanjiWatsuki/Kunoichi-DPO-v2-7B mistral 7B 2024-01-13 1216 -586.90 72.40
757 SanjiWatsuki/Kunoichi-7B mistral 7B 2024-01-04 1230 -579.35 72.13
758 SanjiWatsuki/Loyal-Macaroni-Maid-7B mistral 7B 2023-12-24 1297 -575.31 71.68
759 SanjiWatsuki/Silicon-Maid-7B mistral 7B 2023-12-27 1578 -580.90 70.31
760 SanjiWatsuki/Sonya-7B mistral 7B 2023-12-31 5399 -594.55 68.48
761 sarvamai/OpenHathi-7B-Hi-v0.1-Base llama-2 6B 2023-12-13 1766 -673.13 46.64
762 scaledown/ScaleDown-7B-slerp-v0.1 mistral 7B 2024-01-01 1206 -538.58 71.57
763 scb10x/llama-3-typhoon-v1.5-8b-instruct Pipatanakul et al. (2023) llama-3 8B 2024-05-06 6088 -590.46 65.62
764 scb10x/typhoon-7b Pipatanakul et al. (2023) mistral 7B 2023-12-20 1908 -617.60 58.05
765 SciPhi/SciPhi-Self-RAG-Mistral-7B-32k Mukherjee et al. (2023); Longpre et al. (2023); Asai et al. (2023) mistral 7B 2023-10-27 1214 -664.82 56.46
766 SeaLLMs/SeaLLM-7B-v2.5 Zhang et al. (2023); Nguyen et al. (2024) gemma 8B 2024-04-03 13928 -565.35 69.07
767 SeaLLMs/SeaLLM-7B-v2 Kojima et al. (2023); Zheng et al. (2023); Zhang et al. (2023); Nguyen et al. (2024) mistral 7B 2024-01-29 6294 -548.56 67.57
768 selfrag/selfrag_llama2_7b Asai et al. (2023) llama-2 7B 2023-10-18 4388 -605.61 51.30
769 senseable/WestLake-7B-v2 mistral 7B 2024-01-22 1189 -594.85 74.68
770 sethuiyer/Medichat-Llama3-8B llama-3 8B 2024-04-22 4542 -535.77 66.03
771 Severian/ANIMA-Phi-Neptune-Mistral-7B mistral 7B 2023-10-11 1191 -631.73 55.54
772 shadowml/BeagSake-7B mistral 7B 2024-01-31 11842 -562.99 75.38
773 shanchen/llama3-8B-slerp-biomed-chat-chinese llama-3 8B 2024-04-30 2708 -591.11 63.00
774 shanchen/llama3-8B-slerp-med-chinese llama-3 8B 2024-04-30 8028 -593.26 58.99
775 shanchen/llama3-8B-slerp-med-262k llama-3 8B 2024-04-30 2697 -599.35 53.65
776 shenzhi-wang/Llama3-8B-Chinese-Chat Wang et al. (2024b) llama-3 8B 2024-04-21 55718 -550.92 67.10
777 shibing624/chinese-alpaca-plus-7b-hf Ming (2023) llama-1 7B 2023-05-01 1527 -741.70 44.77
778 shitshow123/tinylamma-20000 llama 1B 2024-01-09 1195 -1213.36 27.95
779 SJ-Donald/llama3-passthrough-chat llama-3 11B 2024-05-17 2241 -607.16 60.15
780 SJ-Donald/SJ-SOLAR-10.7b-DPO llama 10B 2024-01-25 2306 -533.83 72.67
781 SJ-Donald/SOLAR-10.7B-slerp llama 10B 2024-01-12 2296 -533.67 72.58
782 skfrost19/BioMistralMerged mistral 7B 2024-04-21 4013 -600.14 57.44
783 speakleash/Bielik-7B-Instruct-v0.1 Levine et al. (2020); Granziol et al. (2021); Ociepa et al. (2024b); Wang et al. (2024a); Ociepa et al. (2024c) mistral 7B 2024-03-30 3573 -871.06 51.24
784 speakleash/Bielik-7B-v0.1 Ociepa et al. (2024a, c) mistral 7B 2024-03-30 2822 -721.18 50.01
785 stabilityai/japanese-stablelm-base-gamma-7b Jiang et al. (2023) mistral 7B 2023-10-16 2072 -581.58 52.59
786 stabilityai/japanese-stablelm-instruct-gamma-7b Jiang et al. (2023) mistral 7B 2023-10-16 1412 -584.50 52.82
787 stabilityai/StableBeluga-13B Touvron et al. (2023b); Mukherjee et al. (2023) llama 13B 2023-07-27 6154 -540.85 57.05
788 stabilityai/StableBeluga-7B Touvron et al. (2023b); Mukherjee et al. (2023) llama 6B 2023-07-27 6691 -565.19 53.56
789 stabilityai/stablelm-base-alpha-3b Andonian et al. (2021) gpt_neox 3B 2023-04-17 1728 -677.47 31.50
790 stabilityai/stablelm-base-alpha-7b-v2 Shazeer (2020); Rajbhandari et al. (2020); Gao et al. (2020); Li et al. (2023a); Su et al. (2023a); Tow (2023) stablelm_alpha 6B 2023-08-04 2209 -505.45 46.18
791 stabilityai/stablelm-base-alpha-7b Andonian et al. (2021) gpt_neox 7B 2023-04-11 1750 -623.09 34.37
792 stabilityai/stablelm-tuned-alpha-3b Taori et al. (2023); Anand et al. (2023); Chiang et al. (2023) gpt_neox 3B 2023-04-19 2145 -736.44 32.14
793 stabilityai/stablelm-tuned-alpha-7b Taori et al. (2023); Anand et al. (2023); Chiang et al. (2023) gpt_neox 7B 2023-04-19 3765 -694.37 34.04
794 stabilityai/stablelm-zephyr-3b Zheng et al. (2023); Rafailov et al. (2024) stablelm 2B 2023-11-21 8261 -777.98 53.43
795 stabilityai/stablelm-2-zephyr-1_6b Stability AI Language Team (2024); Rafailov et al. (2024) stablelm 1B 2024-01-19 18239 -660.56 49.99
796 stabilityai/stablelm-2-12b-chat Stability AI Language Team (2024); Rafailov et al. (2024) stablelm 12B 2024-04-04 4843 -588.55 68.38
797 stabilityai/stablelm-2-1_6b-chat Stability AI Language Team (2024); Rafailov et al. (2024) stablelm 1B 2024-04-08 4156 -683.91 50.71
798 stabilityai/stablelm-3b-4e1t Ba et al. (2016); Zhang and Sennrich (2019); Rajbhandari et al. (2020); Gao et al. (2020); Black et al. (2022); Li et al. (2023a); Su et al. (2023a); Tow et al. (2023); Touvron et al. (2023b) stablelm 2B 2023-09-29 11112 -510.56 46.58
799 stabilityai/stable-code-3b Rajbhandari et al. (2020); Black et al. (2022); Li et al. (2023a); Su et al. (2023a); Touvron et al. (2023b); Pinnaparaju et al. (2024); Azerbayev et al. (2024); Yu et al. (2024b) stablelm 2B 2024-01-09 5670 -616.25 41.53
800 starmpcc/Asclepius-Llama2-13B Kweon et al. (2024) llama-2 13B 2023-09-19 1175 -645.45 50.25
801 starmpcc/Asclepius-Llama2-7B Kweon et al. (2024) llama-2 7B 2023-09-19 1231 -688.58 47.15
802 statking/zephyr-7b-sft-full-orpo mistral 7B 2024-05-18 2278 -548.89 53.16
803 StudentLLM/Alpagasus-2-13b-QLoRA-merged Chen et al. (2024a) llama 13B 2023-09-02 1303 -533.28 54.31
804 SuperAGI/SAM mistral 7B 2023-12-22 1235 -550.64 59.30
805 swap-uniba/LLaMAntino-3-ANITA-8B-Inst-DPO-ITA Basile et al. (2023); AI@Meta (2024); Polignano et al. (2024) llama-3 8B 2024-04-29 5995 -635.92 75.12
806 S4sch/zephyr-neural-chat-frankenmerge11b mistral 11B 2023-11-28 1181 -618.27 58.57
807 TaylorAI/Flash-Llama-13B llama 13B 2023-08-19 1181 -530.82 53.67
808 TaylorAI/Flash-Llama-3B llama 3B 2023-08-13 1177 -578.13 40.13
809 TaylorAI/Flash-Llama-7B llama 7B 2023-08-19 1184 -549.87 49.73
810 TeeZee/Bielik-SOLAR-LIKE-10.7B-Instruct-v0.1 mistral 10B 2024-04-10 1566 -854.79 53.50
811 TehVenom/Dolly_Malion-6b gptj 6B 2023-03-27 1201 -486.77 39.77
812 TehVenom/Dolly_Shygmalion-6b-Dev_V8P2 Wang and Komatsuzaki (2021) gptj 6B 2023-05-23 1197 -487.65 40.11
813 TehVenom/Dolly_Shygmalion-6b gptj 6B 2023-03-29 1195 -491.76 39.89
814 TehVenom/GPT-J-Pyg_PPO-6B-Dev-V8p4 gptj 6B 2023-03-26 1191 -493.17 39.61
815 TehVenom/GPT-J-Pyg_PPO-6B gptj 6B 2023-03-05 1203 -495.88 39.60
816 TehVenom/Metharme-13b-Merged llama-1 13B 2023-05-18 1199 -561.32 54.15
817 TehVenom/Moderator-Chan_GPT-JT-6b gptj 6B 2023-03-19 1185 -502.43 42.17
818 TehVenom/PPO_Pygway-V8p4_Dev-6b Wang and Komatsuzaki (2021) gptj 6B 2023-03-17 1191 -489.49 39.85
819 TehVenom/PPO_Shygmalion-V8p4_Dev-6b gptj 6B 2023-03-23 1188 -490.13 39.85
820 TehVenom/PPO_Shygmalion-6b gptj 6B 2023-03-23 1199 -489.16 39.35
821 TehVenom/Pygmalion-Vicuna-1.1-7b llama-1 6B 2023-05-02 1246 -572.52 49.25
822 TehVenom/Pygmalion-13b-Merged llama-1 13B 2023-05-18 1207 -588.35 48.49
823 TehVenom/Pygmalion_AlpacaLora-7b llama-1 7B 2023-04-30 1195 -605.84 46.49
824 teilomillet/MiniMerlin-3B llama 3B 2023-12-15 1168 -681.62 47.63
825 teknium/OpenHermes-13B llama-2 13B 2023-09-06 1594 -543.02 55.24
826 teknium/OpenHermes-2.5-Mistral-7B mistral 7B 2023-10-29 100997 -560.57 61.45
827 Telugu-LLM-Labs/Indic-gemma-7b-finetuned-sft-Navarasa-2.0 gemma 8B 2024-03-17 2025 -650.29 55.74
828 TencentARC/LLaMA-Pro-8B-Instruct llama-2 8B 2024-01-06 1423 -634.71 58.06
829 TencentARC/LLaMA-Pro-8B llama-2 8B 2024-01-05 1319 -557.58 51.67
830 TheBloke/airoboros-13B-HF llama-1 13B 2023-05-23 1214 -581.17 54.05
831 TheBloke/airoboros-7b-gpt4-fp16 llama-1 7B 2023-06-04 1208 -603.39 47.70
832 TheBloke/CodeLlama-13B-Instruct-fp16 llama-2 13B 2023-08-24 2193 -579.86 45.82
833 TheBloke/CodeLlama-13B-Python-fp16 llama-2 13B 2023-08-24 2114 -587.68 37.52
834 TheBloke/gpt4-alpaca-lora-13B-HF llama-1 13B 2023-04-17 1181 -547.42 53.98
835 TheBloke/guanaco-13B-HF llama-1 13B 2023-05-25 1222 -588.92 53.54
836 TheBloke/guanaco-7B-HF llama-1 7B 2023-05-25 1267 -586.15 47.34
837 TheBloke/koala-13B-HF llama-1 13B 2023-04-07 2501 -594.23 51.16
838 TheBloke/koala-7B-HF llama-1 7B 2023-04-07 1238 -617.95 44.29
839 TheBloke/Llama-2-13B-fp16 llama-2 13B 2023-07-18 6470 -530.82 53.67
840 TheBloke/Nous-Hermes-13B-SuperHOT-8K-fp16 llama-1 13B 2023-06-26 1225 -633.76 52.18
841 TheBloke/Planner-7B-fp16 llama-1 7B 2023-06-05 1219 -562.04 45.65
842 TheBloke/stable-vicuna-13B-HF von Werra et al. (2023); Anand et al. (2023); Chiang et al. (2023); Taori et al. (2023); Touvron et al. (2023a) llama-1 13B 2023-04-28 1337 -590.21 51.64
843 TheBloke/tulu-13B-fp16 Köpf et al. (2023); Conover et al. (2023); Wang et al. (2023c); Peng et al. (2023a); Longpre et al. (2023); Chaudhary (2023); Touvron et al. (2023a) llama-1 13B 2023-06-10 1231 -583.06 53.58
844 TheBloke/tulu-7B-fp16 Köpf et al. (2023); Conover et al. (2023); Wang et al. (2023c); Peng et al. (2023a); Longpre et al. (2023); Chaudhary (2023); Touvron et al. (2023a) llama-1 7B 2023-06-10 4079 -616.51 50.24
845 TheBloke/UltraLM-13B-fp16 Ding et al. (2023) llama-1 13B 2023-06-29 1211 -567.38 54.62
846 TheBloke/Vicuna-13B-CoT-fp16 Lacoste et al. (2019) llama-1 13B 2023-06-08 1219 -646.61 53.28
847 TheBloke/WizardLM-13B-V1-1-SuperHOT-8K-fp16 Xu et al. (2023a) llama-1 13B 2023-07-07 1225 -626.40 53.16
848 TheBloke/wizard-vicuna-13B-HF llama-1 13B 2023-05-04 1219 -610.98 52.75
849 TheBloke/Wizard-Vicuna-13B-Uncensored-HF llama-1 13B 2023-05-13 1648 -579.62 54.14
850 TheBloke/Wizard-Vicuna-7B-Uncensored-HF llama-1 7B 2023-05-18 1976 -607.14 48.27
851 TheTravellingEngineer/llama2-7b-chat-hf-dpo llama-2 7B 2023-08-14 1201 -659.39 50.38
852 TheTravellingEngineer/llama2-7b-chat-hf-guanaco llama-2 6B 2023-08-02 1209 -597.96 50.02
853 TheTravellingEngineer/llama2-7b-chat-hf-v2 llama-2 6B 2023-08-08 1208 -549.87 49.73
854 TheTravellingEngineer/llama2-7b-chat-hf-v3 llama-2 6B 2023-08-10 1203 -557.97 48.81
855 TheTravellingEngineer/llama2-7b-chat-hf-v4 llama-2 6B 2023-08-10 1213 -549.87 49.78
856 TheTravellingEngineer/llama2-7b-hf-guanaco llama-2 6B 2023-07-25 1206 -554.41 50.12
857 TigerResearch/tigerbot-7b-base llama 7B 2023-08-19 1228 -587.17 47.93
858 TIGER-Lab/MAmmoTH2-7B-Plus Yue et al. (2024b) mistral 7B 2024-05-06 10155 -673.56 67.75
859 TIGER-Lab/MAmmoTH2-8B-Plus Yue et al. (2024b) llama 8B 2024-05-06 12819 -595.71 67.49
860 TIGER-Lab/TIGERScore-13B Jiang et al. (2024) llama 13B 2023-11-26 1429 -548.56 56.79
861 tiiuae/falcon-rw-1b Brown et al. (2020); Dao et al. (2022); Press et al. (2022); Penedo et al. (2023) falcon 1B 2023-04-26 22921 -688.36 37.07
862 tiiuae/falcon-7b-instruct Shazeer (2019); Brown et al. (2020); Dao et al. (2022); Su et al. (2023a); Penedo et al. (2023); Almazrouei et al. (2023) falcon 7B 2023-04-25 179952 -621.07 43.16
863 tiiuae/falcon-7b Shazeer (2019); Brown et al. (2020); Gao et al. (2020); Dao et al. (2022); Su et al. (2023a); Penedo et al. (2023); Almazrouei et al. (2023) falcon 7B 2023-04-24 104010 -549.44 44.17
864 timpal0l/Mistral-7B-v0.1-flashback-v2 mistral 7B 2023-12-04 1275 -578.05 57.53
865 TinyLlama/TinyLlama-1.1B-Chat-v0.6 llama 1B 2023-11-20 15745 -642.54 34.94
866 TinyLlama/TinyLlama-1.1B-Chat-v1.0 llama 1B 2023-12-30 1078500 -619.71 37.17
867 TinyLlama/TinyLlama-1.1B-intermediate-step-1195k-token-2.5T llama 1B 2023-12-11 1365 -613.87 36.26
868 TinyLlama/TinyLlama-1.1B-intermediate-step-1431k-3T llama 1B 2023-12-28 798206 -608.46 36.42
869 TinyLlama/TinyLlama-1.1B-intermediate-step-955k-token-2T llama 1B 2023-11-19 8119 -646.34 34.56
870 togethercomputer/GPT-JT-6B-v0 gptj 6B 2022-11-22 1422 -484.70 44.05
871 togethercomputer/GPT-JT-6B-v1 Tay et al. (2022, 2023) gptj 6B 2022-11-24 5765 -503.02 43.13
872 togethercomputer/Llama-2-7B-32K-Instruct Liu et al. (2023) llama-2 7B 2023-08-08 5596 -564.83 50.02
873 togethercomputer/LLaMA-2-7B-32K llama-2 7B 2023-07-26 8884 -530.20 47.07
874 togethercomputer/Pythia-Chat-Base-7B gpt_neox 7B 2023-03-22 7040 -522.45 39.81
875 togethercomputer/RedPajama-INCITE-Base-3B-v1 gpt_neox 3B 2023-05-04 2635 -590.97 38.54
876 togethercomputer/RedPajama-INCITE-Chat-3B-v1 gpt_neox 3B 2023-05-05 1647 -610.76 39.53
877 togethercomputer/RedPajama-INCITE-Instruct-3B-v1 gpt_neox 3B 2023-05-05 2014 -552.27 39.06
878 togethercomputer/RedPajama-INCITE-7B-Base gpt_neox 7B 2023-05-04 1487 -560.75 41.49
879 togethercomputer/RedPajama-INCITE-7B-Chat gpt_neox 7B 2023-05-04 1583 -931.82 39.37
880 togethercomputer/RedPajama-INCITE-7B-Instruct gpt_neox 7B 2023-05-05 1274 -525.00 42.38
881 TomGrc/FusionNet_linear llama 10B 2023-12-31 1167 -561.26 74.43
882 TomGrc/FusionNet llama 10B 2023-12-31 1166 -561.26 74.38
883 totally-not-an-llm/EverythingLM-13b-16k llama-2 13B 2023-08-12 2132 -557.79 52.33
884 totally-not-an-llm/PuddleJumper-13b-V2 llama 13B 2023-09-21 1162 -633.45 54.19
885 Toten5/Marcoroni-neural-chat-7B-v2 mistral 7B 2023-12-12 1201 -557.34 72.50
886 TsinghuaC3I/Llama-3-8B-UltraMedical Zhang et al. (2024c) llama-3 8B 2024-04-27 3949 -558.79 63.73
887 Unbabel/TowerBase-7B-v0.1 Alves et al. (2024) llama 6B 2024-01-03 1856 -582.06 49.11
888 Unbabel/TowerInstruct-7B-v0.1 Alves et al. (2024) llama 6B 2024-01-04 2230 -590.30 52.39
889 Undi95/Meta-Llama-3-8B-hf AI@Meta (2024) llama-3 8B 2024-04-18 11376 -514.33 62.35
890 Undi95/Mistral-11B-v0.1 mistral 10B 2023-10-09 1196 -556.96 58.05
891 Undi95/MLewdBoros-L2-13B llama 13B 2023-09-09 1224 -542.84 56.51
892 Undi95/MLewd-Chat-v2-13B llama 13B 2023-09-26 1187 -569.88 57.23
893 Undi95/MLewd-L2-Chat-13B llama 13B 2023-09-16 1177 -551.09 57.75
894 Undi95/MLewd-L2-13B llama 13B 2023-09-04 1172 -666.07 53.12
895 Undi95/MLewd-v2.4-13B llama 13B 2023-09-26 1280 -571.03 56.37
896 Undi95/Nous-Hermes-13B-Code llama 13B 2023-09-02 1207 -602.47 55.93
897 Undi95/OpenRP-13B llama 13B 2023-09-11 1242 -536.49 56.57
898 Undi95/ReMM-SLERP-L2-13B llama 13B 2023-09-04 1511 -585.33 56.03
899 Undi95/ReMM-v2-L2-13B llama 13B 2023-09-09 1207 -565.06 56.99
900 Undi95/ReMM-v2.1-L2-13B llama 13B 2023-09-12 1217 -564.83 56.71
901 Undi95/ReMM-v2.2-L2-13B llama 13B 2023-09-21 1384 -567.22 57.10
902 Undi95/UndiMix-v1-13b llama 13B 2023-08-31 1220 -649.19 55.50
903 Undi95/UndiMix-v4-13B llama 13B 2023-09-12 1212 -569.06 56.93
904 Undi95/Unholy-v1-12L-13B llama 13B 2023-09-10 1210 -541.04 57.47
905 Undi95/X-MythoChronos-13B llama 13B 2023-11-18 1203 -604.69 58.43
906 unsloth/mistral-7b-v0.2 mistral 7B 2024-03-24 4119 -532.63 60.34
907 unsloth/Phi-3-medium-4k-instruct mistral 13B 2024-05-23 3994 -552.43 73.57
908 unsloth/Phi-3-mini-4k-instruct mistral 3B 2024-04-29 12224 -575.74 69.86
909 unsloth/tinyllama-chat llama 1B 2024-02-14 5959 -619.71 37.24
910 upstage/SOLAR-10.7B-Instruct-v1.0 Kim et al. (2024b, a) llama 10B 2023-12-12 67725 -557.67 74.20
911 upstage/SOLAR-10.7B-v1.0 Kim et al. (2024b) llama 10B 2023-12-12 24478 -525.51 66.04
912 uukuguy/speechless-code-mistral-7b-v1.0 mistral 7B 2023-10-10 4393 -540.50 58.85
913 uukuguy/speechless-llama2-hermes-orca-platypus-wizardlm-13b Touvron et al. (2023b) llama-2 13B 2023-09-01 2101 -609.55 57.52
914 uukuguy/speechless-llama2-hermes-orca-platypus-13b Touvron et al. (2023b) llama-2 13B 2023-09-01 1368 -587.97 57.17
915 uukuguy/speechless-zephyr-code-functionary-7b mistral 7B 2024-01-23 4098 -529.18 62.93
916 uukuguy/zephyr-7b-alpha-dare-0.85 mistral 7B 2023-11-23 6141 -529.44 62.35
917 uygarkurt/llama-3-merged-linear llama-3 8B 2024-05-09 12676 -570.63 73.93
918 VAGOsolutions/Llama-3-SauerkrautLM-8b-Instruct llama-3 8B 2024-04-19 55400 -579.40 73.74
919 VAGOsolutions/SauerkrautLM-Gemma-7b gemma 8B 2024-02-27 5691 -561.27 67.83
920 VAGOsolutions/SauerkrautLM-SOLAR-Instruct llama 10B 2023-12-20 1175 -560.76 74.21
921 VAGOsolutions/SauerkrautLM-7b-HerO mistral 7B 2023-11-24 1226 -553.88 64.49
922 varox34/Bio-Saul-Dolphin-Beagle-Breadcrumbs mistral 7B 2024-05-01 2679 -610.01 48.72
923 vibhorag101/llama-2-13b-chat-hf-phr_mental_therapy llama-2 13B 2023-09-17 1269 -630.34 42.50
924 vicgalle/CarbonBeagle-11B-truthy mistral 10B 2024-02-10 13887 -554.84 76.10
925 vicgalle/CarbonBeagle-11B Wortsman et al. (2022) mistral 10B 2024-01-21 6696 -549.83 74.64
926 vicgalle/ConfigurableBeagle-11B Gallego (2024) mistral 10B 2024-02-17 6045 -548.90 75.40
927 vicgalle/ConfigurableHermes-7B Gallego (2024) mistral 7B 2024-02-17 6085 -572.68 68.89
928 vicgalle/ConfigurableSOLAR-10.7B Gallego (2024) llama 10B 2024-03-10 5135 -559.06 73.94
929 vicgalle/Configurable-Hermes-2-Pro-Llama-3-8B Gallego (2024) llama-3 8B 2024-05-02 10273 -580.31 70.10
930 vicgalle/Configurable-Llama-3-8B-v0.3 Gallego (2024) llama-3 8B 2024-04-20 6030 -555.07 68.79
931 vicgalle/Configurable-Yi-1.5-9B-Chat Gallego (2024) llama 8B 2024-05-12 6537 -633.71 70.50
932 viethq188/LeoScorpius-7B mistral 7B 2023-12-12 1194 -552.34 72.21
933 viethq188/Rabbit-7B-v2-DPO-Chat mistral 7B 2023-12-12 1181 -579.23 69.36
934 vihangd/dopeyplats-1.1b-2T-v1 llama 1B 2023-11-26 1191 -682.92 35.28
935 vihangd/dopeyshearedplats-1.3b-v1 llama-2 1B 2023-12-12 1168 -766.02 36.74
936 vihangd/dopeyshearedplats-2.7b-v1 llama-2 2B 2023-12-16 1173 -709.66 42.90
937 vihangd/neuralfalcon-1b-v1 falcon 1B 2023-12-17 1185 -895.18 29.72
938 vihangd/shearedplats-1.3b-v1 llama-2 1B 2023-11-16 1184 -706.16 35.97
939 vihangd/shearedplats-2.7b-v2 llama-2 2B 2023-11-18 2081 -647.66 41.61
940 vihangd/smartyplats-3b-v1 llama 3B 2023-09-11 1180 -600.04 40.00
941 vihangd/smartyplats-3b-v2 llama 3B 2023-09-14 1179 -597.07 40.29
942 vikash06/llama-2-7b-small-model-new llama-2 6B 2023-12-22 1174 -925.37 46.62
943 vmajor/Orca2-13B-selfmerge-26B llama 13B 2023-12-01 2041 -653.10 62.24
944 vmajor/Orca2-13B-selfmerge-39B llama 13B 2023-12-01 1208 -653.10 62.24
945 VMware/open-llama-0.7T-7B-open-instruct-v1.1 llama-1 7B 2023-05-31 1161 -652.26 41.11
946 VMware/open-llama-7b-open-instruct llama-1 7B 2023-06-08 6470 -642.10 42.59
947 Voicelab/trurl-2-13b-academic llama-2 13B 2023-09-18 2774 -568.46 53.94
948 Voicelab/trurl-2-7b llama-2 7B 2023-08-16 2861 -602.32 50.58
949 vonjack/Qwen-LLaMAfied-HFTok-7B-Chat llama-2 7B 2023-08-09 1189 -839.96 50.64
950 v1olet/v1olet_marcoroni-go-bruins-merge-7B mistral 7B 2023-12-11 1225 -559.92 72.81
951 v1olet/v1olet_merged_dpo_7B mistral 7B 2023-12-12 1210 -592.71 70.26
952 wang7776/Llama-2-7b-chat-hf-10-sparsity Sun et al. (2024) llama-2 6B 2023-12-11 1178 -660.87 52.48
953 wang7776/Llama-2-7b-chat-hf-20-sparsity Sun et al. (2024) llama-2 7B 2023-12-13 1175 -668.34 52.01
954 wang7776/Llama-2-7b-chat-hf-30-sparsity Sun et al. (2024) llama-2 6B 2023-12-11 1179 -676.50 51.02
955 webbigdata/ALMA-7B-Ja-V2 Xu et al. (2024a) llama-2 7B 2023-10-21 1177 -599.69 47.85
956 wei123602/Llama-2-13b-FINETUNE4_TEST llama-2 13B 2023-09-18 1174 -538.02 53.62
957 wenbopan/Faro-Yi-9B-DPO OpenAI et al. (2024) llama 8B 2024-04-07 6121 -597.40 68.77
958 wenbopan/Faro-Yi-9B OpenAI et al. (2024) llama 8B 2024-03-27 6127 -594.07 66.37
959 wenge-research/yayi-7b-llama2 llama-2 7B 2023-07-21 1195 -562.60 49.88
960 wenge-research/yayi-7b bloom 7B 2023-06-02 1192 -653.40 41.88
961 Weyaxi/ChatAYT-Lora-Assamble-Marcoroni llama 13B 2023-09-14 1203 -569.45 57.76
962 Weyaxi/Dolphin2.1-OpenOrca-7B mistral 7B 2023-10-11 1221 -557.64 60.47
963 Weyaxi/Instruct-v0.2-Seraph-7B mistral 7B 2023-12-12 1203 -574.63 68.48
964 Weyaxi/Luban-Marcoroni-13B-v2 llama 13B 2023-09-13 1212 -566.43 57.92
965 Weyaxi/Luban-Marcoroni-13B-v3 llama 13B 2023-09-13 1214 -566.44 57.94
966 Weyaxi/MetaMath-Chupacabra-7B-v2.01-Slerp mistral 7B 2023-12-08 1210 -546.54 70.26
967 Weyaxi/MetaMath-NeuralHermes-2.5-Mistral-7B-Linear mistral 7B 2023-12-05 1205 -560.95 67.60
968 Weyaxi/MetaMath-NeuralHermes-2.5-Mistral-7B-Ties mistral 7B 2023-12-05 1212 -604.36 67.03
969 Weyaxi/MetaMath-neural-chat-7b-v3-2-Slerp mistral 7B 2023-12-08 1208 -544.26 69.79
970 Weyaxi/MetaMath-neural-chat-7b-v3-2-Ties mistral 7B 2023-12-05 1212 -590.97 67.54
971 Weyaxi/MetaMath-OpenHermes-2.5-neural-chat-v3-3-Slerp mistral 7B 2023-12-10 1224 -548.15 69.92
972 Weyaxi/MetaMath-Tulpar-7b-v2-Slerp mistral 7B 2023-12-08 1217 -555.63 70.07
973 Weyaxi/MetaMath-una-cybertron-v2-bf16-Ties mistral 7B 2023-12-06 1218 -602.74 68.88
974 Weyaxi/neural-chat-7b-v3-1-OpenHermes-2.5-7B mistral 7B 2023-12-01 1205 -570.29 67.19
975 Weyaxi/openchat-3.5-1210-Seraph-Slerp mistral 7B 2023-12-27 1210 -553.81 71.82
976 Weyaxi/OpenHermes-2.5-neural-chat-7b-v3-1-7B mistral 7B 2023-11-24 1230 -567.74 67.84
977 Weyaxi/OpenHermes-2.5-neural-chat-7b-v3-2-7B mistral 7B 2023-12-03 1221 -576.02 68.71
978 Weyaxi/OpenOrca-Zephyr-7B mistral 7B 2023-10-11 1208 -564.12 64.97
979 Weyaxi/Samantha-Nebula-7B mistral 7B 2023-10-05 1166 -584.51 54.58
980 Weyaxi/SauerkrautLM-UNA-SOLAR-Instruct llama 10B 2023-12-21 1225 -561.98 74.26
981 Weyaxi/Seraph-openchat-3.5-1210-Slerp mistral 7B 2023-12-27 1210 -567.06 70.89
982 Weyaxi/Seraph-7B mistral 7B 2023-12-11 1214 -546.73 71.86
983 Weyaxi/SlimOpenOrca-Mistral-7B mistral 7B 2023-10-11 1220 -587.26 60.84
984 WhiteRabbitNeo/WhiteRabbitNeo-13B-v1 llama-2 13B 2023-12-17 2090 -615.73 49.11
985 winglian/Llama-2-3b-hf llama-2 3B 2023-09-19 1565 -1376.16 29.53
986 winglian/llama-2-4b llama-2 4B 2023-09-19 1193 -676.31 34.23
987 w601sxs/b1ade-1b gpt_neox 1B 2023-07-17 1204 -929.29 32.59
988 xDAN-AI/xDAN-L1-Chat-RL-v1 mistral 7B 2023-12-20 1178 -574.92 68.38
989 Xwin-LM/Xwin-LM-13B-V0.1 Xwin-LM Team (2023) llama-2 13B 2023-09-15 1401 -559.13 55.29
990 Xwin-LM/Xwin-LM-7B-V0.1 Xwin-LM Team (2023) llama-2 7B 2023-09-15 1474 -592.93 52.08
991 yam-peleg/Hebrew-Mistral-7B mistral 7B 2024-04-26 5993 -625.28 58.76
992 yanolja/Bookworm-10.7B-v0.4-DPO Mukherjee et al. (2023); Lian et al. (2023g); Cui et al. (2024) llama 10B 2024-01-18 2239 -586.77 66.59
993 yanolja/EEVE-Korean-Instruct-10.8B-v1.0 Lian et al. (2023g); Mukherjee et al. (2023); Kim et al. (2024c); Cui et al. (2024) llama 10B 2024-02-22 13241 -555.03 66.48
994 yeen214/llama2_7b_small_tuning_v1 llama-2 7B 2023-10-02 3289 -1402.04 28.56
995 yeen214/test_llama2_ko_7b llama-2 7B 2023-10-02 3285 -1405.43 29.99
996 yeen214/test_llama2_7b llama-2 7B 2023-09-30 3289 -549.87 49.73
997 yhyhy3/open_llama_7b_v2_med_instruct llama-1 7B 2023-07-09 1197 -561.62 46.24
998 Yhyu13/chimera-inst-chat-13b-hf llama-1 13B 2023-05-11 1295 -582.39 52.86
999 Yhyu13/LMCocktail-10.7B-v1 Xiao et al. (2023) llama-2 10B 2023-12-20 3332 -556.38 74.06
1000 Yukang/Llama-2-7b-longlora-32k-ft Chen et al. (2024b) llama-2 7B 2023-09-12 2423 -1387.05 29.20
1001 Yukang/LongAlpaca-13B Chen et al. (2023b, 2024b) llama 13B 2023-10-08 1896 -832.57 41.74
1002 Yukang/LongAlpaca-7B Chen et al. (2023b, 2024b) llama 6B 2023-10-07 2773 -794.46 39.36
1003 yulan-team/YuLan-Chat-2-13b-fp16 YuLan-Team (2023) llama 13B 2023-08-04 1165 -645.56 57.01
1004 yunconglong/DARE_TIES_13B Yadav et al. (2023a); Yu et al. (2024a) mixtral 12B 2024-01-30 7029 -611.23 77.10
1005 yunconglong/MoE_13B_DPO mixtral 12B 2024-01-28 3939 -603.88 77.05
1006 yunconglong/Truthful_DPO_TomGrc_FusionNet_7Bx2_MoE_13B mixtral 12B 2024-01-21 7965 -598.84 77.44
1007 01-ai/Yi-1.5-6B-Chat 01. AI et al. (2025) llama 6B 2024-05-11 19443 -645.33 66.17
1008 01-ai/Yi-1.5-6B 01. AI et al. (2025) llama 6B 2024-05-11 5102 -525.57 61.57
1009 01-ai/Yi-1.5-9B-Chat-16K 01. AI et al. (2025) llama 8B 2024-05-15 14262 -601.39 66.98
1010 01-ai/Yi-1.5-9B-Chat 01. AI et al. (2025) llama 8B 2024-05-10 20108 -626.90 69.56
1011 01-ai/Yi-1.5-9B-32K 01. AI et al. (2025) llama 8B 2024-05-15 9287 -510.38 55.22
1012 01-ai/Yi-1.5-9B 01. AI et al. (2025) llama 8B 2024-05-11 20252 -513.39 66.73
1013 01-ai/Yi-6B-200K Zhang et al. (2024b); Yue et al. (2024a); 01. AI et al. (2025) llama 6B 2023-11-06 8370 -529.12 56.76
1014 01-ai/Yi-6B Zhang et al. (2024b); Yue et al. (2024a); 01. AI et al. (2025) llama 6B 2023-11-01 6997 -515.62 54.02
1015 01-ai/Yi-9B-200K Zhang et al. (2024b); Yue et al. (2024a); 01. AI et al. (2025) llama 8B 2024-03-15 8915 -517.82 61.94
1016 42dot/42dot_LLM-PLM-1.3B 42dot Inc. (2023) llama 1B 2023-09-04 1268 -600.03 35.70
1017 42dot/42dot_LLM-SFT-1.3B 42dot Inc. (2023) llama 1B 2023-09-04 1453 -605.30 36.61
1018 922-CA/monika-ddlc-7b-v1 llama-2 7B 2023-10-13 1191 -635.19 50.49